### v1

# Limit Laws for Poincaré Recurrence and the Shrinking Target Problem

Abstract.
Let $(X,T,\mu,d)$ be a metric measure-preserving system. If $B(x,r_{n}(x))$ is a sequence of balls such that, for each $n$, the measure of $B(x,r_{n}(x))$ is constant, then we obtain a self-norming CLT for recurrence for systems satisfying a multiple decorrelation property. When $\mu$ is absolutely continuous, we obtain a distributional limit law for recurrence for the sequence of balls $B(x,r_{n})$. In the latter case, the density of the limiting distribution is an average over Gaussian densities. An important assumption in the CLT for recurrence is that the CLT holds for the shrinking target problem. Because of this, we also prove an ASIP for expanding and Axiom A systems for non-autonomous Hölder observables and apply it to the shrinking target problem, thereby obtaining a CLT.
Key words and phrases:
Recurrence, Central limit theorem, Shrinking target problem, Almost sure invariance principle
2020 Mathematics Subject Classification:
37B20 (Primary); 37D20, 37A05 (Secondary)
1. Introduction
A fundamental result in dynamical systems is the Poincaré Recurrence Theorem, which gives conditions for when a given point returns infinitely often to a shrinking neighbourhood of its initial position. More precisely, if $(X,T,\mu,d)$ is a separable metric measure-preserving system with Borel measure $\mu$, then $\mu$-a.e. point $x\in X$ is recurrent, meaning that $\liminf_{n\to\infty}d(x,T^{n}x)=0$. While this theorem offers a qualitative conclusion under minimal assumptions, it leaves open the question of quantitative recurrence. For example, can one – perhaps under additional conditions – establish the rate at which $T^{n}x$ returns to $x$? This question has motivated a significant body of recent work in quantitative recurrence.
The first result establishing a rate of convergence of $\liminf_{n\to\infty}d(x,T^{n}x)=0$ is due to Boshernitzan [3] in 1993 and, similar to the Poincaré Recurrence theorem, has minimal assumptions. Namely, under the additional assumption that there exists $\alpha>0$ such that $X$ is $\sigma$-finite with respect to the Hausdorff measure $H_{\alpha}$, one has that
(1.1) $\liminf_{n\to\infty}n^{-1/\alpha}d(x,T^{n}x)<\infty\quad\text{$\mu$-a.e.\ $x$}.$
Following Boshernitzan’s work, Barreira and Saussol [1] and Urbánski [38] obtain a connection between recurrence rates and the pointwise dimension of the measure for various systems. More than 20 years since Boshernitzan’s pioneering work, several authors have improved the conclusion by imposing stricter conditions on the underlying systems, see for instance [30, 21, 23, 24].
Another way of viewing (1.1) is the following: for $\mu$-a.e. $x$, there exists $C(x)>0$ such that
$\sum_{k=1}^{\infty}\mathbf{1}_{B(x,C(x)k^{-1/\alpha})}(T^{k}x)=\infty.$
More generally, one may consider the sum
(1.2) $\sum_{k=1}^{n}\mathbf{1}_{B(x,r_{k}(x))}(T^{k}x),$
for more general sequences of radii $r_{n}(x)$. When the $r_{n}$ are such that $\mu(B(x,r_{n}(x)))$ is constant for each $n$, recent results [33, 35, 34] establish the almost-sure asymptotic behaviour of (1.2) in the form of Strong Borel–Cantelli lemmas for recurrence (RSBC). More precisely, the authors obtain that, for a class of systems satisfying exponential decay of $3$-fold correlations,
$\lim_{n\to\infty}\frac{\sum_{k=1}^{n}\mathbf{1}_{B(x,r_{k}(x))}(T^{k}x)}{\sum_{k=1}^{n}\mu\bigl(B(x,r_{k}(x))\bigr)}=1$
under some decay conditions on the sequence $M_{n}=\mu(B(x,r_{n}(x)))$. Persson and the author [34] also establish an error term. Considering radii $r_{n}$ that do not depend on the centre of the ball, He [15], when $\mu$ is absolutely continuous, and Levesley, Li, Simmons, Velani [27], when $\mu=m$ is the Lebesgue measure, obtain RSBCs for more general sets with error terms, while also relaxing conditions on the decay of $r_{n}$. In particular, we bring attention to the main result of He, which restricted to balls implies that
(1.3) $\lim_{n\to\infty}\frac{\sum_{k=1}^{n}\mathbf{1}_{B(x,r_{k})}(T^{k}x)}{\sum_{k=1}^{n}m(B(x,r_{k}))}=h(x)\quad\text{$\mu$-a.e.\ $x$},$
whenever $r_{k}\to 0$ and $\sum_{k=1}^{\infty}m(B(x,r_{k}))=\infty$.
One may view a Strong Borel–Cantelli lemma as a non-autonomous Strong Law of Large Numbers. As such, it is natural to ask if the sum in (1.2) satisfies a distributional law when normalized appropriately. Answering this question is the primary purpose of this paper. With this in mind, Theorem 3.3 in Conze and Le Borgne [9] can be applied to obtain a distributional law for the sum in (1.2) for a class of systems when $X$ is a compact group and $r_{k}(x)=r$ is constant in $x\in X$ and $k\geq 0$. More precisely, when $X$ is a compact group and the system $(X,T,\mu)$ satisfies a multiple decorrelation property, one can apply Conze and Le Borgne’s result to obtain that $x\mapsto n^{-1/2}\sum_{k=1}^{n}(\mathbf{1}_{B(x,r)}\circ T^{k}x-\mu(B(x,r)))$ converges in distribution to a random variable with characteristic function
$t\mapsto\int_{X}\exp\Bigl(-\frac{t^{2}}{2\sigma_{x}^{2}}\Bigr)\mathop{}\!d\mu(x),$
where $\sigma_{x}^{2}=\lim_{n\to\infty}n^{-1}\operatorname{Var}(\sum_{k=1}^{n}\mathbf{1}_{B(x,r)}\circ T^{k})$ is the asymptotic variance of $\mathbf{1}_{B(x,r)}$. This suggests that, in the more general setting, the sum in (1.2) satisfies a similar distributional law.
At the same time, Holland and Todd [20] proved recently that, for measure-preserving systems with absolutely continuous measure $\mu$ satisfying exponential decay of correlations for $\mathrm{BV}$ versus $L^{1}(\mu)$ observables, the sum $\sum_{k=1}^{n}\mathbf{1}_{B(x,r_{n})}(T^{k}x)$ converges in distribution to an averaged Poisson distribution, where $r_{n}=\tau/2n$ for some constant $\tau>0$. More precisely, Holland and Todd show that
(1.4) $\lim_{n\to\infty}\mu\biggl\{x\in X:\sum_{j=1}^{n}\mathbf{1}_{B(x,r_{n})}(T^{j}x)=k\biggr\}=\int_{X}\frac{\tau^{k}h(x)^{k+1}}{k!}e^{-h(x)\tau}\mathop{}\!dx,$
where $h$ is the density of $\mu$. The appearance of the density of $\mu$ in both (1.3) and (1.4) suggests that the distributional limit of (1.2) should involve the density, at least in the case when $r_{n}$ does not depend on the centre of the ball.
Note the similarity between the recurrence problem of studying the asymptotic behaviour of the left-hand side of (1.2) and the shrinking target problem, which concerns – among other things – the asymptotic behaviour of the sum
(1.5) $\sum_{k=1}^{n}\mathbf{1}_{B(y_{k},r_{k})}\circ T^{k},$
where $B(y_{k},r_{k})$ is a fixed sequence of shrinking balls. The shrinking target problem has been studied for a large range of systems. We bring particular attention to the work of Chernov and Kleinbock [6] who prove a Strong Borel–Cantelli lemma for Axiom A diffeomorphisms, and Haydn, Nicol, Török and Vaienti [13], who improve upon [14] and prove an almost sure invariance principle (see definition below) for the shrinking target problem in the setting of expanding interval maps. The secondary purpose of this paper is to establish an almost sure invariance principle for the shrinking target problem for a larger class of measure-preserving systems including Axiom A systems. In fact, this will be used in establishing the distributional law for recurrence.
We now state the limit laws that we will concern ourselves with when considering sums of the form (1.2) and (1.5). We take inspiration from the probabilistic setting and consider two limit laws of increasing strength. Suppose $(\Omega,\mathbf{P})$ is a probability space. We say that a sequence of real-valued random variables $(X_{k})$ with $\sigma_{n}^{2}=\operatorname{Var}(\sum_{k=1}^{n}X_{k})$ satisfies the (self-norming) central limit theorem (CLT) if $\sigma_{n}^{-1}\sum_{k=1}^{n}\tilde{X}_{k}$, where $\tilde{X}=X-\mathbf{E}[X]$, converges in distribution to a standard Gaussian variable, i.e., if
(CLT) $\lim_{n\to\infty}\mathbf{P}\biggl\{\frac{1}{\sigma_{n}}\sum_{k=1}^{n}\tilde{X}_{k}<t\biggr\}=\frac{1}{\sqrt{2\pi}}\int_{-\infty}^{t}e^{-\frac{s^{2}}{2}}\mathop{}\!ds\quad\text{for all $t\in\mathbb{R}$}.$
Alternatively, Lévy’s continuity theorem ([22, Theorem 5.3]) implies that the condition (CLT) is equivalent to
$\lim_{n\to\infty}\mathbf{E}\biggl[\exp\biggl(\frac{it}{\sigma_{n}}\sum_{k=1}^{n}\tilde{X}_{k}\biggr)\biggr]=e^{-\frac{t^{2}}{2}}\quad\text{for all $t\in\mathbb{R}$}.$
A stronger result than the CLT, is the almost sure invariance principle (ASIP), which can be stated simply as an almost sure version of the CLT. The particular version of the ASIP we are interested in is the following. We say that a sequence of random variables $(X_{n})$ satisfies the ASIP with rate $\beta\in(0,1)$ if there exists a sequence of standard Gaussian random variables $(Z_{n})$ (by enlarging the underlying probability space if necessary) such that
(ASIP) $\sup_{1\leq k\leq n}\biggl|\sum_{i=1}^{k}\tilde{X}_{i}-\sum_{i=1}^{k}Z_{i}\biggr|=o(\sigma_{n}^{1-\beta})\quad\text{a.s.}$
and $\sum_{k=1}^{n}\mathbf{E}[Z_{k}^{2}]\sim\sigma_{n}^{2}$, where $\sim$ denotes asymptotic equivalence. The ASIP is a strong asymptotic estimate and implies other limit laws, such as a functional CLT and the law of the iterated logarithm (LIL).
In this paper, we are interested in the case when our underlying space is an m.m.p.s. $(X,T,\mu,d)$ and the random variables $X_{k}$ correspond to observables related to the recurrence or the shrinking target problem. The primary purpose of the paper is to establish distributional laws for the recurrence problem; the secondary purpose is to establish (ASIP) for the shrinking target problem.
Our first main result is Theorem 2.1, which establishes (CLT) for the sequence $(\mathbf{1}_{B(x,r_{n}(x))}(T^{n}x))_{n}$ for certain implicitly defined $r_{n}\colon X\to(0,\infty)$ under decay conditions on the measure of the balls $B(x,r_{n}(x))$. The main assumptions include exponential decay of multiple correlations and a thin annuli assumption for $\mu$. These properties are satisfied for various expanding and uniformly hyperbolic systems. In Theorem 2.2, under the assumption that the variance is well behaved, we show that the observables $(\mathbf{1}_{B(x,r_{n})}(T^{n}x))_{n}$ for sequences of scalars $(r_{n})$ satisfies a distributional law that is in some sense an average of Gaussian distributions. In particular in Corollary 2.1, when $\mu$ is absolutely continuous, we explicitly obtain an expression for the limiting distribution, which involves the density of $\mu$. An important ingredient in the proofs of these results is that the shrinking target problem satisfies a CLT, which motivates the secondary purpose of this paper.
Our third and fourth main results are Theorems 2.3 and 2.5, which imply (ASIP) for the shrinking target problem for higher dimensional expanding and uniformly hyperbolic systems (stated explicitly in Corollaries 2.2 and 2.3). More precisely, we establish, for $\mu$-a.e. $y$, (ASIP) for the sequences $(\mathbf{1}_{B(y,r_{n}(y))}\circ T^{n})_{n}$ and $(\mathbf{1}_{B(y,r_{n})}\circ T^{n})_{n}$ for non-invertible systems with a spectral gap and for Axiom A diffeomorphisms equipped with a Gibbs measure. This verifies one of the main assumptions of the results of the preceding paragraph.
Theorems 2.3 and 2.5 extend that of Haydn, Nicol, Török and Vaienti [13, Theorem 6.1, Corollary 6.2] who employ an abstract framework due to Cuny and Merlevède [10, Theorem 2.3] (see Theorem 4.1), in order to prove (ASIP) for non-stationary Hölder continuous observables on non-invertible and invertible hyperbolic systems, respectively. The main drawback of their results is the requirement that the observables have uniformly bounded Hölder norm, which prevents the application of their results to the shrinking target problem. Nevertheless, the proofs of Theorems 2.3 and 2.5 are modifications of [13, Theorem 6.1, Corollary 6.2]. By improving their results, we also manage to prove a CLT for the shrinking target problem in the setting of uniformly hyperbolic systems, which hitherto had not been proven. We also note that Chen, Yang and Zhang [5] establish (ASIP) for non-uniformly hyperbolic systems equipped with the absolutely continuous invariant measure. We delay a more in-detail discussion of these results for when we state our results.
1.1. Structure of the Paper
We begin with establishing relevant notation and stating our assumptions in Section 2. We then state our results on the distributional law for the recurrence problem and the ASIP for non-autonomous observables, which is then applied to the shrinking target problem. In Section 3, we state various results on the behaviour of the variances that appear in the both the recurrence problem and shrinking target problem. Section 4 is devoted to stating various supporting results that are mostly found elsewhere. We devote Sections 58 to the proofs of the CLT, ASIP and variance results. Section 9 is devoted to a discussion on assumptions used in Theorem 2.1, Corollary 2.1 and variance Lemmas 3.4 and 3.6. More precisely, the section is devoted to establishing a short returns estimate when $\mu$ is absolutely continuous.
2. Main Results
2.1. Assumptions
Throughout the paper, we assume that $(X,d)$ is a $\dim X$-dimensional smooth compact Riemannian manifold, and that the quintuplet $(X,T,\mu,\mathcal{F},d)$ is a metric measure-preserving dynamical system (m.m.p.s.). That is, $(X,\mu,\mathcal{F})$ is a probability space, $\mu$ is a Borel measure and $T\colon X\to X$ is a measurable transformation which preserves $\mu$. Unless relevant, we choose to suppress the mention of the $\sigma$-algebra $\mathcal{F}$. We denote the normalised transfer operator corresponding to the system by $P$, i.e., the $P\colon L^{2}(\mu)\to L^{2}(\mu)$ is the operator defined by $\int_{X}\varphi(\psi\circ T)\mathop{}\!d\mu=\int_{X}(P\varphi)\psi\mathop{}\!d\mu$ for all $\varphi,\psi\in L^{2}(\mu)$.
For an observable $f\colon X\to\mathbb{R}$, we write $\mu(f)=\int_{X}f\mathop{}\!d\mu$ and we denote by $\tilde{f}$ the centred observable, i.e., $\tilde{f}=f-\mu(f)$. If $g\colon X\to\mathbb{R}$ is another observable, we define $\operatorname{Cov}(f,g)=\mu(fg)-\mu(f)\mu(g)$ and $\operatorname{Var}(f)=\operatorname{Cov}(f,f)$. By the characteristic function of $f$, we mean the function $\psi_{f}\colon\mathbb{R}\to\mathbb{C}$ given by $\psi_{f}(t)=\int_{X}\exp(itf(x))\mathop{}\!d\mu(x)$. For $\alpha\in(0,1]$, we denote by $\mathcal{H}_{\alpha}$ the space of real-valued $\alpha$-Hölder continuous functions on $X$ and equip it with the norm $\|\cdot\|_{\alpha}$ defined by $\|f\|_{\alpha}=\|f\|_{\infty}+\lvert f\rvert_{\alpha}$ for $f\in\mathcal{H}_{\alpha}$, where $\|f\|_{\infty}:=\sup_{x\in X}|f(x)|$ and
$\lvert f\rvert_{\alpha}:=\sup_{x\neq y}\frac{|f(x)-f(y)|}{d(x,y)^{\alpha}}.$
For a set $A$, denote by #A\#{A} the cardinality of $A$ and, for a set $Q\subset X$, let $\lvert Q\rvert:=\sup\{d(x,y):x,y\in Q\}$ be the diameter of $Q$. If $\mathcal{Q}$ is a collection of subsets of $X$, we define $\lvert\mathcal{Q}\rvert:=\sup_{Q\in\mathcal{Q}}\lvert Q\rvert$. We write $a\lesssim_{x}b$ for quantities $a,b,x$ if there exists a constant $C>0$ that depends on $x$ such that $a\leq Cb$. We write $a\lesssim b$ if $C>0$ is a universal constant, i.e., it depends only on the system $(X,T,\mu,d)$. We define $\gtrsim$ analogously, and $a\asymp b$ if $a\lesssim b$ and $a\gtrsim b$. For sequences $a_{n}$ and $b_{n}$, we write $a_{n}\sim b_{n}$ to mean that $a_{n}$ and $b_{n}$ are asymptotically equivalent, i.e., $a_{n}/b_{n}\to 1$.
Throughout the article, $(r_{n})$ will denote a sequence of radius functions $r_{n}\colon X\to(0,\infty)$ and we will consider the corresponding recurrence sets $E_{n}=\{x\in X:T^{n}x\in B(x,r_{n}(x))\}$, where $B(x,r)$ is the open ball with centre $x$ and radius $r$. We will predominantly consider two cases: one in which the functions $(r_{n})$ are defined implicitly with reference to a given sequence $(M_{n})$; and another in which $(r_{n})$ is a sequence of scalars, given without a reference to another sequence. More precisely, in the first case, given a sequence of real numbers $(M_{n})$ in $(0,1]$, define – whenever possible – the functions $(r_{n})$ by $\mu(B(x,r_{n}(x)))=M_{n}$ for all $x\in X$ and $n\geq 0$. We will broadly refer to the first case as the implicit radii problem, and to the second case as the explicit radii problem. In the latter problem, we will denote the recurrence sets as $\hat{E}_{n}=\{x\in X:T^{n}x\in B(x,r_{n})\}$ to clearly delineate between the problems. We will also denote related quantities with a hat. When considering explicit radii, we will find it useful to consider the sequence $(\hat{M}_{n})$ given by $\hat{M}_{n}=\int\mu(B(x,r_{n}))\mathop{}\!d\mu(x)$. In the implicit radii setting, we will see that the functions $r_{n}$ are well defined under certain assumptions on $\mu$ and that $r_{n}\to 0$ uniformly on $\operatorname{supp}\mu$. Note that, since $X$ is compact, we have that $r_{k}(x)\leq\lvert X\rvert$, whenever $r_{k}(x)$ is defined.
We will frequently, though not always, rely on the following condition for the sequences $(M_{n})$ and $(\hat{M}_{n})$, where $\gamma,\upsilon\geq 0$ are constants to be specified.
Condition $\mathcal{S}_{\gamma,\upsilon}$.
The sequence $(a_{n})$ is monotone decreasing and, for all $n\in\mathbb{N}$,
$\frac{1}{n^{\gamma}}\lesssim a_{n}\lesssim\frac{1}{(\log n)^{\upsilon}}.$
We restrict ourselves to rapidly mixing systems. At different times throughout the article, we consider three different decorrelation properties.
Condition D.1 (Transfer Operator).
For every $\alpha\in(0,1]$, the transfer operator $P$ restricts to an operator $P\colon\mathcal{H}_{\alpha}\to\mathcal{H}_{\alpha}$ such that there exists $\tau_{0}>0$ for which
$\|P^{n}\varphi\|_{\alpha}\lesssim_{\alpha}\|\varphi\|_{\alpha}e^{-\tau_{0}n}$
for all $n\in\mathbb{N}$ and all $\varphi\in\mathcal{H}_{\alpha}$ with $\mu(\varphi)=0$.
Condition D.1 implies decay of correlations for Hölder versus $L^{1}$ observables, i.e., for all $\varphi\in\mathcal{H}_{\alpha}$ and $\psi\in L^{1}$,
(2.1) $|\operatorname{Cov}(\varphi,\psi\circ T^{n})|\lesssim_{\alpha}\|\varphi\|_{\alpha}\|\psi\|_{L^{1}}e^{-\tau_{0}n}.$
For invertible systems, the transfer operator is given by $Pf=f\circ T^{-1}$ and as such one cannot hope for Condition D.1 to hold. In such cases, we will assume the following decorrelation property.
Condition D.2 (Decay of Correlations).
For every $\alpha\in(0,1]$, there exists $\tau_{0}>0$ such that, for all $n\in\mathbb{N}$ and all $\varphi,\psi\in\mathcal{H}_{\alpha}$,
$|\operatorname{Cov}(\varphi,\psi\circ T^{n})|\lesssim_{\alpha}\|\varphi\|_{\alpha}\|\psi\|_{\alpha}e^{-\tau_{0}n}.$
For the proof of the CLTs in particular, we assume the following multiple decorrelation property.
Condition D.3 (Multiple Decorrelation).
For every $\alpha\in(0,1]$, there exists $r\geq 1$, a polynomial function $P_{r}$ with real non-negative coefficients and a real number $\tau_{r}>0$ such that, for all integers $m,m^{\prime}$, for all functions $\varphi_{1},\ldots,\varphi_{m+m^{\prime}}\in\mathcal{H}_{\alpha}$, for all non-negative sequence of integers $k_{1}\leq\ldots\leq k_{m}$ and $l_{1}\leq\ldots\leq l_{m^{\prime}}$, and for all non-negative integers $N$, we have
$\operatorname{Cov}\biggl(\prod_{i=1}^{m}\varphi_{i}\circ T^{k_{i}},\prod_{j=1}^{m^{\prime}}\varphi_{m+j}\circ T^{N+l_{j}}\biggr)\\ \leq\biggl(\prod_{i=1}^{m+m^{\prime}}\|\varphi_{i}\|_{\infty}+\sum_{i=1}^{m+m^{\prime}}\lvert\varphi_{i}\rvert_{\alpha}\prod_{j\neq i}\|\varphi_{j}\|_{\infty}\biggr)P_{r}(l_{m^{\prime}})e^{-\tau_{r}(N-rk_{m})}.$
In addition to good mixing properties, we require certain continuity properties of the measure, primarily for the purpose of approximating indicator functions of balls by Hölder continuous functions.
Condition M.1 (Frostman property).
There exists $s_{0}>0$ such that, for all $x\in X$ and $r>0$,
$\mu\bigl(B(x,r)\bigr)\lesssim r^{s_{0}}.$
Condition M.2 (Thin annuli).
There exists $\alpha_{0},\rho_{0}>0$ such that, for all $x\in X$ and $\varepsilon\leq r\leq\rho_{0}$,
$\mu\bigl(B(x,r+\varepsilon)\setminus B(x,r)\bigr)\lesssim\varepsilon^{\alpha_{0}}.$
The conditions on the measure imply that, when defined implicitly, $r_{n}$ is well defined for each $n\in\mathbb{N}$ and that $r_{n}$ Lipschitz continuous on $\operatorname{supp}\mu$ with $\lvert r_{n}\rvert_{\textup{Lip}}=1$. In the case that $M_{n}\to 0$, then $r_{n}\to 0$ uniformly on $\operatorname{supp}\mu$. (The proofs of these facts are straightforward [35, Lemmas 3.1, 3.2].) When $r_{n}\to 0$ uniformly on $\operatorname{supp}\mu$, we may therefore without loss of generality assume that $r_{n}\leq\rho_{0}$ $\mu$-a.e. for all $n\in\mathbb{N}$.
We conclude this section with some remarks on the conditions.
Remark 1 (Decay of Correlations).
For non-invertible systems, one can sometimes show that the transfer operator has a spectral gap, which implies Condition D.1. Systems for which the transfer operator has a spectral gap include sub-shifts of finite type (see e.g. Bowen [4]), and piecewise uniformly expanding maps on $[0,1]$ equipped with an absolutely continuous invariant measure (see [26, 36]).
A weaker decorrelation property is Condition D.2, which, in addition to the aforementioned systems, holds for piecewise uniformly expanding maps on $[0,1]$ equipped with a Gibbs measure (see [28]), and Axiom A diffeomorphisms equipped with a Gibbs measure corresponding to a Hölder continuous observable (see Bowen [4]). It is a standard assumption when considering statistical properties of dynamical systems.
Condition D.3 lies in between Conditions D.1 and D.2 in terms of strength, and is a version of a multiple decorrelation property. The version we use comes from Pène [32], in which it is referred to as ‘Property $(\mathcal{P}_{r})$’ for a real number $r\geq 1$. For dynamical systems to which Young’s method of [39] can be applied, it can be shown that Condition D.3 is satisfied for any $r>1$ (see Corollary B.2. of [31]). This includes Axiom A diffeomorphisms equipped with the SRB measure. Conze and Le Borgne [9] assume a stronger multiple decorrelation property (Property 3.1 as it appears in [9]), namely they assume Condition D.3 holds with $r=1$ and $P_{r}=C$ for some constant $C>0$. This version holds for billiards with bounded horizon and no corners [7, Theorem 7.41] for ‘dynamically Hölder continuous observables’.
We note that in the proof of Theorem 2.1, it is important that the multiple decorrelation estimates are uniform in the number of observables, i.e., that $P_{r}$ and $\tau_{r}$ do not depend on $m,m^{\prime}$. Versions of a multiple decorrelation of any order, but for which the multiplicative constant depends on the number of observables are proven in the following works: Dolgopyat [11] for hyperbolic systems for partially hyperbolic systems that are ‘strongly $u$-transitive with exponential rate’; Kotani–Sunada [25] for Axiom A diffeomorphisms equipped with a Gibbs measure corresponding to a Hölder continuous observable; and Björklund–Einsiedler–Gorodnik [2] for semi-simple Lie groups, semi-simple $S$-algebraic groups, and semi-simple adele groups.
Remark 2 (Continuity of Measure).
Conditions M.1 and M.2 are standard assumptions; they are assumed, either implicitly or explicitly, in the aforementioned works establishing RSBCs [15, 20, 27, 33, 34, 35]. If $\mu$ is absolutely continuous with density $h\in L^{q}(m)$ for some $q>1$, then both conditions hold. Furthermore, if $X$ is a Riemannian manifold with dimension an $\dim X$ and $\mu$ satisfies Condition M.1 for $s>\dim X-1$, then Condition M.2 holds for $\alpha_{0}=\dim X+1-s$ (see the proof of Theorem 1.1 in [35]).
2.2. Distributional Limit Laws for Recurrence
For the first theorem, we consider the implicit radii problem and obtain (CLT). Recall that $(X,d)$ is a smooth compact Riemannian manifold and that the sets $E_{n}$ are defined in terms of a reference sequence $(M_{n})$. Let $\sigma_{n}^{2}=\operatorname{Var}(\sum_{k=1}^{n}\mathbf{1}_{E_{k}})$ and $s_{n}^{2}(y)=\operatorname{Var}(\sum_{k=1}^{n}\mathbf{1}_{B(y,r_{k}(y))}\circ T^{k})$.
Theorem 2.1.
Suppose that $(X,T,\mu,d)$ is an m.m.p.s. satisfying Conditions D.3, M.1 and M.2 and suppose that the sequence $(M_{n})$ satisfies Condition $\mathcal{S}_{\gamma,\upsilon}$ for some $\gamma\in(0,1)$ and $\upsilon>1$. Assume that
(2.2) $\frac{1}{s_{n}(y)}\sum_{k=1}^{n}\bigl(\mathbf{1}_{B(y,r_{k}(y))}\circ T^{k}-M_{k}\bigr)\xrightarrow[]{\mathcal{D}}N(0,1)\quad\text{$\mu$-a.e.\ $y\in X$}$
and
(2.3) $\frac{s_{n}^{2}}{\sigma_{n}^{2}}\xrightarrow[]{\mathcal{D}}1.$
Then
$\frac{1}{\sigma_{n}}\sum_{k=1}^{n}\bigl(\mathbf{1}_{B(x,r_{k}(x))}(T^{k}x)-\mu(E_{k})\bigr)$
converges in distribution to a standard Gaussian variable.
Remark 3.
We can replace the condition of multiple decorrelation for Hölder continuous observables with decorrelation for $\mathrm{BV}$ versus $L^{1}$ observables. The proof is simpler in that case.
Remark 4.
The condition in (2.3) that the ratio of variances converge in distribution can be verified if the system satisfies a short returns estimate of the following form: there exist $\delta_{0}>0$ and $\upsilon_{0}>1$ such that
$\mu(E_{k}\cap E_{j})\lesssim\mu(E_{k})^{1+\delta_{0}}.$
This will be proven in Lemma 3.4. The short returns estimate will be discussed further in Section 9.
Theorem 2.1 can be applied to interval maps equipped with a Gibbs measure, and Axiom A systems equipped with SRB measures satisfying (M.1) for $s>\dim X-1$.
We also establish a distributional limit in for the explicit radii problem. Recall that $\hat{E}_{n}=\{x\in X:T^{n}x\in B(x,r_{n})\}$ and $\hat{M}_{n}=\int\mu(B(x,r_{n}))\mathop{}\!d\mu(x)$. Let $\hat{\sigma}_{n}^{2}=\operatorname{Var}(\sum_{k=1}^{n}\mathbf{1}_{\hat{E}_{k}})$ and $\hat{s}_{n}^{2}=\operatorname{Var}(\sum_{k=1}^{n}\mathbf{1}_{B(y,r_{k})}\circ T^{k})$.
Theorem 2.2.
Suppose that $(X,T,\mu,d)$ is an m.m.p.s. satisfying Conditions D.3, M.1 and M.2. Suppose further that $r_{n}\to 0$ and that $(\hat{M}_{n})$ satisfies Condition $\mathcal{S}_{\gamma,\upsilon}$ for some $\gamma\in(0,1)$ and $\upsilon>1$. Assume that
$\frac{1}{\hat{s}_{n}(y)}\sum_{k=1}^{n}\bigl(\mathbf{1}_{B(y,r_{k})}\circ T^{k}-\mu(B(y,r_{k}))\bigr)\xrightarrow[]{\mathcal{D}}N(0,1)\quad\text{$\mu$-a.e.\ $y\in X$}$
and
(2.4) $\lim_{n\to\infty}\int_{X}\Bigl|\frac{\hat{s}_{n}}{\hat{\sigma}_{n}}-g\Bigr|\mathop{}\!d\mu=0$
for some $g\in L^{1}(\mu)$. Then the characteristic function of
(2.5) $x\mapsto\frac{1}{\hat{\sigma}_{n}}\sum_{k=1}^{n}\bigl(\mathbf{1}_{B(x,r_{k})}(T^{k}x)-\mu(B(x,r_{k}))\bigr)$
converges everywhere to
$t\mapsto\int_{X}e^{-\frac{1}{2}g(x)^{2}t^{2}}\mathop{}\!d\mu(x).$
Remark 5.
In the setting of [15], He obtains that
$\lim_{n\to\infty}\frac{\sum_{k=1}^{n}\mathbf{1}_{B(x,r_{k})}(T^{k}x)}{\sum_{k=1}^{n}\mu(B(x,r_{k}))}=1\quad\text{$\mu$-a.e.\ $x$},$
which justifies the form of the sum in (2.5).
For general measures, (2.4) may be difficult to verify or may fail completely. Similarly, verifying that $(\hat{M}_{n})$ satisfies Condition $\mathcal{S}_{\gamma,\upsilon}$ is in general difficult. This is in part because the measure of the sets $B(x,r_{k})$ may vary wildly as $x$ varies. However, if one assumes that the measure $\mu$ is absolutely continuous, then studying the asymptotic behaviour of $\hat{s}_{n}/\hat{\sigma}_{n}$ becomes tractable. In particular, under the additional assumption of a short returns estimate, the asymptotic relationship between $\hat{s}_{n}$ and $\hat{\sigma}_{n}$ is clear (see Lemma 3.6) and one obtains the following corollary to Theorem 2.2.
Corollary 2.1.
Suppose that $(X,T,\mu,d)$ is an m.m.p.s. satisfying Condition D.3 and that $\mathop{}\!d\mu=h\mathop{}\!dm$ with density $h\in L^{2}(m)$ such that $h\geq c>0$ for some constant $c>0$. Assume that
$\frac{1}{\hat{s}_{n}(y)}\sum_{k=1}^{n}\bigl(\mathbf{1}_{B(y,r_{k})}\circ T^{k}-\mu(B(y,r_{k}))\bigr)\xrightarrow[]{\mathcal{D}}N(0,1)\quad\text{$\mu$-a.e.\ $y\in X$}.$
and that there exist constants $\delta_{0}>0$ and $p_{0}>1$ such that
(2.6) $\int_{X}\mu\bigl(B(y,r_{k})\cap T^{-l}B(y,r_{k})\bigr)\mathop{}\!d\mu(y)\lesssim\hat{M}_{k}^{1+\delta_{0}}$
for all sufficiently large $k$ and all $l=1,\ldots,(\log k)^{p_{0}}$. If $(r_{n})$ satisfies Condition $\mathcal{S}_{\gamma,\upsilon}$ for some $\gamma\in(0,d)$ and $\upsilon>2(\delta_{0}\dim X)^{-1}$, then the characteristic function of
$x\mapsto\frac{1}{\hat{\sigma}_{n}}\sum_{k=1}^{n}\bigl(\mathbf{1}_{B(x,r_{k})}(T^{k}x)-\mu(B(x,r_{k}))\bigr)$
converges everywhere to
$t\mapsto\int_{X}\exp\Bigl(-\frac{h(x)}{2\mu(h)}t^{2}\Bigr)\mathop{}\!d\mu(x).$
Remark 6.
Although the result holds for $h\in L^{2}(m)$ given that (2.6) holds, we can only verify (2.6) when $h\in L^{\infty}(m)$. See §9 for further details. Since $h$ is bounded away from 0, the limiting distribution has probability density function $f\colon\mathbb{R}\to(0,\infty)$ given by
$f(t)=\frac{1}{\sqrt{2\pi}}\int_{X}\frac{1}{\sqrt{h(y)}}\exp\Bigl(-\frac{t^{2}}{2h(y)}\Bigr)\mathop{}\!d\mu(y)$
by the inversion formula.
Remark 7.
Comparing with the result of Holland and Todd [20], namely (1.4), we see that both our probability density functions involve averaging over the ‘expected’ distributions. More precisely, the density in (1.4) is an average of Poisson distributions with variance $\tau h(x)$. In our case, our density is an average of Gaussian distributions with variance $h(x)$.
Comparing with the result of Conze and Le Borgne [9, Theorem 3.3], we recover their result when $X$ is a compact group and $r_{k}=r\in(0,\infty)$ for all $k$.
2.3. Almost Sure Invariance Principles for Non-invertible Systems
In establishing distributional laws for recurrence, we assume that the CLT holds for restricted versions of the shrinking target problem. Our results in this section establish that these assumptions are met for the systems under our consideration by establishing a stronger property than the CLT. We establish more general results for non-autonomous Hölder observables, which we then apply to the shrinking target problem. The first result is an ASIP for a non-stationary sequence of observables that are uniformly bounded in $\|\cdot\|_{\infty}$ but polynomially growing in $\|\cdot\|_{\alpha}$.
Theorem 2.3.
Let $(X,T,\mu,d)$ be an m.m.p.s. satisfying Condition D.1 and suppose $(\varphi_{n})$ is a sequence of $\alpha$-Hölder continuous functions satisfying $\sup_{n\geq 1}\|\varphi_{n}\|_{\infty}<\infty$ and $\|\varphi_{n}\|_{\alpha}\lesssim n^{p}$ for some constant $p>0$. Let $M_{n}=\mu(\varphi_{n})$ and $\sigma_{n}^{2}=\operatorname{Var}(\sum_{k=1}^{n}\varphi_{k}\circ T^{k})$. Suppose that $(M_{n})$ satisfies Condition $\mathcal{S}_{\gamma,\upsilon}$ for some $\gamma\in(0,1)$ and $\upsilon>1$, and that $\mu(|\varphi_{n}|)\lesssim M_{n}$. Then $(\tilde{\varphi}_{n}\circ T^{n})$ satisfies (ASIP) for any $\beta<1/2$ with $\sum_{k=1}^{n}\mathbf{E}[Z_{k}^{2}]=\sigma_{n}^{2}+\mathcal{O}(\sigma_{n}\log\sigma_{n})$.
As mentioned in the introduction, Haydn, Nicol, Vaienti and Zhang [13] prove an ASIP (and therefore a CLT) for the shrinking target problem when the transfer operator of the system contracts in the $\mathrm{BV}$ norm, i.e., under Condition D.1 with $\|\cdot\|_{\alpha}$ replaced with $\|\cdot\|_{\textup{BV}}$. Thus, their result can be used to establish the conclusion of Theorem 2.1 for such systems; however, in order to prove the conclusion for a larger class of systems – including uniformly hyperbolic systems – one needs to weaken their assumptions. In [13, Theorem 6.1], the authors assume, for their sequence of observables $(\varphi_{n})$, that $\sup_{n\geq 1}\|\varphi_{n}\|_{\alpha}<\infty$. Therefore, they are unable to apply their result to the shrinking target problem. Theorem 2.3 relaxes this condition, which enables us to use an approximation argument and establish an ASIP for the shrinking target problem. Additionally, we improve [13, Theorem 6.1] by weakening their condition that $\sigma_{n}^{2}\geq n^{\delta}$ for some $\delta>\frac{\sqrt{17}-1}{4}\approx 0.781$ to $\sigma_{n}^{2}\geq n^{\delta}$ for any $\delta>0$. (We will see that $\sigma_{n}^{2}\gtrsim n^{1-\gamma}$.)
Corollary 2.2.
Let $(X,T,\mu,d)$ be a m.m.p.s. satisfying Conditions D.1, M.1 and M.2. Let $B_{n}=B(y_{n},r_{n})$ be a sequence of balls with centres $y_{n}\in\operatorname{supp}\mu$ and write $M_{n}=\mu(B_{n})$ and $\sigma_{n}^{2}=\operatorname{Var}(\sum_{k=1}^{n}\mathbf{1}_{B_{n}}\circ T^{n})$. Assume that the sequence $(M_{n})$ satisfies Condition $\mathcal{S}_{\gamma,\upsilon}$ for some $\gamma\in(0,1)$ and $\upsilon>1$. Then $(\tilde{\mathbf{1}}_{B_{n}}\circ T^{n})$ satisfies (ASIP) for any $\beta<1/2$ with $\sum_{k=1}^{n}\mathbf{E}[Z_{k}^{2}]=\sigma_{n}^{2}+\mathcal{O}(\sigma_{n}\log\sigma_{n})$.
In extending Theorem 2.3 to Axiom A systems, one cannot directly apply the transfer operator method because the system is invertible. Instead, the problem of establishing an ASIP for Axiom A systems can be reduced to the non-invertible setting via a non-autonomous version of the Sinai trick (see Lemma 7.2). However, the resulting sequence of observables on the non-invertible space is not uniformly bounded in $\|\cdot\|_{\infty}$. Consequently, it is necessary to relax the uniform bound assumption in Theorem 2.3, at the cost of restricting the lower growth rate of $\sigma_{n}^{2}$.
Theorem 2.4.
Let $(X,T,\mu,d)$ be a m.m.p.s. satisfying Condition D.1 and suppose that $(\varphi_{n})$ is a sequence of $\alpha$-Hölder continuous observables such that $\|\varphi_{n}\|_{\infty}\lesssim(\log n)^{q}$ and $\|\varphi_{n}\|_{\alpha}\lesssim n^{p}$ for some constants $p,q>0$. Suppose also that $\sigma_{n}^{2}:=\operatorname{Var}(\sum_{k=1}^{n}\varphi_{k}\circ T^{k})\gtrsim n^{\delta}$ for some $\delta>1/2$. Then $(\tilde{\varphi}_{n}\circ T^{n})$ satisfies (ASIP) for any $\beta<\min\{\frac{1}{2},1-\frac{1}{2\delta}\}$ with $\sum_{k=1}^{n}\mathbf{E}[Z_{k}^{2}]=\sigma_{n}^{2}+\mathcal{O}(\sigma_{n}(\log\sigma_{n})^{1+q})$.
2.4. Almost Sure Invariance Principles for Axiom A Diffeomorphisms
For the following results, we assume that $(X,T,\mu,d)$ is an Axiom A system restricted to a mixing component, where $\mu$ is a Gibbs measure corresponding to a Hölder continuous observable. We first state a non-stationary ASIP for a sequence of Hölder continuous observables whose norms are polynomially growing.
Theorem 2.5.
Suppose that $T\colon X\to X$ is an Axiom A diffeomorphism restricted to a mixing component and that $\mu$ is a Gibbs measure corresponding to a Hölder continuous observable. Let $(\varphi_{n})$ be a sequence of $\alpha$-Hölder continuous functions such that $\|\varphi_{n}\|_{\infty}\lesssim(\log n)^{q}$ and $\|\varphi_{n}\|_{\alpha}\lesssim n^{p}$ for some constants $p,q>0$. Assume that $\sigma_{n}^{2}:=\operatorname{Var}(\sum_{k=1}^{n}\varphi_{k}\circ T^{k})\geq n^{\delta}$ for some $\delta>1/2$. Then $(\tilde{\varphi}_{n}\circ T^{n})$ satisfies (ASIP) for any $\beta<\min\{\frac{1}{2},1-\frac{1}{2\delta}\}$ with $\sum_{k=1}^{n}\mathbf{E}[Z_{k}^{2}]=\sigma_{n}^{2}+\mathcal{O}(\sigma_{n}(\log\sigma_{n})^{2+q})$.
Remark 8.
We compare Theorem 2.5 to other known results. The improvement over [13, Corollary 6.2] is two-fold. First, we allow the sequence $(\|\varphi_{n}\|_{\alpha})_{n}$ of Hölder norms to grow at a polynomial rate, and second, we relax the condition on $\delta$ from $\delta>\frac{\sqrt{17}-1}{4}\approx 0.781$ to $\delta>1/2$.
Compare also Theorem 2.5 to the work of Chen, Yang and Zhang [5] who consider 2-dimensional hyperbolic systems with singularities, but with $\mu$ being the SRB measure. They also require that $\sigma_{n}^{2}>n^{\delta}$ for some $\delta>1/2$.
Finally, we obtain the ASIP for the shrinking target problem in the case of Axiom A systems.
Corollary 2.3.
Suppose that $T\colon X\to X$ is an Axiom A diffeomorphism restricted to a mixing component and that $\mu$ is a Gibbs measure corresponding to a Hölder continuous observable. Suppose that $\mu$ satisfies Conditions M.1 and M.2. Let $(B_{n})$ be a sequence of balls with centres in $\operatorname{supp}\mu$ and let $M_{n}=\mu(B_{n})$ and $\sigma_{n}^{2}=\operatorname{Var}(\sum_{k=1}^{n}\mathbf{1}_{B_{k}}\circ T^{k})$. If $(M_{n})$ satisfies Condition $\mathcal{S}_{\gamma,\upsilon}$ for some $\gamma<1/2$ and $\upsilon>1$, then $(\tilde{\mathbf{1}}_{B_{n}}\circ T^{n})$ satisfies (ASIP) for any $\beta<1-\frac{1}{2(1-\gamma)}$ with $\sum_{k=1}^{n}\mathbf{E}[Z_{k}^{2}]=\sigma_{n}^{2}+\mathcal{O}(\sigma_{n}(\log\sigma_{n})^{2})$.
3. Variance Estimates
We devote this section to asymptotic estimates of the variances appearing in the previous sections. In particular, we establish their lower growth rates and the asymptotic relationship between them. These estimates will be used in the proofs of the main results. As in the previous section, we assume that $(X,d)$ is a smooth compact Riemannian manifold. We let $\mathcal{M}_{n}=\sum_{k=1}^{n}M_{k}$ for a sequence of scalars $(M_{n})$. We begin with a pointwise variance estimate.
Lemma 3.1.
Suppose that $(X,T,\mu,d)$ is an m.m.p.s. satisfying Conditions D.2, M.1 and M.2. Then the following hold.
  1. (1)
    If $B_{n}=B(y_{n},r_{n})$ is a sequence of balls with centres $y\in\operatorname{supp}\mu$ such that the sequence $M_{n}=\mu(B_{n})$ satisfies Condition $\mathcal{S}_{\gamma,\upsilon}$ for some $\gamma\in(0,1)$ and $\upsilon>1$, then
    $\liminf_{n\to\infty}\frac{\sigma_{n}^{2}}{\mathcal{M}_{n}}\geq 1,$
    where $\sigma_{n}^{2}=\operatorname{Var}(\sum_{k=1}^{n}\mathbf{1}_{B_{k}}\circ T^{k})$. As a consequence, $\sigma_{n}^{2}\gtrsim n^{1-\gamma}$.
  2. (2)
    If $(M_{n})$ is a sequence of satisfying Condition $\mathcal{S}_{\gamma,\upsilon}$ for some $\gamma\in(0,1)$ and $\upsilon>1$ and $B(y,r_{n}(y))$ is a sequence of balls such that $\mu(B(x,r_{n}(x)))=M_{n}$ for all $x\in X$ and $n\in\mathbb{N}$, then
    $\liminf_{n\to\infty}\frac{s_{n}^{2}}{\mathcal{M}_{n}}\geq 1\quad\text{uniformly on $\operatorname{supp}\mu$},$
    where $s_{n}^{2}(y)=\operatorname{Var}(\sum_{k=1}^{n}\mathbf{1}_{B(y,r_{k}(y))}\circ T^{k})$. As a result, $s_{n}^{2}(y)\gtrsim n^{1-\gamma}$ for all $y\in\operatorname{supp}\mu$.
3.1. Variance estimates for implicit radii
The following result concerns the implicit radii problem. In the following results, we assume that $(X,T,\mu,d)$ is an m.m.p.s. satisfying Conditions M.1, M.2 and Conditions D.3 restricted to $m+m^{\prime}=3$. We also assume that $(M_{n})$ is a sequence satisfying Condition $\mathcal{S}_{\gamma,\upsilon}$ for some $\gamma\in(0,1)$ and $\upsilon>1$, which will be restricted if necessary. Recall that we consider the recurrence sets $E_{n}=\{x\in X:T^{n}x\in B(x,r_{n}(x))\}$, where $\mu(B(x,r_{n}(x)))=M_{n}$ for all $x\in X$ and $n\in\mathbb{N}$. We let $\mathcal{E}_{n}=\sum_{k=1}^{n}\mu(E_{k})$, $\sigma_{n}^{2}=\operatorname{Var}(\sum_{k=1}^{n}\mathbf{1}_{E_{k}})$ and $s_{n}^{2}(y)=\operatorname{Var}(\sum_{k=1}^{n}\mathbf{1}_{B(y,r_{k}(y))}\circ T^{k})$. We begin with a lower growth result for $\sigma_{n}^{2}$.
Lemma 3.2.
Given the above assumptions, we have that
$\liminf_{n\to\infty}\frac{\sigma_{n}^{2}}{\mathcal{E}_{n}}\geq 1.$
In particular, $\sigma_{n}^{2}\gtrsim n^{1-\gamma}$.
We also have that $\sigma_{n}^{2}$ is asymptotically equal to the average of $s_{n}^{2}$.
Lemma 3.3.
Given the above assumptions, we have that
$\lim_{n\to\infty}\int_{X}\frac{s_{n}^{2}(y)}{\sigma_{n}^{2}}\mathop{}\!d\mu(y)=1.$
Under a short returns estimate, we obtain convergence in measure of the quantity $s_{n}^{2}/\sigma_{n}^{2}$.
Lemma 3.4.
In addition to the above assumptions, suppose that, for all $D>0$, there exists $\delta_{0}>0$ such that
(3.1) $\sum_{k=1}^{n}\sum_{j=k+1}^{k+D\log k}\int_{X}\mu\bigl(B(y,r_{k}(y))\cap T^{-(j-k)}B(y,r_{k}(y))\bigr)\mathop{}\!d\mu(y)=o(\sigma_{n}^{2-\delta_{0}}).$
Then, $s_{n}^{2}/\sigma_{n}^{2}\xrightarrow[]{\mu}1$, i.e., for all $\varepsilon>0$,
$\lim_{n\to\infty}\mu\biggl\{y:\Bigl|\frac{s_{n}^{2}(y)}{\sigma_{n}^{2}}-1\Bigr|>\varepsilon\biggr\}=0.$
3.2. Variance estimates for explicit radii
For the explicit radii problem, we obtain analogous results. Recall that we let $(r_{n})$ be a sequence in $(0,\infty)$ and define $\hat{M}_{n}=\int\mu(B(x,r_{n}))\mathop{}\!d\mu$ We consider the recurrence sets $\hat{E}_{n}=\{x\in X:T^{n}x\in B(x,r_{n})\}$, and let $\hat{\sigma}_{n}^{2}=\operatorname{Var}(\sum_{k=1}^{n}\mathbf{1}_{\hat{E}_{k}}-\mu(B(x,r_{k})))$ and $\hat{s}_{n}^{2}(y)=\operatorname{Var}(\sum_{k=1}^{n}\mathbf{1}_{B(y,r_{k})}\circ T^{k})$. We also let $\hat{\mathcal{E}}_{n}=\sum_{k=1}^{n}\mu(\hat{E}_{k})$ and $\hat{\mathcal{M}}_{n}=\sum_{k=1}^{n}\hat{M}_{k}$. As before, we assume that the m.m.p.s. $(X,T,\mu,d)$ satisfies Conditions M.1, M.2 and D.3 restricted to $m+m^{\prime}=3$.
Lemma 3.5.
Given the above assumptions on the system $(X,T,\mu)$, suppose that $r_{n}\lesssim(\log n)^{-\upsilon/s_{0}}$ for some $\upsilon>1$ and that $\hat{M}_{n}\gtrsim n^{-\gamma}$ for some $\gamma\in(0,1)$. Then
$\liminf_{n\to\infty}\frac{\hat{\sigma}_{n}^{2}}{\hat{\mathcal{E}}_{n}}\geq 1.$
In particular, $\hat{\sigma}_{n}^{2}\gtrsim n^{1-\gamma}$.
Remark 9.
If $\mu$ is absolutely continuous with density $h\in L^{p}(m)$ for some $p>1$ and if $h$ is bounded away from 0, then $r^{\dim X}\lesssim\hat{M}_{n}\lesssim r^{\dim X/q}$, where q=p/(p1)q=p/(p-1). We can therefore express the assumptions on $\hat{M}_{n}$ in terms of $r_{n}$, which is more natural in this setting.
In the case when $\mu$ is absolutely continuous and under a short returns estimate, we obtain convergence in $L^{1}$ of the quantity $\hat{s}_{n}^{2}/\hat{\sigma}_{n}^{2}$. Recall that $m$ denotes the volume measure on $X$.
Lemma 3.6.
In addition to the above assumptions, suppose that $\mathop{}\!d\mu=h\mathop{}\!dm$ with density $h\in L^{2}(m)$ such that $h\geq c>0$ for some constant $c>0$. Additionally, suppose that the following short returns estimate holds: there exist constants $\delta_{0}>0$ and $p_{0}>1$ such that
(3.2) $\int_{X}\mu\bigl(B(y,r_{k})\cap T^{-l}B(y,r_{k})\bigr)\mathop{}\!d\mu(y)\lesssim\hat{M}_{k}^{1+\delta_{0}}.$
for all sufficiently large $k$ and all $l=1,\ldots,(\log k)^{p_{0}}$. If $(r_{n})$ satisfies Condition $\mathcal{S}_{\gamma,\upsilon}$ for some $\gamma\in(0,d)$ and $\upsilon>2(\delta_{0}\dim X)^{-1}$, then
$\lim_{n\to\infty}\int_{X}\biggl|\frac{\hat{s}_{n}}{\hat{\sigma}_{n}}-\frac{\sqrt{h}}{\sqrt{\mu(h)}}\biggr|\mathop{}\!d\mu(y)=0.$
As mentioned after Corollary 2.1, we can only verify (3.2) when $h\in L^{\infty}(m)$. See §9 for further details.
4. Supporting Results
The method for establishing (ASIP) for the shrinking target problem has its roots in the following result by Cuny and Merlevède [10, Theorem 2.3] that establishes (ASIP) for a sequence of random variables $(X_{n})$. We incorporate [10, Remark 2.4] into the statement, which weakens the conditions on boundedness of the sequence $(\mathbf{E}X_{n}^{2})_{n}$. This is the method the authors in [13] employ to establish (ASIP) for the shrinking target problem in the case of non-invertible systems satisfying Condition D.1 with $\|\cdot\|_{\alpha}$ replaced with $\|\cdot\|_{\textup{BV}}$ [13, Theorem 5.1].
Theorem 4.1 ([10, Theorem 2.3]).
Let $(X_{n})$ be a sequence of square-integrable random variables adapted to a non-increasing filtration $(\mathcal{G}_{n})_{n}$. Assume that $\mathbf{E}[X_{n}|\mathcal{G}_{n+1}]=0$ a.s., that $\sigma_{n}^{2}:=\sum_{k=1}^{n}\mathbf{E}[X_{k}^{2}]\to\infty$ and that there exists $s\in[0,1)$ such that $\mathbf{E}[X_{n}^{2}]=\mathcal{O}(\sigma_{n}^{2s})$ for all $n$. Let $(a_{n})_{n}$ be a non-decreasing sequence of positive numbers such that $(a_{n}/\sigma_{n}^{2})_{n}$ is non-increasing and $(a_{n}/\sigma_{n})_{n}$ is non-decreasing. Assume that
(A) $\sum_{k=1}^{n}(\mathbf{E}[X_{k}^{2}|\mathcal{G}_{k+1}]-\mathbf{E}[X_{k}^{2}])=o(a_{n})\quad\text{$\mathbf{P}$-a.s.,}$
(B) $\sum_{n=1}^{\infty}a_{n}^{-v}\mathbf{E}[|X_{n}|^{2v}]<\infty\quad\text{for some $1\leq v\leq 2$}.$
Then, enlarging our probability space if necessary, it is possible to find a sequence $(Z_{k})_{k\geq 1}$ of independent centred Gaussian variables with $\mathbf{E}[Z_{k}^{2}]=\mathbf{E}[X_{k}^{2}]$ such that
$\sup_{1\leq k\leq n}\biggl|\sum_{i=1}^{k}X_{i}-\sum_{i=1}^{k}Z_{i}\biggr|=o\bigl((a_{n}(\log(\sigma_{n}^{2}/a_{n})+\log\log a_{n}))^{1/2}\bigr)\quad\text{$\mathbf{P}$-a.s.}$
The following result is a standard tool for establishing strong Borel–Cantelli lemmas – a non-autonomous version of a strong law of large numbers – by reducing the problem to a second moment analysis. Here it will be used to establish Condition A in Theorem 4.1. We state the result as formulated by Harman [12].
Lemma 4.1 (Gàl–Koksma).
Let $\mu$ be a probability measure on the space $X$. Let $(f_{k})$ be a sequence of non-negative $\mu$-measurable functions, and let $(g_{k})$ and $(\gamma_{k})$ be sequences of real numbers such that
$0\leq g_{k}\leq\gamma_{k}\quad k\in\mathbb{N}.$
Write $\Gamma(N)=\sum_{k=1}^{N}\gamma_{k}$ and suppose that $\Gamma(N)\to\infty$ as $N\to\infty$. Suppose that, for arbitrary integers $1\leq m\leq n$,
$\int_{X}\biggl(\sum_{k=m}^{n}(f_{k}(x)-g_{k})\biggr)^{2}\mathop{}\!d\mu(x)\lesssim\sum_{k=m}^{n}\gamma_{k}.$
Then, for any given $\varepsilon>0$ and for $\mu$-a.e. $x$, we have
$\biggl|\sum_{k=1}^{N}f_{k}(x)-\sum_{k=1}^{N}g_{k}\biggr|\lesssim_{x,\varepsilon}\Gamma(N)^{1/2}\bigl(\log\Gamma(N)\bigr)^{3/2+\varepsilon}+\max_{1\leq k\leq N}g_{k}.$
For the shrinking target problem, the hitting sets are of the form $A_{n}=\{x\in X:T^{n}x\in B(y_{n},r_{n})\}$. Since $T$ preserves $\mu$, one obtains that $\mu(A_{n})=\mu(B(y_{n},r_{n}))$. In the recurrence setting, the targets depend on the initial point and one cannot directly use the measure-preserving property. Nevertheless, if the system is rapidly mixing, then one can obtain estimates for the difference between $\mu(E_{n})$ and $\mu(B(x,r_{n}(x)))$, and $\mu(\hat{E}_{n})$ and $\hat{M}_{n}=\int\mu(B(x,r_{n}))\mathop{}\!d\mu(x)$. The following lemma was established by Kirsebom, Kunde and Persson [23, Lemma 4.1] for interval systems. In [35, Proposition 4.1], the author establishes (4.1) for uniformly hyperbolic systems. A simple rewriting of the proof of [35, Proposition 4.1], gives (4.2) as well. Note that Kirsebom, Kunde and Persson’s result as it appears in [23] is slightly stronger when restricting to their setting.
Lemma 4.2.
Suppose that Conditions D.2, M.1 and M.2 hold and that the sequences $(M_{n})$ and $(\hat{M}_{n})$ converge to 0. Then there exists $\eta_{0}>0$ such that, for all $n\in\mathbb{N}$,
(4.1) $|\mu(E_{n})-M_{n}|\lesssim e^{-\eta_{0}n}.$
and
(4.2) $\Bigl|\mu(\hat{E}_{n})-\int_{X}\mu(B(x,r_{n}))\mathop{}\!d\mu(x)\Bigr|\lesssim e^{-\eta_{0}n}.$
In order to leverage Conditions D.2 or D.3, which apply to Hölder continuous observables, we would like to obtain a partition of $X$ for which the indicator function of each set can be well approximated by Hölder continuous observables. This turns out to be difficult in general; however, by removing a subset of $X$ with small measure, Conze and Le Borgne obtain a partition of the remaining space with the desired properties. The statement and proof was originally communicated in private correspondence, but can be found in [34, Lemma 4].
Lemma 4.3 (Conze–Le Borgne).
Suppose that $X$ is a compact $d$-dimensional smooth manifold and let $\kappa_{0}>0$. Then, for all $n\in\mathbb{N}$, there exists a subset $Y_{n}\subset X$ and a partition $\mathcal{Q}_{n}$ of $X\setminus Y_{n}$ consisting entirely of sets with positive measure such that
$\#\mathcal{Q}_{n}$ $\leq n^{d\kappa_{0}},$
$|\mathcal{Q}_{n}|$ $\lesssim n^{-\kappa_{0}},$
$\mu(Y_{n})$ $\leq n^{-\kappa_{0}}.$
Moreover, for each $Q\in\mathcal{Q}_{n}$, there exists a density function $\rho_{n,Q}\colon X\to[0,\infty)$ such that the estimates
$\|\rho_{Q,n}\|_{\infty}$ $\lesssim n^{(d+1)\kappa_{0}},$
$\lvert\rho_{Q,n}\rvert_{\textup{Lip}}$ $\lesssim n^{2(d+2)\kappa_{0}},$
$\|\rho_{Q,n}-\mu(Q)^{-1}\mathbf{1}_{Q}\|_{L^{1}}$ $\lesssim n^{-\kappa_{0}},$
hold uniformly in $n$.
We call the set $\mathcal{Q}_{n}$ a quasi-partition of $X$ if it satisfies the conclusion of Lemma 4.3.
We state a lemma regarding a Hölder continuous approximation of indicator functions of balls as it will be used in many of the proofs. The proof is similar to that of the approximation argument in [23, Lemma 3.1]. Recall the constants $\alpha_{0}$ and $\rho_{0}$ appearing in Condition M.2.
Lemma 4.4.
Assume that $\mu$ satisfies Condition M.2 and let $\alpha\in(0,1]$. Suppose that $\varepsilon>0$ and $r\colon X\to\mathbb{R}$ is Lipschitz continuous with $\lvert r\rvert_{\textup{Lip}}\leq 1$ and such that $0<\varepsilon\leq r(x)\leq\rho_{0}$ for all $x\in\operatorname{supp}\mu$. Then there exists an $\alpha$-Hölder continuous function $g\colon X\times X\to[0,1]$ such that $\|g\|_{\alpha}\lesssim\varepsilon^{-1}$,
(4.3) $\mathbf{1}_{B(x,r(x))}(y)\leq g(x,y)\leq\mathbf{1}_{B(x,r(x)+\varepsilon)}(y)\quad\forall x,y\in X$
and
(4.4) $\biggl|\int_{X}g(x,y)\mathop{}\!d\mu(y)-\mu\bigl(B(x,r(x))\bigr)\biggr|\lesssim\varepsilon^{\alpha_{0}}\quad\forall x\in\operatorname{supp}\mu.$
Alternatively, one may replace (4.3) with
$\mathbf{1}_{B(x,r(x)-\varepsilon)}(y)\leq g(x,y)\leq\mathbf{1}_{B(x,r(x))}(y)\quad\forall x,y\in X.$
Proof.
Let $Z=\{(x,y)\in X\times X_{0}:d(x,y)<r(x)+\varepsilon\}$ and notice that $Z$ is open by the Lipschitz continuity of $r$. Define
$g(x,y)=\begin{cases}0&(x,y)\notin Z,\\ \min\bigl\{1,\varepsilon^{-1}\operatorname{dist}\bigl((x,y),(X\times X)\setminus Z\bigr)\bigr\}&(x,y)\in Z.\end{cases}$
Then $g$ is a bounded Lipschitz continuous on $X\times X$ (and therefore $\alpha$-Hölder continuous) with $\|g\|_{\alpha}\lesssim\varepsilon^{-1}$. Furthermore, it is clear that (4.3) holds. The inequality in (4.4) is then a straightforward application of Condition M.2. ∎
Finally, we prove the following lemma, which will be employed several times throughout the proofs. It uses Lemma 4.4.
Lemma 4.5.
Assume that the system $(X,\mu,T)$ satisfies Condition D.2 and fix $\alpha\in(0,1]$. There exists a constants $C_{1},C_{2},p>0$ such that, for all $k\leq n$ and any $\alpha$-Hölder continuous function $g\colon X\times X\to[0,1]$ such that $\|g\|_{\alpha}\leq C_{1}n^{p}$, we have the estimate
$\Bigl|\int g(x,T^{k}x)\mathop{}\!d\mu(x)-\iint g(x,y)\mathop{}\!d\mu(x)\mathop{}\!d\mu(y)\Bigr|\leq C_{2}\bigl(n^{-p}+n^{4(d+3)p/\alpha}e^{-\tau_{0}k}\bigr),$
where $d$ is the dimension of $X$.
Under the additional assumption that Condition D.3 holds, we obtain that, for all $k\leq j\leq n$ and any function $\alpha$-Hölder continuous functions $g,h\colon X\times X\to[0,1]$ such that $\|g\|_{\alpha},\|h\|_{\alpha}\leq C_{1}n^{p}$, we have the estimate
$\Bigl|\int g(x,T^{k}x)h(x,T^{j}x)\mathop{}\!d\mu(x)-\iint g(y,T^{k}x)h(y,T^{j}x)\mathop{}\!d\mu(x)\mathop{}\!d\mu(y)\Bigr|\\ \leq C_{2}\bigl(n^{-p}+n^{4(d+3)p/\alpha}e^{-\tau_{r}k}\bigr).$
Proof.
For the first part of the lemma, obtain the quasi-partitions $(\mathcal{Q}_{n})$ of $X$ and densities $(\rho_{Q,n})_{Q\in\mathcal{Q}_{n}}$ by applying Lemma 4.4 to $\kappa_{0}=2p/\alpha$. For all $Q\in\mathcal{Q}_{n}$ and any $y,y_{Q}\in Q$ and all $x\in X$, we have that
(4.5) $|g(y_{Q},x)-g(y,x)|\leq C_{1}n^{p}\lvert\mathcal{Q}_{n}\rvert^{\alpha}\lesssim n^{-p}.$
Additionally, since $\bigcup_{Q\in\mathcal{Q}_{n}}Q=X\setminus Y_{n}$ with $\mu(Y_{n})\lesssim n^{-2p/\alpha}\leq n^{-p}$, we obtain that
$\begin{split}\int_{X}g(x,T^{k}x)\mathop{}\!d\mu(x)&=\sum_{Q\in\mathcal{Q}_{n}}\int_{X}\mathbf{1}_{Q}(x)g(x,T^{k}x)\mathop{}\!d\mu(x)+\mathcal{O}(n^{-p})\\ &=\sum_{Q\in\mathcal{Q}_{n}}\int_{X}\mathbf{1}_{Q}(x)g(y_{Q},T^{k}x)\mathop{}\!d\mu(x)+\mathcal{O}(n^{-p})\end{split}$
Using the densities $\rho_{Q,n}$ and the fact that $\|\rho_{Q,n}-\mu(Q)^{-1}\mathbf{1}_{Q}\|_{L^{1}}\lesssim n^{-2p/\alpha}\lesssim n^{-p}$, we obtain that
$\int_{X}\mu(Q)^{-1}\mathbf{1}_{Q}(x)g(y_{Q},T^{k}x)\mathop{}\!d\mu(x)=\int_{X}\rho_{Q,n}(x)g(y_{Q},T^{k}x)\mathop{}\!d\mu(x)\\ +\mathcal{O}(n^{-p}).$
Hence,
(4.6) $\int_{X}g(x,T^{k}x)\mathop{}\!d\mu(x)=\sum_{Q\in\mathcal{Q}_{n}}\mu(Q)\int_{X}\rho_{Q,n}(x)g(y_{Q},T^{k}x)\mathop{}\!d\mu(x)+\mathcal{O}(n^{-p}).$
Using Condition D.2, we obtain that
$\begin{split}\int_{X}\rho_{Q,n}(x)g(&y_{Q},T^{k}x)\mathop{}\!d\mu(x)\\ &=\int_{X}\rho_{Q,n}\mathop{}\!d\mu\int_{X}g(y_{Q},x)\mathop{}\!d\mu(x)+\mathcal{O}(\|\rho_{Q,n}\|_{\alpha}\|g\|_{\alpha}e^{-\tau_{0}k})\\ &=\int_{X}g(y_{Q},x)\mathop{}\!d\mu(x)+\mathcal{O}(n^{-p}+n^{4(d+3)p/\alpha}e^{-\tau_{0}k}),\end{split}$
where we have used that $\|\rho_{Q,n}\|_{L^{1}}=1+\mathcal{O}(n^{-p})$, $\|\rho_{Q,n}\|_{\textup{Lip}}\lesssim n^{4(d+2)p/\alpha}$ and $\|g\|_{\alpha}\lesssim n^{p}$. In view of (4.6), we therefore have that
$\int_{X}g(x,T^{k}x)\mathop{}\!d\mu(x)=\sum_{Q\in\mathcal{Q}_{n}}\mu(Q)\int_{X}g(y_{Q},x)\mathop{}\!d\mu(x)\\ +\mathcal{O}(n^{-p}+n^{4(d+3)p/\alpha}e^{-\tau_{0}k}).$
Using (4.5) again, we obtain that
$\Bigl|\mu(Q)\int_{X}g(y_{Q},x)\mathop{}\!d\mu(x)-\int_{Q}\int_{X}g(y,x)\mathop{}\!d\mu(x)\mathop{}\!d\mu(y)\Bigr|\lesssim n^{-p}$
Therefore,
$\int_{X}g(x,T^{k}x)\mathop{}\!d\mu(x)=\int_{X\setminus Y_{n}}\int_{X}g(y,x)\mathop{}\!d\mu(x)\mathop{}\!d\mu(y)\\ +\mathcal{O}(n^{-p}+n^{4(d+3)p/\alpha}e^{-\tau_{0}k}),$
where we have again used that $\bigcup_{Q\in\mathcal{Q}_{n}}Q=X\setminus Y_{n}$ with $\mu(Y_{n})\lesssim n^{-p}$. Since none of the suppressed constants depends on $g$, we obtain the desired result.
For the second part, assume that Condition D.3 holds. Using the same partition as before and in the same way (4.6) was established, we obtain
(4.7) $\int_{X}g(x,T^{k}x)h(x,T^{j}x)\mathop{}\!d\mu(x)\\ =\sum_{Q\in\mathcal{Q}_{n}}\mu(Q)\int_{X}\rho_{Q,n}(x)g(y_{Q},T^{k}x)h(y_{Q},T^{j}x)\mathop{}\!d\mu(x)+\mathcal{O}(n^{-p}),$
where $y_{Q}\in Q$ are arbitrary points. Using Condition D.3,
$\begin{split}\int_{X}\rho_{Q,n}(x)g(y_{Q},T^{k}x)h(y_{Q},&T^{j}x)\mathop{}\!d\mu(x)\\ &=\int_{X}\rho_{Q,n}\mathop{}\!d\mu\int_{X}g(y_{Q},T^{k}x)h(y_{Q},T^{j}x)\mathop{}\!d\mu(x)\\ &\quad+\mathcal{O}(\|\rho_{Q,n}\|_{\alpha}\|g\|_{\alpha}\|h\|_{\alpha}e^{-\tau_{r}k}).\end{split}$
Since $\mu(\rho_{Q,n})=1+\mathcal{O}(n^{-2p/\alpha})$, we have that
$\int_{X}\rho_{Q,n}(x)g(y_{Q},T^{k}x)h(y_{Q},T^{j}x)\mathop{}\!d\mu(x)=\int_{X}g(y_{Q},T^{k}x)h(y_{Q},T^{j}x)\mathop{}\!d\mu(x)\\ +\mathcal{O}(n^{-2p/\alpha}+n^{4(d+3)p/\alpha}e^{-\tau_{r}k}),$
where we have used the estimates $\|\rho_{Q,n}\|_{\alpha}\lesssim n^{4(d+2)p/\alpha}$ and $\|g\|_{\alpha},\|h\|_{\alpha}\lesssim n^{p}$. In view of (4.7), we therefore have that
$\int_{X}g(x,T^{k}x)h(x,T^{j}x)\mathop{}\!d\mu(x)\\ =\sum_{Q\in\mathcal{Q}_{n}}\mu(Q)\int_{X}g(y_{Q},T^{k}x)h(y_{Q},T^{j}x)\mathop{}\!d\mu(x)+\mathcal{O}(n^{-p}+n^{4(d+3)p/\alpha}e^{-\tau_{r}k}),$
Following the proof of the first part, we obtain
$\int_{X}g(x,T^{k}x)h(x,T^{j}x)\mathop{}\!d\mu(x)=\iint_{X\times X}g(y,T^{k}x)h(y,T^{j}x)\mathop{}\!d\mu(x)\mathop{}\!d\mu(y)\\ +\mathcal{O}(n^{-p}+n^{4(d+3)p/\alpha}e^{-\tau_{0}k}),$
as desired. ∎
5. Proof of CLT for Recurrence
We devote this section to the proof of Theorems 2.1 and 2.2. A central argument in the proof of Theorem 2.1 is Lévy’s continuity theorem: the sequence of observables $(f_{n})$ converges in distribution to $f$ if and only if the characteristic functions $\psi_{f_{n}}(t)$ converge to $\psi_{f}(t)$ as $n\to\infty$ for all $t\in\mathbb{R}$ (see e.g. [22, Theorem 5.3]). In particular, we use the fact that if $(f_{n})$ and $(g_{n})$ are sequences of observables such that $g_{n}\to g$ in distribution and $|\psi_{f_{n}}(t)-\psi_{g_{n}}(t)|\to 0$ for all $t\in\mathbb{R}$, then $f_{n}\to g$ in distribution. Another argument that we repeatedly use is the following. Let $f,g$ be two observables. Then, for all $t\in\mathbb{R}$,
(5.1) $|\psi_{f}(t)-\psi_{g}(t)|\leq\int_{X}\bigl|e^{itf(x)}-e^{itg(x)}\bigr|\mathop{}\!d\mu(x)\leq|t|\cdot\|f-g\|_{L^{1}}.$
We momentarily consider a general sequence of functions $(r_{n})$ in the following lemma, which we will then apply separately to the implicit and explicit radii problems.
Lemma 5.1.
Suppose that $(X,T,\mu,d)$ satisfies Conditions D.3, M.1 and M.2. Suppose that $(r_{n})$ is a sequence of Lipschitz continuous functions $r_{n}\colon X\to[0,\infty)$ such that $\lvert r_{n}\rvert_{\textup{Lip}}\leq 1$, $r_{n}\to 0$ uniformly on $\operatorname{supp}\mu$, and $r_{n}\gtrsim n^{-q}$ for some constant $q>0$. Let $(a_{n})$ be a sequence in $(0,\infty)$ such that $a_{n}\lesssim n$. Then, for all $t\in\mathbb{R}$ and all $n\in\mathbb{N}$,
$\biggl|\int_{X}\exp\biggl(\frac{it}{a_{n}}\sum_{k=1}^{n}\mathbf{1}_{B(x,r_{k}(x))}(T^{k}x)-\mu\bigl(B(x,r_{k}(x))\bigr)\biggr)\mathop{}\!d\mu(x)\\ -\iint_{X\times X}\exp\biggl(\frac{it}{a_{n}}\sum_{k=1}^{n}\mathbf{1}_{B(y,r_{k}(y))}(T^{k}x)-\mu\bigl(B(y,r_{k}(y))\bigr)\biggr)\mathop{}\!d\mu(x)\mathop{}\!d\mu(y)\biggr|\\ \lesssim_{t}\frac{\log n}{a_{n}}.$
Recall the constants $s_{0},\rho_{0}\alpha_{0}>0$ from Conditions M.1 and M.2, and the constant $\tau_{r}$ and polynomial $P_{r}$ from Condition D.3. Recall that we may assume, without loss of generality, that $r_{n}\leq\rho_{0}$ $\mu$-a.e. for all $n\in\mathbb{N}$.
Proof of Lemma 5.1.
Let $B_{k}(x)=B(x,r_{k}(x))$ and denote by $\psi_{n}(t)$ and $\psi_{n}(t,y)$ the characteristic functions of
$x\mapsto\frac{1}{a_{n}}\sum_{k=1}^{n}\mathbf{1}_{B_{k}(x)}(T^{k}x)-\mu(B_{k}(x)),$
$x\mapsto\frac{1}{a_{n}}\sum_{k=1}^{n}\mathbf{1}_{B_{k}(y)}(T^{k}x)-\mu(B_{k}(y)),$
respectively. The conclusion of the lemma can now be formulated as
(5.2) $\Bigl|\psi_{n}(t)-\int_{X}\psi_{n}(t,y)\mathop{}\!d\mu(y)\Bigr|\lesssim_{t}\frac{\log n}{a_{n}}.$
Let $\kappa>\max\{q,2,2/\alpha_{0}\}$ and obtain the quasi-partition $\mathcal{Q}_{n}$ using Lemma 4.4 applied to $\kappa_{0}=\kappa$. Let $\varepsilon_{n}=\max\{2\lvert\mathcal{Q}_{n}\rvert^{1/2},2n^{-\kappa/2}\}$ and notice that $\varepsilon_{n}\asymp n^{-\kappa/2}$ since $\lvert\mathcal{Q}_{n}\rvert\lesssim n^{-\kappa}$. Let $\beta_{n}=D\log n$ for some constant $D>0$ satisfying
(5.3) $n^{1+2(d+3)\kappa}P_{r}(n)e^{-\tau_{r}\beta_{n}}\lesssim n^{-1},$
By the choice of $\kappa>0$ and since $r_{n}\to 0$ uniformly on $\operatorname{supp}\mu$, we have that $\varepsilon_{n}\leq r_{n}\leq\rho_{0}$ on $\operatorname{supp}\mu$ for all sufficiently large $n$. We can therefore, for each $k\leq n$, apply Lemma 4.4 with $\varepsilon=\varepsilon_{n}$ and $r=r_{k}$ to obtain functions $g_{k,n},h_{k,n}\colon X\times X\to[0,1]$ satisfying $\|g_{k,n}\|_{\textup{Lip}},\|h_{k,n}\|_{\textup{Lip}}\lesssim\varepsilon_{n}^{-1}\lesssim n^{\kappa/2}$,
(5.4) $\Bigl|\int_{X}g_{k,n}(x,y)\mathop{}\!d\mu(y)-\mu\bigl(B(x,r_{k}(x))\bigr)\Bigr|$ $\lesssim\varepsilon_{n}^{\alpha_{0}}\lesssim n^{-\alpha_{0}\kappa/2},$
(5.5) $\biggl|\int_{X}h_{n,k}(x,y)\mathop{}\!d\mu(y)-\mu\bigl(B(x,r_{k}(x))\bigr)\biggr|$ $\lesssim\varepsilon_{n}^{\alpha_{0}}\lesssim n^{-\alpha_{0}\kappa/2},$
uniformly in $x\in\operatorname{supp}\mu$, and
$\mathbf{1}_{B(y,r_{k}(y))}(x)$ $\leq g_{k,n}(x,y)\leq\mathbf{1}_{B(y,r_{k}(y)+\varepsilon_{n})}(x)$ $\forall x,y\in X,$
$\mathbf{1}_{B(y,r_{k}(y)-\varepsilon_{n})}(x)$ $\leq h_{k,n}(x,y)\leq\mathbf{1}_{B(y,r_{k}(y))}(x)$ $\forall x,y\in X.$
For later use, let $\hat{\psi}_{n}(t)$ and $\hat{\psi}_{n}(t,y)$ denote, respectively, the characteristic of
$x$ $\mapsto\frac{1}{a_{n}}\sum_{k=1}^{n}g_{n,k}(x,T^{k}x)-\mu\bigl(g_{n,k}(x,\cdot)\bigr),$
$x$ $\mapsto\frac{1}{a_{n}}\sum_{k=1}^{n}g_{n,k}(y,T^{k}x)-\mu\bigl(g_{n,k}(y,\cdot)\bigr).$
We begin by shifting time by $\beta_{n}$ and considering the characteristic function $\psi_{n,\beta_{n}}(t)$ of
$x\mapsto\frac{1}{a_{n}}\sum_{k=1+\beta_{n}}^{n+\beta_{n}}\mathbf{1}_{B_{k}(x)}(T^{k}x)-\mu\bigl(B_{k}(x)\bigr).$
Let $\psi_{n,\beta_{n}}(t,y),\hat{\psi}_{n,\beta_{n}}(t)$ and $\hat{\psi}_{n,\beta_{n}}(t,y)$ denote the characteristic functions of the corresponding sums. By (5.1), we have that
(5.6) $|\psi_{n}(t)-\psi_{n,\beta_{n}}(t)|\lesssim_{t}\frac{\beta_{n}}{a_{n}}.$
By (5.4) and the fact that $\mathbf{1}_{B_{k}(x)}(y)\leq g_{k,n}(x,y)$,
(5.7) $\int_{X}\bigl|\mu\bigl(B_{k}(x)\bigr)-\mu\bigl(g_{k,n}(x,\cdot)\bigr)\bigr|\mathop{}\!d\mu(x)\lesssim n^{-\alpha_{0}\kappa/2}.$
On the other hand,
$\|\mathbf{1}_{B_{k}(\cdot)}(T^{k}\cdot)-g_{n,k}(\cdot,T^{k}\cdot)\|_{L^{1}}\leq\int_{X}g_{n,k}(x,T^{k}x)-h_{n,k}(x,T^{k}x)\mathop{}\!d\mu(x)$
since $h_{n,k}(x,y)\leq\mathbf{1}_{B_{k}(x)}(y)\leq g_{n,k}(x,y)$. By Lemma 4.5 (with $\alpha=1$ and $p=\kappa/2$), we obtain the estimate
$\biggl|\int g_{k,n}(x,T^{k}x)\mathop{}\!d\mu-\iint g_{k,n}(x,y)\mathop{}\!d\mu(x)\mathop{}\!d\mu(y)\biggr|\lesssim n^{-\kappa/2}+n^{2(d+3)\kappa}e^{-\tau_{r}k}$
and the corresponding estimate for $h_{n,k}$. Furthermore, by (5.4) and (5.5),
$\Bigl|\iint g_{k,n}\mathop{}\!d\mu\mathop{}\!d\mu-\iint h_{k,n}\mathop{}\!d\mu\mathop{}\!d\mu\Bigr|\lesssim n^{-\alpha_{0}\kappa/2}$
Combining the three previous equations, yields
(5.8) $\|\mathbf{1}_{B_{k}(\cdot)}(T^{k}\cdot)-g_{n,k}(\cdot,T^{k}\cdot)\|_{L^{1}}\lesssim n^{-\alpha_{0}\kappa/2}+n^{-\kappa/2}+n^{2(d+3)\kappa}e^{-\tau_{r}k}.$
Combining (5.7) and (5.8), and using that $n^{-\alpha_{0}\kappa/2},n^{\kappa/2}\lesssim n^{-1}$, yields
(5.9) $\begin{split}|\psi_{n,\beta_{n}}(t)-\hat{\psi}_{n,\beta_{n}}(t)|&\lesssim_{t}\frac{1}{a_{n}}\sum_{k=1+\beta_{n}}^{n+\beta_{n}}\bigl(n^{-1}+n^{2(d+3)\kappa}e^{-\tau_{r}k}\bigr)\\ &\lesssim_{t}\frac{1}{a_{n}}\bigl(1+n^{2(d+3)\kappa}e^{-\tau_{r}\beta_{n}}\bigr)\\ &\lesssim_{t}\frac{1}{a_{n}},\end{split}$
where we have used (5.3) in the third line. We employ the quasi-partition $\mathcal{Q}_{n}$ and the Lipschitz continuity of $g_{n,k}$ to obtain that, for all $Q\in\mathcal{Q}_{n}$ and $y_{Q}\in Q$,
$\sum_{Q\in\mathcal{Q}_{n}}\mathbf{1}_{Q}(x)\bigl(g(x,T^{k}x)-g(y_{Q},T^{k}x)\bigr)\leq\sum_{Q\in\mathcal{Q}_{n}}\mathbf{1}_{Q}(x)\varepsilon_{n}^{-1}\lvert\mathcal{Q}_{n}\rvert\lesssim n^{-\kappa/2}$
and, similarly,
$\sum_{Q\in\mathcal{Q}_{n}}\mathbf{1}_{Q}(x)\bigl(\mu(g(x,\cdot))-\mu(g(y_{Q},\cdot))\bigr)\lesssim n^{-\kappa/2}$
since $\varepsilon_{n}\asymp n^{-\kappa/2}$ and $\lvert\mathcal{Q}_{n}\rvert\lesssim n^{-\kappa}$. Therefore, if we let $\hat{\psi}_{n,\beta_{n}}^{\mathcal{Q}_{n}}$ denote the characteristic function of
$x\mapsto a_{n}^{-1}\sum_{k=1+\beta_{n}}^{n+\beta_{n}}\sum_{Q\in\mathcal{Q}_{n}}\mathbf{1}_{Q}(x)\bigl(g(y_{Q},T^{k}x)-\mu(g(y_{Q},\cdot))\bigr)$
then
(5.10) $|\hat{\psi}_{n,\beta_{n}}(t)-\hat{\psi}_{n,\beta_{n}}^{\mathcal{Q}_{n}}(t)|\lesssim_{t}\frac{1}{a_{n}}\sum_{k=1+\beta_{n}}^{n+\beta_{n}}(n^{-\kappa/2}+\mu(Y_{n}))\lesssim_{t}\frac{n^{1-\kappa/2}}{a_{n}}\lesssim_{t}\frac{1}{a_{n}}$
since $\mu(Y_{n})\lesssim n^{-\kappa}$ and $\kappa>2$. Using that elements of $\mathcal{Q}_{n}$ are disjoint, we obtain that
(5.11) $\hat{\psi}_{n,\beta_{n}}^{\mathcal{Q}_{n}}(t)=\sum_{Q}\int_{X}\mathbf{1}_{Q}(x)\exp\biggl(\frac{it}{a_{n}}\sum_{k=1+\beta_{n}}^{n+\beta_{n}}g_{n,k}(y_{Q},T^{k}x)-\mu(g_{n,k}(y_{Q},\cdot))\biggr)\mathop{}\!d\mu(x)$
Using the densities $(\rho_{Q,n})$ and the fact that $\|\rho_{Q,n}-\mu(Q)^{-1}\mathbf{1}_{Q}\|_{L^{1}}\lesssim n^{-\kappa}$, we obtain that, for all $Q\in\mathcal{Q}_{n}$,
(5.12) $\biggl|\mu(Q)^{-1}\int_{X}\mathbf{1}_{Q}(x)\exp\biggl(\frac{it}{a_{n}}\sum_{k=1+\beta_{n}}^{n+\beta_{n}}g_{n,k}(y_{Q},T^{k}x)-\mu(g_{n,k}(y_{Q},\cdot))\biggr)\mathop{}\!d\mu(x)\\ -\int_{X}\rho_{Q,n}\exp\biggl(\frac{it}{a_{n}}\sum_{k=1+\beta_{n}}^{n+\beta_{n}}g_{n,k}(y_{Q},T^{k}x)-\mu(g_{n,k}(y_{Q},\cdot))\biggr)\mathop{}\!d\mu(x)\biggr|\\ \lesssim n^{-\kappa}.$
Using the property that $\exp(f+g)=\exp(f)\exp(g)$ for any observables $f,g$, we can apply Condition D.3 to $N=\beta_{n}$, and the observables $\rho_{Q,n}$ and $\{e^{ita_{n}^{-1}(g_{n,k}(y_{Q},\cdot)-\mu(g_{n,k}(y_{Q},\cdot)))}\}_{k=1+\beta_{n}}^{n+\beta_{n}}$ to obtain that
$\biggl|\int_{X}\rho_{Q,n}\exp\biggl(\frac{it}{a_{n}}\sum_{k=1+\beta_{n}}^{n+\beta_{n}}g_{n,k}(y_{Q},T^{k}x)-\mu(g_{n,k}(y_{Q},\cdot))\biggr)\mathop{}\!d\mu(x)\\ -\mu(\rho_{Q,n})\int_{X}\exp\biggl(\frac{it}{a_{n}}\sum_{k=1+\beta_{n}}^{n+\beta_{n}}g_{n,k}(y_{Q},T^{k}x)-\mu(g_{n,k}(y_{Q},\cdot))\biggr)\mathop{}\!d\mu(x)\biggr|\\ \lesssim_{t}(n^{2(d+2)\kappa}+n^{1+(d+1)\kappa+\kappa/2})P_{r}(n)e^{-\tau_{r}\beta_{n}},$
where we have used that $\|\rho_{Q,n}\|_{\infty}\lesssim n^{(d+1)}\kappa$, $\|\rho_{Q,n}\|_{\textup{Lip}}\lesssim n^{2(d+1)\kappa}$ and $\|g_{n,k}\|_{\alpha}\lesssim n^{\kappa/2}$. Since $|\mu(\rho_{Q,n})-1|\lesssim n^{-\kappa}$, the previous equation together with (5.11) and (5.12) imply that
$\Bigl|\hat{\psi}_{n,\beta_{n}}^{\mathcal{Q}_{n}}(t)-\sum_{Q\in\mathcal{Q}_{n}}\mu(Q)\hat{\psi}_{n,\beta_{n}}(t,y_{Q})\Bigr|\lesssim_{t}n^{-\kappa}+n^{1+2(d+2)\kappa}P_{r}(n)e^{-\tau_{r}\beta_{n}}.$
By (5.3) and since $\kappa>2$, we can bound the right-hand side by $n^{-1}\leq a_{n}^{-1}$. Therefore,
(5.13) $\Bigl|\hat{\psi}_{n,\beta_{n}}^{\mathcal{Q}_{n}}(t)-\sum_{Q\in\mathcal{Q}_{n}}\mu(Q)\hat{\psi}_{n,\beta_{n}}(t,y_{Q})\Bigr|\lesssim_{t}\frac{1}{a_{n}}.$
As before, we may use the Lipschitz continuity of $g_{n,k}$, and hence that of $y\mapsto\hat{\psi}_{n,\beta_{n}}(t,y)$, to obtain that, for each $Q\in\mathcal{Q}_{n}$,
(5.14) $\Bigl|\mu(Q)\hat{\psi}_{n,\beta_{n}}(t,y_{Q})-\int_{Q}\hat{\psi}_{n,\beta_{n}}(t,y)\mathop{}\!d\mu(y)\Bigr|\lesssim_{t}\frac{1}{a_{n}}.$
Using that $Y_{n}=X\setminus\bigcup_{Q\in\mathcal{Q}_{n}}Q$ has measure $\mu(Y_{n})\leq n^{-\kappa}$, we can combine (5.13) and (5.14), and obtain that
$\Bigl|\hat{\psi}_{n,\beta_{n}}^{\mathcal{Q}_{n}}(t)-\int_{X}\hat{\psi}_{n,\beta_{n}}(t,y)\mathop{}\!d\mu(y)\Bigr|\lesssim_{t}\frac{1}{a_{n}}.$
Using this estimate together with (5.6), (5.9) and (5.10), yields
$\Bigl|\psi_{n}(t)-\int_{X}\hat{\psi}_{n,\beta_{n}}(t,y)\mathop{}\!d\mu(y)\Bigr|\lesssim_{t}\frac{\beta_{n}}{a_{n}}.$
A similar argument to the one establishing (5.9) implies that
$\Bigl|\psi_{n}(t)-\int_{X}\hat{\psi}_{n}(t,y)\mathop{}\!d\mu(y)\Bigr|\lesssim_{t}\frac{\beta_{n}}{a_{n}}.$
Since $\beta_{n}\lesssim\log n$, we obtain (5.2) and we conclude the proof. ∎
5.1. Proof of Theorem 2.1
We now consider the implicit radii problem. Let $\psi_{n}(t)$, $\psi_{n}^{\sigma}(t,y)$ and $\psi_{n}^{s}(t,y)$ denote, respectively, the characteristic functions of
$x\mapsto\frac{1}{\sigma_{n}}\sum_{k=1}^{n}(\mathbf{1}_{E_{k}}(x)-\mu(E_{k})),$
$x\mapsto\frac{1}{\sigma_{n}}\sum_{k=1}^{n}(\mathbf{1}_{B(y,r_{k}(y))}(T^{k}x)-M_{k}),$
$x\mapsto\frac{1}{s_{n}(y)}\sum_{k=1}^{n}(\mathbf{1}_{B(y,r_{k}(y))}(T^{k}x)-M_{k}),$
where $\sigma_{n}^{2}=\operatorname{Var}(\sum_{k=1}^{n}\mathbf{1}_{E_{k}})$ and $s_{n}^{2}(y)=\operatorname{Var}(\sum_{k=1}^{n}\mathbf{1}_{B(y,r_{n}(y))}\circ T^{k})$. As before, we assume, without loss of generality, that $r_{n}\leq\rho_{0}$ $\mu$-a.e. for all $n\in\mathbb{N}$.
Proof of Theorem 2.1.
By [35, Lemma 3.1, Lemma 3.2], the radii functions $r_{n}\colon X\to(0,\infty)$ are Lipschitz continuous with $\lvert r_{n}\rvert_{\textup{Lip}}=1$ and $r_{n}\to 0$ uniformly on $\operatorname{supp}\mu$. Using Condition M.1 together with $M_{n}\gtrsim n^{-\gamma}$, one obtains that $r_{n}(x)\gtrsim n^{-\gamma/s_{0}}$ for all $x\in X$. Furthermore, $\sigma_{n}^{2}\lesssim n^{2}$. Therefore, we may apply Lemma 5.1 with $a_{n}=\sigma_{n}$ and to obtain that
$\biggl|\int_{X}\exp\biggl(\frac{it}{\sigma_{n}}\sum_{k=1}^{n}\mathbf{1}_{E_{k}}(x)-M_{k}\biggr)\mathop{}\!d\mu(x)-\int_{X}\psi_{n}^{\sigma}(t,y)\mathop{}\!d\mu(y)\biggr|\lesssim_{t}\frac{\log n}{\sigma_{n}}.$
Since $|\mu(E_{k})-M_{k}|\lesssim e^{-\eta_{0}k}$ by Lemma 4.2, we may use (5.1) to obtain that
$\biggl|\psi_{n}(t)-\int_{X}\exp\biggl(\frac{it}{\sigma_{n}}\sum_{k=1}^{n}\mathbf{1}_{E_{k}}(x)-M_{k}\biggr)\mathop{}\!d\mu(x)\biggr|\lesssim_{t}\frac{\log n}{\sigma_{n}}.$
Hence,
(5.15) $\Bigl|\psi_{n}(t)-\int_{X}\psi_{n}^{\sigma}(t,y)\mathop{}\!d\mu(y)\Bigr|\lesssim_{t}\frac{\log n}{\sigma_{n}}$
We will show that
(5.16) $\lim_{n\to\infty}\int_{X}\bigl|\psi_{n}^{\sigma}(t,y)-\psi_{n}^{s}(t,y)\bigr|\mathop{}\!d\mu(y)=0\quad\text{for all $t\in\mathbb{R}$}$
and that
(5.17) $\lim_{n\to\infty}\int_{X}\psi_{n}^{s}(t,y)\mathop{}\!d\mu(y)=e^{-t^{2}/2}\quad\text{for all $t\in\mathbb{R}$}.$
By Lemma 3.2, there exists $c>0$ such that $\sigma_{n}^{2}\geq c\mathcal{E}_{n}$ for all sufficiently large $n$. At the same time, we have that $\mathcal{E}_{n}\gtrsim n^{1-\gamma}$ by the assumption that $M_{n}\gtrsim n^{-\gamma}$. Hence, $\sigma_{n}\gtrsim n^{(1-\gamma)/2}$ and $\lim_{n\to\infty}\sigma_{n}^{-1}\log n=0$ as required. Therefore, (5.15)–(5.17) imply that $\psi_{n}(t)$ converges to the characteristic function of a standard normal distribution. By Lévy’s continuity theorem, this will imply the theorem.
We first show (5.16). Let $S_{n,y}=\sum_{k=1}^{n}\mathbf{1}_{B(y,r_{k}(y))}\circ T^{k}-M_{k}$. By (5.1),
$\begin{split}\int_{X}\bigl|\psi_{n}^{\sigma}(t,y)-\psi_{n}^{s}(t,y)\bigr|\mathop{}\!d\mu(y)&\lesssim_{t}\int_{X}\Bigl|\frac{1}{\sigma_{n}}-\frac{1}{s_{n}(y)}\Bigr|\int_{X}|S_{n,y}(x)|\mathop{}\!d\mu(x)\mathop{}\!d\mu(y)\\ &\leq\int_{X}\Bigl|\frac{1}{\sigma_{n}}-\frac{1}{s_{n}(y)}\Bigr||s_{n}(y)|\mathop{}\!d\mu(y)\\ &=\int_{X}\Bigl|\frac{s_{n}(y)}{\sigma_{n}}-1\Bigr|\mathop{}\!d\mu(y)\end{split}$
since $\int|S_{n,y}|\mathop{}\!d\mu\leq s_{n}(y)$ by the Cauchy–Schwarz inequality. The assumption that $s_{n}^{2}/\sigma_{n}^{2}\to 1$ in distribution implies that $s_{n}^{2}/\sigma_{n}^{2}\to 1$ in measure. Together with Lemma 3.3, namely that $\|s_{n}^{2}/\sigma_{n}^{2}\|_{L^{1}}\to 1$, one obtains that $s_{n}^{2}/\sigma_{n}^{2}\to 1$ in $L^{1}$ (see e.g. [22, Proposition 4.12]). Since $|s_{n}/\sigma_{n}+1|\geq 1$, we therefore have that
$\lim_{n\to\infty}\int_{X}\Bigl|\frac{s_{n}(y)}{\sigma_{n}}-1\Bigr|\mathop{}\!d\mu(y)\leq\lim_{n\to\infty}\int_{X}\Bigl|\frac{s_{n}^{2}(y)}{\sigma_{n}^{2}}-1\Bigr|\mathop{}\!d\mu(y)=0.$
This establishes (5.16).
To establish (5.17), notice that, for fixed $t\in\mathbb{R}$, the function $y\mapsto\psi_{n}^{s}(t,y)$ is uniformly bounded by $1$ and converges to $\exp(-t^{2}/2)$ as $n\to\infty$ by (2.2). Therefore, by the Dominated Convergence Theorem, we obtain (5.17), as desired. ∎
5.2. Proof of Theorem 2.2 and Corollary 2.1
We now consider the explicit radii problem. Recall that we assume that $r_{n}\to 0$. Therefore, we may assume, without loss of generality, that $r_{n}\leq\rho_{0}$ for all $n\in\mathbb{N}$. Let $\psi_{n}(t)$, $\psi_{n}^{\sigma}(t,y)$ and $\psi_{n}^{s}(t,y)$ denote, respectively, the characteristic functions of
$x\mapsto\frac{1}{\hat{\sigma}_{n}}\sum_{k=1}^{n}\bigl(\mathbf{1}_{B(x,r_{k})}(T^{k}x)-\mu(B(x,r_{k}))\bigr),$
$x\mapsto\frac{1}{\hat{\sigma}_{n}}\sum_{k=1}^{n}\bigl(\mathbf{1}_{B(y,r_{k})}(T^{k}x)-\mu(B(y,r_{k}))\bigr),$
$x\mapsto\frac{1}{\hat{s}_{n}(y)}\sum_{k=1}^{n}\bigl(\mathbf{1}_{B(y,r_{k})}(T^{k}x)-\mu(B(y,r_{k}))\bigr),$
where $\hat{\sigma}_{n}^{2}=\operatorname{Var}(\sum_{k=1}^{n}\mathbf{1}_{E_{k}})$ and $\hat{s}_{n}^{2}(y)=\operatorname{Var}(\sum_{k=1}^{n}\mathbf{1}_{B(y,r_{k})}\circ T^{k})$. As before, we also assume, without loss of generality, that $r_{n}\leq\rho_{0}$ for all $n\in\mathbb{N}$.
Proof of Theorem 2.2.
Since $\int\mu(B(x,r_{n}))\mathop{}\!d\mu(x)=M_{n}\gtrsim n^{-\gamma}$, we can use Condition M.1 to obtain that $r_{n}\gtrsim n^{-\gamma/s_{0}}$. Clearly, $r_{n}\to 0$ and $\lvert r_{n}\rvert_{\textup{Lip}}=0$. Since $\hat{\sigma}_{n}^{2}\lesssim n^{2}$, we may therefore apply Lemma 5.1 with $a_{n}=\hat{\sigma}_{n}$ to obtain that
$\Bigl|\psi_{n}(t)-\int_{X}\psi_{n}^{\sigma}(t,y)\mathop{}\!d\mu(y)\Bigr|\lesssim_{t}\frac{\log n}{\hat{\sigma}_{n}}.$
Let $S_{n,y}=\sum_{k=1}^{n}\mathbf{1}_{B(y,r_{k})}\circ T^{k}-\mu(B(y,r_{k}))$. Using that $\int|S_{n,y}|\mathop{}\!d\mu\leq\hat{s}_{n}(y)$ by the Cauchy–Schwarz inequality, we have that
$\begin{split}\int_{X}\bigl|\psi_{n}^{\sigma}(t,y)-\psi_{n}^{s}(t,y)\bigr|\mathop{}\!d\mu(y)&\lesssim_{t}\int_{X}\Bigl|\frac{1}{\hat{\sigma}_{n}}-\frac{g}{\hat{s}_{n}(y)}\Bigr|\int_{X}|S_{n,y}(x)|\mathop{}\!d\mu(x)\mathop{}\!d\mu(y)\\ &\leq\int_{X}\Bigl|\frac{1}{\hat{\sigma}_{n}}-\frac{g}{\hat{s}_{n}(y)}\Bigr||\hat{s}_{n}(y)|\mathop{}\!d\mu(y)\\ &=\int_{X}\Bigl|\frac{\hat{s}_{n}(y)}{\hat{\sigma}_{n}}-g\Bigr|\mathop{}\!d\mu(y),\end{split}$
the right-hand side of which converges to 0 as $n\to\infty$ by our assumptions. For fixed $t\in\mathbb{R}$, the functions $y\mapsto\psi_{n}^{s}(t,y)$ are uniformly bounded by $1$ and, as $n\to\infty$, converges to $\exp(-g(y)^{2}t^{2}/2)$ for $\mu$-a.e. $y\in X$ by our assumptions. Since $\hat{\sigma}_{n}^{-1}\log n\to 0$ by Lemma 3.2, we conclude. ∎
We can now prove Corollary 2.1. Recall that $\mu$ is absolutely continuous with density $h$ such that $h\in L^{p}(m)$ for some $p>1$ and $h\geq c>0$ for some constant $c>0$.
Proof of Corollary 2.1.
The conditions on $h$ imply that $\mu$ satisfies Conditions M.1 and M.2. Furthermore, the assumptions of Lemma 3.6 are satisfied. Hence, $\hat{s}_{n}/\hat{\sigma}_{n}\to g$ in $L^{1}(\mu)$. The assumptions of Theorem 2.2 are therefore satisfied and we conclude. ∎
6. Proof of ASIP for Non-invertible Systems
We devote this section to proving the ASIP results for non-invertible systems, i.e., the results stated in Section 2.3. Recall that we assume that the transfer operator $P$ exists and satisfies Condition D.1. We begin by proving Theorem 2.3 in which $(\varphi_{n})$ is a sequence of observables and $\sigma_{n}^{2}=\operatorname{Var}(\sum_{k=1}^{n}\varphi_{k}\circ T^{k})$. Recall that the sequence $(M_{n})$ is defined by $M_{n}=\mu(\varphi_{n})$ and $\mathcal{M}_{n}=\sum_{k=1}^{n}M_{k}$.
Proof of Theorem 2.3.
By the abstract framework provided by Theorem 4.1, we need only verify that Conditions A and B hold for an appropriate sequence $(X_{n})$ of reverse martingale differences. We first establish all the assumptions apart from Conditions A and B. For each $n\geq 1$, define $h_{n}=\sum_{k=1}^{n}P^{k}\varphi_{n-k}$ and
$\psi_{n}=\varphi_{n}+h_{n}-h_{n+1}\circ T=\tilde{\varphi}_{n}+\tilde{h}_{n}-\tilde{h}_{n+1}\circ T.$
Then $\mathbf{E}[\psi_{n}\circ T^{n}|T^{-n-1}\mathcal{B}]=\mathbf{E}[\psi_{n}|T^{-1}\mathcal{B}]\circ T^{n}=(P\psi_{n})\circ T^{n}=0$ $\mu$-a.e.  which is the crucial property that allows us the define a reverse martingale difference.
We will show that
(6.1) $\|h_{n}\|_{\infty}$ $\lesssim\log n,$
(6.2) $\|h_{n}\|_{\alpha}$ $\lesssim n^{p},$
but momentarily delay the proof in order to establish some basic properties of the sequence $(\psi_{n}\circ T^{n})$. Let $X_{n}=\psi_{n}\circ T^{n}$ and $\mathcal{B}_{n}=T^{-n}\mathcal{B}$ so that $(X_{n})$ is a sequence of reverse martingale differences with respect to the filtration $(\mathcal{B}_{n})$. Define $\hat{\sigma}_{n}^{2}=\sum_{k=1}^{n}\mathbf{E}[X_{k}^{2}]$ and notice that
$\int_{X}X_{k}X_{j}\mathop{}\!d\mu=\int_{X}\psi_{k}\psi_{j}\circ T^{j-k}\mathop{}\!d\mu=\int_{X}(P\psi_{k})\psi_{j}\circ T^{j-k-1}\mathop{}\!d\mu=0$
for $j>k$. Hence,
$\mathbf{E}\biggl(\sum_{k=1}^{n}X_{k}\biggr)^{2}=\sum_{k=1}^{n}\mathbf{E}[X_{k}^{2}]=\hat{\sigma}_{n}^{2}.$
Since it is also the case that
$\biggl(\sum_{k=1}^{n}X_{k}\biggr)^{2}=\biggl(\sum_{k=1}^{n}\varphi_{k}\circ T^{k}\biggr)^{2}+2\biggl(\sum_{k=1}^{n}\varphi_{k}\circ T^{k}\biggr)(h_{1}\circ T-h_{n+1}\circ T^{n+1})\\ +(h_{1}\circ T-h_{n+1}\circ T^{n+1})^{2},$
we obtain by (6.1) that
$|\hat{\sigma}_{n}^{2}-\sigma_{n}^{2}|\lesssim\sigma_{n}\log n+(\log n)^{2}.$
Since $\sigma_{n}^{2}\gtrsim\mathcal{M}_{n}\gtrsim n^{1-\gamma}$ by Lemma 3.1, we have that $\sigma_{n}\to\infty$ as $n\to\infty$ and $\log n\lesssim\log\sigma_{n}$, which implies that
(6.3) $|\hat{\sigma}_{n}^{2}-\sigma_{n}^{2}|\lesssim\sigma_{n}\log\sigma_{n}.$
Therefore $\hat{\sigma}_{n}^{2}\to\infty$ as $n\to\infty$. Furthermore, by (6.1), we have that
(6.4) $\mathbf{E}[X_{n}^{2}]\lesssim(\log n)^{2}\lesssim(\log\hat{\sigma}_{n})^{2}\lesssim\hat{\sigma}_{n}.$
Finally, let $\varepsilon\in(0,1)$ and define $a_{n}=\hat{\sigma}_{n}^{2-\varepsilon}$. Then $a_{n}/\hat{\sigma}_{n}^{2}=\hat{\sigma}_{n}^{-\varepsilon}$ is non-increasing and $a_{n}/\hat{\sigma}_{n}=\hat{\sigma}_{n}^{1-\varepsilon}$ is non-decreasing. With these properties established, what remains, in order to apply Theorem 4.1, is to prove (6.1) and (6.2), and Conditions A and B.
We begin with establishing (6.1) and (6.2). Let
(6.5) $D>\frac{2(2p+1)}{\tau_{0}}$
and set $\beta_{n}=D\log n$. We have that
(6.6) $\|h_{n}\|_{\infty}\leq\sum_{k=1}^{\beta_{n}}\|\varphi_{n-k}\|_{\infty}+\sum_{k=\beta_{n}+1}^{n}\|P^{k}\varphi_{n-k}\|_{\alpha},$
where we have used that $\|Pf\|_{\infty}\leq\|f\|_{\infty}$ and $\|f\|_{\infty}\leq\|f\|_{\alpha}$. Since $\|\varphi_{n}\|_{\infty}\lesssim 1$ and $\|P^{k}\varphi_{n-k}\|_{\alpha}\lesssim\|\varphi_{n-k}\|_{\alpha}e^{-\tau_{0}k}\lesssim(n-k)^{p}e^{-\tau_{0}k}$ by Condition D.1, we have that
$\|h_{n}\|_{\infty}\lesssim\beta_{n}+\sum_{k=\beta_{n}+1}^{n}(n-k)^{p}e^{-\tau_{0}k}\lesssim\log n+n^{p+1}e^{-\tau_{0}\beta_{n}}.$
By the choice of $D$ (see (6.5)), we have that $n^{p+1}e^{-\tau_{0}\beta_{n}}=n^{p+1-\tau_{0}D}$ converges to 0 and we have therefore established (6.1). For (6.2), notice that
$\|h_{n}\|_{\alpha}\leq\sum_{k=1}^{n}\|P^{k}\varphi_{n-k}\|_{\alpha}\lesssim\sum_{k=1}^{n}\|\varphi_{n-k}\|_{\alpha}e^{-\tau_{0}k}\lesssim\sum_{k=1}^{n}(n-k)^{p}e^{-\tau_{0}k}\lesssim n^{p}.$
It follows directly from (6.1) that
(6.7) $\|\psi_{n}\|_{\infty}\lesssim\log n.$
We also obtain from (6.2) that, upto replacing $P(\psi_{n}^{2})$ by a $\mu$-a.e. identical observable,
(6.8) $\|P(\psi_{n}^{2})\|_{\alpha}\lesssim n^{p}\log n.$
To see the second inequality, notice that the right-hand side of
$P(\psi_{n}^{2})=P\bigl((\varphi_{n}+h_{n})^{2}\bigr)-2h_{n+1}P(\varphi_{n}+h_{n})+h_{n+1}^{2}\quad\text{$\mu$-a.e.}$
consists entirely of Hölder continuous observables. Using that $\|\varphi_{n}\|_{\alpha}\lesssim n^{p}$ together with (6.1) and (6.2), we obtain the desired estimate. We will apply (6.8) only after integrating, and hence the substitution of $P(\varphi_{n}^{2})$ with a Hölder continuous representative does not pose a problem.
We now establish Condition A using Lemma 4.1 (Gàl–Koksma). As such, it suffices to show that
(6.9) $\int_{X}\biggl(\sum_{k=m}^{n}\mathbf{E}[X_{k}^{2}|\mathcal{B}_{k+1}]-\mathbf{E}[X_{k}^{2}]\biggr)^{2}\mathop{}\!d\mu\lesssim\sum_{k=m}^{n}\gamma_{k}\quad\text{for all $m\leq n$},$
where $\gamma_{k}\geq\mathbf{E}[X_{k}^{2}]$ and
(6.10) $\Gamma(N):=\sum_{k=1}^{N}\gamma_{k}\lesssim\hat{\sigma}_{N}^{2}(\log\hat{\sigma}_{N})^{3}.$
Indeed, by Lemma 4.1, we would then obtain, for all $\varepsilon>0$, that
$\begin{split}\biggl|\sum_{k=1}^{N}\mathbf{E}[X_{k}^{2}|\mathcal{B}_{k+1}]-\mathbf{E}[X_{k}^{2}]\biggr|&\lesssim_{\varepsilon}\Gamma(N)^{1/2}(\log\Gamma(N))^{3/2+\varepsilon}+\max_{1\leq k\leq N}\mathbf{E}[X_{k}^{2}]\\ &\lesssim_{\varepsilon}\hat{\sigma}_{N}(\log\hat{\sigma}_{N})^{3+\varepsilon}\\ &=o_{\varepsilon}(a_{n})\end{split}$
since $\max_{1\leq k\leq N}\mathbf{E}[X_{k}^{2}]\lesssim\hat{\sigma}_{N}$ by (6.4), and $a_{n}=\hat{\sigma}_{n}^{2-\varepsilon}$.
To establish (6.9) write the left-hand side as $A_{m,n}+2B_{m,n}^{(1)}+2B_{m,n}^{(2)}$, where
$A_{m,n}$ $=\sum_{k=m}^{n}\int_{X}(\mathbf{E}[X_{k}^{2}|\mathcal{B}_{k+1}]-\mathbf{E}[X_{k}^{2}])^{2}\mathop{}\!d\mu,$
$B_{m,n}^{(1)}$ $=\sum_{k=m}^{n-1}\sum_{j=k+1}^{k+\beta_{k}}\int_{X}\mathbf{E}[X_{k}^{2}|\mathcal{B}_{k+1}]\mathbf{E}[X_{j}^{2}|\mathcal{B}_{j+1}]-\mathbf{E}[X_{k}^{2}]\mathbf{E}[X_{j}^{2}]\mathop{}\!d\mu,$
$B_{m,n}^{(2)}$ $=\sum_{k=m}^{n-1}\sum_{j=k+\beta_{k}+1}^{n}\int_{X}\mathbf{E}[X_{k}^{2}|\mathcal{B}_{k+1}]\mathbf{E}[X_{j}^{2}|\mathcal{B}_{j+1}]-\mathbf{E}[X_{k}^{2}]\mathbf{E}[X_{j}^{2}]\mathop{}\!d\mu.$
We proceed by estimating each term separately.
It is straightforward to estimate $A_{m,n}$, namely,
(6.11) $\begin{split}|A_{m,n}|&\leq\sum_{k=m}^{n}\int_{X}\mathbf{E}[X_{k}^{2}|\mathcal{B}_{k+1}]^{2}\mathop{}\!d\mu+\mathbf{E}[X_{k}^{2}]^{2}\\ &\leq\sum_{k=m}^{n}\|X_{k}^{2}\|_{\infty}\mathbf{E}[X_{k}^{2}]\\ &\lesssim\sum_{k=m}^{n}(\log k)^{2}\mathbf{E}[X_{k}^{2}],\end{split}$
where we have used $\|X_{k}^{2}\|_{\infty}\leq C(\log k)^{2}$, which follows from (6.7).
For $B_{m,n}^{(1)}$, notice that
$B_{m,n}^{(1)}=\sum_{k=m}^{n-1}\sum_{j=k+1}^{k+\beta_{k}}\int_{X}\mathbf{E}[X_{k}^{2}|\mathcal{B}_{k+1}]\bigl(\mathbf{E}[X_{j}^{2}|\mathcal{B}_{j+1}]-\mathbf{E}[X_{j}^{2}]\bigr)\mathop{}\!d\mu.$
Hence,
(6.12) $\begin{split}|B_{m,n}^{(1)}|&\leq\sum_{k=m}^{n-1}\sum_{j=k+1}^{k+\beta_{k}}\int_{X}\mathbf{E}[X_{k}^{2}|\mathcal{B}_{k+1}]\bigl|\mathbf{E}[X_{j}^{2}|\mathcal{B}_{j+1}]-\mathbf{E}[X_{j}^{2}]\bigr|\mathop{}\!d\mu\\ &\leq\sum_{k=m}^{n-1}\sum_{j=k+1}^{k+\beta_{k}}2\|X_{j}^{2}\|_{\infty}\int_{X}\mathbf{E}[X_{k}^{2}|\mathcal{B}_{k+1}]\mathop{}\!d\mu.\end{split}$
Now $k+\beta_{k}\leq(1+D)k$ for all $k$. In combination with (6.7), we obtain that $\|X_{j}^{2}\|_{\infty}\lesssim(\log k)^{2}$ for all $j\leq k+\beta_{k}$. Hence, from (6.12) we obtain that
(6.13) $\begin{split}|B_{m,n}^{(1)}|&\lesssim\sum_{k=m}^{n-1}\sum_{j=k+1}^{k+\beta_{k}}(\log k)^{2}\mathbf{E}[X_{k}^{2}]\\ &\lesssim\sum_{k=m}^{n}(\log k)^{3}\mathbf{E}[X_{k}^{2}],\end{split}$
For $B_{m,n}^{(2)}$, we have, using the fact that $\mathbf{E}[X_{j}^{2}|\mathcal{B}_{j+1}]$ is $\mathcal{B}_{k+1}$-measurable for $j\geq k$, that
$\begin{split}B_{m,n}^{(2)}&=\sum_{k=m}^{n-1}\sum_{j=k+\beta_{k}+1}^{n}\int_{X}\mathbf{E}[X_{k}^{2}|\mathcal{B}_{k+1}]\mathbf{E}[X_{j}^{2}|\mathcal{B}_{j+1}]\mathop{}\!d\mu-\mathbf{E}[X_{k}^{2}]\mathbf{E}[X_{j}^{2}]\\ &=\sum_{k=m}^{n-1}\sum_{j=k+\beta_{k}+1}^{n}\int_{X}X_{k}^{2}\mathbf{E}[X_{j}^{2}|\mathcal{B}_{j+1}]\mathop{}\!d\mu-\mathbf{E}[X_{k}^{2}]\mathbf{E}[X_{j}^{2}]\\ &=\sum_{k=m}^{n-1}\sum_{j=k+\beta_{k}+1}^{n}\int_{X}\psi_{k}^{2}\mathbf{E}[\psi_{j}^{2}|\mathcal{B}_{1}]\circ T^{j-k}\mathop{}\!d\mu-\mathbf{E}[\psi_{k}^{2}]\mathbf{E}[\psi_{j}^{2}]\\ &=\sum_{k=m}^{n-1}\sum_{j=k+\beta_{k}+1}^{n}\int_{X}P(\psi_{k}^{2})P(\psi_{j}^{2})\circ T^{j-k}\mathop{}\!d\mu-\mathbf{E}[\psi_{k}^{2}]\mathbf{E}[\psi_{j}^{2}],\end{split}$
where we have used that $\mathbf{E}[f|\mathcal{B}_{1}]=\mathbf{E}[f|T^{-1}\mathcal{B}]=(Pf)\circ T$ $\mu$-a.e. Using decay of correlations (see (2.1)), we obtain that
$\biggl|\int_{X}P(\psi_{k}^{2})P(\psi_{j}^{2})\circ T^{j-k}\mathop{}\!d\mu-\mathbf{E}[\psi_{k}^{2}]\mathbf{E}[\psi_{j}^{2}]\biggr|\lesssim\|P(\psi_{k}^{2})\|_{\alpha}\|\psi_{j}^{2}\|_{L^{1}}e^{-\tau_{0}(j-k)}.$
By (6.7) and (6.8), we therefore have that
(6.14) $\begin{split}B_{m,n}^{(2)}&\lesssim\sum_{k=m}^{n-1}\sum_{j=k+\beta_{k}+1}^{n}\|P(\psi_{k}^{2})\|_{\alpha}\|P(\psi_{j}^{2})\|_{\alpha}e^{-\tau_{0}(j-k)}\\ &\lesssim\sum_{k=m}^{n-1}\sum_{j=k+\beta_{k}+1}^{n}k^{p}(\log k)(\log j)e^{-\tau_{0}(j-k)}\\ &\lesssim\sum_{k=m}^{n-1}k^{p+1}e^{-\tau_{0}\beta_{k}}.\end{split}$
Thus, by combining (6.11), (6.13) and (6.14), we see that
$A_{m,n}+2B_{m,n}^{(1)}+2B_{m,n}^{(2)}\lesssim\sum_{k=m}^{n}\gamma_{k},$
where we have let $\gamma_{k}=(\log k)^{3}\mathbf{E}[X_{k}^{2}]+k^{p+1}e^{-\tau_{0}\beta_{k}}$. Hence, (6.9) is satisfied.
To establish (6.10), notice that $k^{p+1}e^{-\tau_{0}\beta_{k}}=k^{p+1-\tau_{0}D}$ is summable by the choice of $D$ (see (6.5)). Hence,
$\Gamma(N)=\sum_{k=1}^{N}\gamma_{k}\lesssim(\log N)^{3}\sum_{k=1}^{N}\mathbf{E}[X_{k}^{2}]\lesssim\hat{\sigma}_{N}^{2}(\log\hat{\sigma}_{N})^{3},$
where we have used that $\log N\lesssim\log\hat{\sigma}_{N}$. We have therefore established Condition A.
We now verify Condition B for $\nu=2$ similarly to how the authors in [13, Theorem 5.2] do. Recall that $a_{k}=\hat{\sigma}_{n}^{2-\varepsilon}$ for some $\varepsilon\in(0,1)$, where $\hat{\sigma}_{n}^{2}=\sum_{k=1}^{n}\mathbf{E}[X_{k}^{2}]$. By (6.1), we have that $\|\tilde{h}_{n}\|_{\infty}\lesssim\log n$. Hence,
(6.15) $\int_{X}X_{n}^{4}\mathop{}\!d\mu=\int_{X}(\tilde{\varphi}_{n}+\tilde{h}_{n}-\tilde{h}_{n+1}\circ T)^{4}\mathop{}\!d\mu\\ \lesssim(\log n)^{3}\|\tilde{\varphi}_{n}\|_{L^{1}}+\|\tilde{h}_{n}-\tilde{h}_{n+1}\circ T\|_{L^{4}}^{4}.$
As established previously (see (6.6)),
$|\tilde{h}_{n}|\lesssim\sum_{k=1}^{\beta_{n}}|P^{k}\tilde{\varphi}_{n-k}|+n^{p}e^{-\tau_{0}\beta_{n}},$
where $n^{p}e^{-\tau_{0}\beta_{n}}=n^{p-\tau_{0}D}\leq n^{-2}$ by (6.5). Therefore,
(6.16) $\|\tilde{h}_{n}\|_{L^{4}}\lesssim\sum_{k=1}^{\beta_{n}}\|\tilde{\varphi}_{n-k}\|_{L^{4}}+n^{-2}.$
Now, for $1\leq k\leq\beta_{n}$ we have, by the monotonicity of $(M_{n})$, that
$\|\tilde{\varphi}_{n-k}\|_{L^{4}}\lesssim M_{n-k}^{1/4}\bigl(\log(n-k)\bigr)^{3/4}\lesssim M_{n-\beta_{n}}^{1/4}(\log n)^{3/4}.$
Hence, using (6.16) we obtain that
$\|\tilde{h}_{n}\|_{L^{4}}\lesssim\beta_{n}M_{n-\beta_{n}}^{1/4}(\log n)^{3/4}\lesssim M_{n-\beta_{n}}^{1/4}(\log n)^{2}.$
Hence, we obtain, by using (6.15), that
$\int_{X}X_{n}^{4}\mathop{}\!d\mu\lesssim M_{n-\beta_{n}}(\log n)^{8}.$
We also have that $a_{n}^{2}=\hat{\sigma}_{n}^{4-2\varepsilon}\gtrsim\sigma_{n}^{4-2\varepsilon}\gtrsim\mathcal{M}_{n}^{2-\varepsilon}$ by (6.3) and Lemma 3.2. Hence,
(6.17) $\sum_{k=1}^{n}a_{k}^{2}\mu(X_{k}^{4})\lesssim\sum_{k=1}^{n}\frac{M_{k-\beta_{k}}(\log k)^{8}}{\mathcal{M}_{k}^{2-\varepsilon}}\lesssim\sum_{k=1}^{n}\frac{M_{k-\beta_{k}}(\log k)^{8}}{\mathcal{M}_{k-\beta_{k}}^{2-\varepsilon}}.$
Now
$\frac{\mathcal{M}_{n}^{2-\varepsilon}}{M_{n}}\geq\biggl(\sum_{k=1}^{n}M_{k}^{1-\frac{1}{2-\varepsilon}}\biggr)^{2-\varepsilon}\gtrsim\biggl(\sum_{k=1}^{n}k^{-\gamma(1-\frac{1}{2-\varepsilon})}\biggr)^{2-\varepsilon}\gtrsim n^{2-\varepsilon-\gamma(1-\varepsilon)}=n^{1+\varepsilon^{\prime}},$
where $\varepsilon^{\prime}=(1-\varepsilon)(1-\gamma)>0$ since $\varepsilon,\gamma<1$. Hence, by (6.17),
$\sum_{k=1}^{n}a_{k}^{2}\mu(X_{k}^{4})\lesssim\sum_{k=1}^{n}\frac{(\log k)^{4}}{(k-\beta_{k})^{1+\varepsilon^{\prime}}}\lesssim\sum_{k=1}^{n}\frac{(\log k)^{4}}{k^{1+\varepsilon^{\prime}}},$
which converges as $n\to\infty$ as required by Condition B.
In conclusion, there exists a sequence of independent centred Gaussian variables $(Z_{k})$ such that $\mathbf{E}[Z_{k}^{2}]=\mathbf{E}[X_{k}^{2}]$ and
$\sup_{1\leq k\leq n}\biggl|\sum_{i=1}^{k}X_{i}-\sum_{i=1}^{k}Z_{i}\biggr|=o\bigl((a_{n}(\log(\sigma_{n}^{2}/a_{n})+\log\log a_{n}))^{1/2}\bigr)\quad\text{$\mathbf{P}$-a.s.}$
In particular,
(6.18) $\sum_{k=1}^{n}\mathbf{E}[Z_{k}^{2}]=\hat{\sigma}_{n}^{2}=\sigma_{n}^{2}+\mathcal{O}(\sigma_{n}\log\sigma_{n})$
by (6.3). Finally, as $a_{n}=\hat{\sigma}_{n}^{2-\varepsilon}$, we have that
$\bigl(a_{n}(\log(\hat{\sigma}_{n}^{2}/a_{n})+\log\log a_{n})\bigr)^{1/2}\lesssim\hat{\sigma}_{n}^{1-\varepsilon/2}\log\hat{\sigma}_{n}.$
Since we are free to choose $\varepsilon\in(0,1)$, we obtain that (ASIP) holds for any $\beta<1/2$. ∎
We now apply Theorem 2.3 to the shrinking target problem. As in Section 5, we may assume without loss of generality that $r_{k}\leq\rho_{0}$ for all $k$ in order to be able to use use Condition M.2 throughout the proof. Indeed, since the centres of the balls $B_{n}=B(y_{n},r_{n})$ have centres $y_{n}\in\operatorname{supp}\mu$, we have that $r_{n}\to 0$. Let $N\in\mathbb{N}$ be such that $r_{n}\leq\rho_{0}$ for all $n\geq N$. Let $S_{n}^{\prime}$ denote the sum obtained by replacing the terms $\tilde{\mathbf{1}}_{B_{1}},\ldots,\tilde{\mathbf{1}}_{B_{N-1}}$ in $S_{n}$ by $\tilde{\mathbf{1}}_{B_{N}}$. Then $S_{n}$ and $S_{n}^{\prime}$ differ at most by a constant. Let $s_{n}^{\prime 2}$ denote the variance of $S_{n}^{\prime}$. A simple calculation shows that
$s_{n}^{2}=s_{n}^{\prime 2}+\mathcal{O}(s_{n}^{\prime}),$
which establishes that $o\bigl(s_{n}^{\prime 1-\beta}\bigr)=o\bigl(s_{n}^{1-\beta}\bigr)$. Therefore, establishing the ASIP for $S_{n}^{\prime}$ implies the ASIP for $S_{n}$.
Proof of Corollary 2.2.
Write $B_{k}=B(y_{k},r_{k})$ for a sequence of points $y_{k}\in\operatorname{supp}\mu$ and a sequence of radii $r_{k}\in[0,\infty)$. By Condition M.1, we have that there exists a constant $C_{r}>0$ such that $r_{k}\geq C_{r}k^{-\gamma/s_{0}}$ for all $k$. Let $\varepsilon_{k}=C_{r}k^{-p}$ for some $p>\max\{\gamma/s_{0},2/\alpha_{0}\}$. Then $\varepsilon_{k}\leq r_{k}\leq\rho_{0}$ for all $k$ and we may therefore use Lemma 4.4 with $r=r_{k}$, $\varepsilon=\varepsilon_{k}$ and a fixed $y=y_{k}$ to obtain $\alpha$-Hölder continuous functions $f_{k}\colon X\to[0,1]$ such that
$\mathbf{1}_{B_{k}(y_{k},r_{k})}\leq f_{k}\leq\mathbf{1}_{B_{k}(y_{k},r_{k}+\varepsilon_{k})}$
and
$\|f_{k}\|_{\alpha}\lesssim\varepsilon_{k}^{-1}\lesssim k^{p}$
for all $k$. As a consequence,
$0\leq\sum_{k=1}^{n}f_{k}\circ T^{k}-\varphi_{k}\circ T^{k}\leq\sum_{k=1}^{n}\mathbf{1}_{A_{k}}\circ T^{k},$
where $A_{k}=B(y_{k},r_{k}+\varepsilon_{k})\setminus B(y_{k},r_{k})$. Since $\mu(A_{k})\lesssim\varepsilon_{k}^{\alpha_{0}}\lesssim k^{-2}$ by Condition M.2, we obtain that $\sum_{k=1}^{\infty}\mu(A_{k})<\infty$. By the Borel–Cantelli lemma, we this implies that $\sum_{k=1}^{\infty}\mathbf{1}_{A_{k}}\circ T^{k}$ is finite $\mu$-a.e. Furthermore, since $\mu(f_{k})-\mu(\varphi_{k})\leq\mu(A_{k})$ is summable, we have that
(6.19) $\biggl|\sum_{k=1}^{n}\tilde{f}_{k}\circ T^{k}(x)-\sum_{k=1}^{n}\tilde{\varphi}_{k}\circ T^{k}(x)\biggr|\lesssim_{x}1\quad\text{$\mu$-a.e.\ $x\in X$}.$
Let $\hat{\sigma}_{n}^{2}=\operatorname{Var}(\sum_{k=1}^{n}\tilde{f}_{k}\circ T^{k})$. Observe that $(\tilde{f}_{k}\circ T^{k})_{k\geq 1}$ satisfies the conditions of Theorem 2.3 since $\mu(f_{k})\geq\mu(B_{k})\gtrsim n^{-\gamma}$, $\mu(|f_{k}|)=\mu(f_{k})$, $\sup_{k\geq 1}\|f_{k}\|_{\infty}\leq 1$ and $\|f_{k}\|_{\alpha}\lesssim k^{p}$. Therefore, there exists (up to extending to a larger probability space if necessary) a sequence of independent centred Gaussian variables $(Z_{k})$ such that $\sum_{k=1}^{n}\mathbf{E}[Z_{k}^{2}]=\hat{\sigma}_{n}^{2}+\mathcal{O}(\hat{\sigma}_{n}\log\hat{\sigma}_{n})$ and for all $\beta<1/2$
$\sup_{1\leq k\leq n}\biggl|\sum_{i=1}^{k}(f_{i}\circ T^{i}-\mu(f_{i}))-\sum_{i=1}^{k}Z_{i}\biggr|=\mathcal{O}\bigl(\hat{\sigma}_{n}^{1-\beta}\bigr)\quad\text{$\mu$-a.e.}$
By (6.19), we therefore obtain that for all $\beta<1/2$
$\sup_{1\leq k\leq n}\biggl|\sum_{i=1}^{k}(\varphi_{i}\circ T^{i}-\mu(\varphi_{i}))-\sum_{i=1}^{k}Z_{i}\biggr|=\mathcal{O}\bigl(\hat{\sigma}_{n}^{1-\beta}\bigr)\quad\text{$\mu$-a.e.}$
It now suffices to show that $\hat{\sigma}_{n}^{2}=\sigma_{n}^{2}+\mathcal{O}(\log\sigma_{n})$ in order to establish (ASIP) for $(\tilde{\varphi}_{k}\circ T^{k})_{k\geq 1}$.
Recall that $M_{k}=\mu(B_{k})$. We have that
$\sigma_{n}^{2}=\sum_{k=1}^{n}M_{k}-M_{k}^{2}+\sum_{k=1}^{n-1}\sum_{j=k+1}^{n}\int_{X}\varphi_{k}\varphi_{j}\circ T^{j-k}\mathop{}\!d\mu-M_{k}M_{j}$
and
$\hat{\sigma}_{n}^{2}=\sum_{k=1}^{n}\mu(f_{k}^{2})-\mu(f_{k})^{2}+\sum_{k=1}^{n-1}\sum_{j=k+1}^{n}\int_{X}f_{k}f_{j}\circ T^{j-k}\mathop{}\!d\mu-\mu(f_{k})\mu(f_{j})$
We have, by the definition of $f_{k}$, that
$\biggl|\sum_{k=1}^{n}\mu(f_{k}^{2})-\mu(f_{k})^{2}-\sum_{k=1}^{n}M_{k}-M_{k}^{2}\biggr|\lesssim 1$
Hence,
(6.20) $|\hat{\sigma}_{n}^{2}-\sigma_{n}^{2}|\lesssim 1+\biggl|\sum_{k=1}^{n-1}\sum_{j=k+1}^{n}\operatorname{Cov}(f_{k},f_{j}\circ T^{j-k})-\operatorname{Cov}(\varphi_{k},\varphi_{j}\circ T^{j-k})\biggr|.$
We split the covariance estimation up into when jkj-k is small and when it is large.
Let $\beta_{n}=D\log n$ for some constant
$D>\frac{2p+1}{\tau_{0}}.$
Since $\varepsilon_{n}\leq\varepsilon_{k}\leq r_{k}\leq\rho_{0}$ for all $k\leq n$, we may apply Lemma 4.4 with $r=r_{k}$, $\varepsilon=\varepsilon_{n}$ and a fixed $y=y_{k}$ to obtain functions $g_{n,k}\colon X\to[0,1]$ with Hölder norm $\|g_{n,k}\|_{\alpha}\lesssim\varepsilon_{n}^{-1}$ and such that
$\mathbf{1}_{B(y_{k},r_{k})}\leq g_{n,k}\leq\mathbf{1}_{B(y_{k},r_{k}+\varepsilon_{n})}$
and
$|\mu(g_{n,k})-M_{k}|\lesssim\varepsilon_{n}^{\alpha_{0}}$
for all $k\leq n$. As a consequence, we obtain that $|\mu(g_{n,k})\mu(g_{n,j})-M_{k}M_{j}|\lesssim\varepsilon_{n}^{\alpha_{0}}$ and
$\Bigl|\int_{X}g_{n,k}g_{n,j}\circ T^{j-k}\mathop{}\!d\mu-M_{k}M_{j}\Bigr|\lesssim\varepsilon_{n}^{\alpha_{0}}+\varepsilon_{n}^{-1}e^{-\tau_{0}(j-k)}$
by decay of correlations (2.1). Hence,
$\begin{split}\sum_{k=1}^{n-1}\sum_{j=k+\beta_{n}+1}^{n}\int_{X}\varphi_{k}&\varphi_{j}\circ T^{j-k}\mathop{}\!d\mu-M_{k}M_{j}\\ &\leq\sum_{k=1}^{n-1}\sum_{j=k+\beta_{n}+1}^{n}\int_{X}g_{n,k}g_{n,j}\circ T^{j-k}\mathop{}\!d\mu-M_{k}M_{j}\\ &\lesssim\sum_{k=1}^{n-1}\sum_{j=k+\beta_{n}+1}^{n}\varepsilon_{n}^{\alpha_{0}}+\varepsilon_{n}^{-1}e^{-\tau_{0}(j-k)}\\ &\lesssim n^{2}\varepsilon_{n}^{\alpha_{0}}+n\varepsilon_{n}^{-1}e^{-\tau_{0}\beta_{n}}.\end{split}$
Since $n^{2}\varepsilon_{n}^{\alpha_{0}}=n^{2-\alpha_{0}p}$ and $n\varepsilon_{n}^{-1}e^{-\tau_{0}\beta_{0}}=n^{1+p-\tau_{0}D}$ converge to 0 by the choice of $p$ and $D$, we have that
$\sum_{k=1}^{n-1}\sum_{j=k+\beta_{n}+1}^{n}\int_{X}\varphi_{k}\varphi_{j}\circ T^{j-k}\mathop{}\!d\mu-M_{k}M_{j}\lesssim 1.$
By instead considering functions bounding $\varphi_{k}$ by below, we may obtain the reverse inequality and conclude that
(6.21) $\biggl|\sum_{k=1}^{n-1}\sum_{j=k+\beta_{n}+1}^{n}\int_{X}\varphi_{k}\varphi_{j}\circ T^{j-k}\mathop{}\!d\mu-M_{k}M_{j}\biggr|\lesssim 1.$
We treat the covariance estimation of the sequence $(f_{k})$ in a more straightforward manner. By decay of correlations, we obtain that
(6.22) $\begin{split}\sum_{k=1}^{n-1}\sum_{j=k+\beta_{n}+1}^{n}\Bigl|\int_{X}f_{k}f_{j}\circ T^{j-k}\mathop{}\!d\mu-\mu(&f_{k})\mu(f_{j})\Bigr|\\ &\lesssim\sum_{k=1}^{n-1}\sum_{j=k+\beta_{n}+1}^{n}\varepsilon_{k}^{-1}e^{-\tau(j-k)}\\ &\lesssim n^{p+1}e^{-\tau_{0}\beta_{n}}\\ &\lesssim 1\end{split}$
since $\varepsilon_{k}^{-1}\lesssim n^{p}$ for $k\leq j\leq n$ and $\beta_{n}=D\log$ for $D>(p+1)/\tau_{0}$. Hence, by combining (6.20)–(6.22), we obtain that
$|\hat{\sigma}_{n}^{2}-\sigma_{n}^{2}|\lesssim 1+\biggl|\sum_{k=1}^{n-1}\sum_{j=k+1}^{k+\beta_{n}}\operatorname{Cov}(f_{k},f_{j}\circ T^{j-k})-\operatorname{Cov}(\varphi_{k},\varphi_{j}\circ T^{j-k})\biggr|.$
By the Cauchy–Schwarz inequality,
$\begin{split}\biggl|\int_{X}f_{k}f_{j}\circ T^{j-k}\mathop{}\!d\mu-\int_{X}\varphi_{k}&\varphi_{j}\circ T^{j-k}\mathop{}\!d\mu\biggr|\\ &\leq\|f_{k}-\varphi_{k}\|_{L^{2}}\|f_{j}\|_{L^{2}}+\|f_{j}-\varphi_{j}\|_{L^{2}}\|\varphi_{k}\|_{L^{2}}\\ &\lesssim\varepsilon_{k}^{\alpha_{0}/2}M_{j}^{1/2}+\varepsilon_{j}^{\alpha_{0}}M_{k}^{1/2}\\ &\lesssim\varepsilon_{k}^{\alpha_{0}/2}\end{split}$
since $\varepsilon_{j}\leq\varepsilon_{k}$ whenever $k\leq j$. Therefore,
(6.23) $\sum_{k=1}^{n-1}\sum_{j=k+1}^{k+\beta_{n}}\Bigl|\int_{X}f_{k}f_{j}\circ T^{j-k}\mathop{}\!d\mu-\int_{X}\varphi_{k}\varphi_{j}\circ T^{j-k}\mathop{}\!d\mu\Bigr|\lesssim\sum_{k=1}^{n-1}\sum_{j=k+1}^{k+\beta_{n}}\varepsilon_{k}^{\alpha_{0}/2}\lesssim\beta_{n}$
since $\varepsilon_{k}^{\alpha_{0}/2}=k^{-\alpha_{0}p/2}$ is summable as we chose $p>2/\alpha_{0}$. At the same time, $|\mu(f_{k})\mu(f_{j})-M_{k}M_{j}|\lesssim\varepsilon_{k}^{\alpha_{0}}$ and
(6.24) $\sum_{k=1}^{n-1}\sum_{j=k+1}^{k+\beta_{n}}|M_{k}M_{j}-\mu(f_{k})\mu(f_{j})|\lesssim\sum_{k=1}^{n-1}\sum_{j=k+1}^{k+\beta_{n}}\varepsilon_{k}^{\alpha_{0}}\lesssim\beta_{n},$
again, since $\varepsilon_{k}^{\alpha_{0}}=k^{-p\alpha_{0}}$ is summable. By combining (6.23) and (6.24), we obtain that
$\sum_{k=1}^{n-1}\sum_{j=k+1}^{k+\beta_{n}}|\operatorname{Cov}(f_{k},f_{j}\circ T^{j-k})-\operatorname{Cov}(\varphi_{k},\varphi_{j}\circ T^{j-k})|\lesssim\beta_{n}$
and therefore, by (6.20),
$|\hat{\sigma}_{n}^{2}-\sigma_{n}^{2}|\lesssim\beta_{n}.$
Finally, since $\beta_{n}=D\log n$ and $\sigma_{n}^{2}\gtrsim\mathcal{M}_{n}\gtrsim n^{1-\gamma}$ by Lemma 3.1, we obtain that
$|\hat{\sigma}_{n}^{2}-\sigma_{n}^{2}|\lesssim\log\sigma_{n}$
as required. ∎
We prove the ASIP for sequences of functions that are not uniformly bounded in $\|\cdot\|_{\infty}$. It follows for the most part the proof of Theorem 2.3.
Proof of Theorem 2.4.
As before for each $n$, let $h_{n}=\sum_{k+1}^{n}P^{k}\varphi_{n-k}$, $\psi_{n}=\varphi_{n}+h_{n}-h_{n+1}\circ T$, $X_{n}=\psi_{n}\circ T^{n}$, $\mathcal{B}_{n}=T^{-n}\mathcal{B}_{n}$ and $\hat{\sigma}_{n}^{2}=\sum_{k=1}^{n}\mathbf{E}[X_{k}^{2}]$. However, in contrast to the proof of Theorem 2.3, recall that $\delta>1/2$ and let $\varepsilon<\min\{1,2-1/\delta\}$ and $a_{n}=\hat{\sigma}_{n}^{2-\varepsilon}$. The reason for this choice of $\varepsilon$ will become apparent when we establish Condition B.
The proof that Condition A holds is identical to that of the same condition in the proof of Theorem 2.3 except for an additional factor of $(\log n)^{q}$ appearing (6.1). This factor propagates throughout the calculations culminating in
$\|X_{n}\|_{\infty}\lesssim(\log n)^{1+q}.$
This factor does not affect the final asymptotic bound of $o(a_{n})$ and so Condition A remains true.
To establish Condition B for $\nu=2$, we use the same approach as employed in [13, Theorem 5.1]. Notice that, just as in the proof of Theorem 2.3, we have that $\hat{\sigma}_{n}^{2}\gtrsim\sigma_{n}^{2}\gtrsim n^{\delta}$. Thus, since $a_{n}^{2}=\hat{\sigma}_{n}^{4-2\varepsilon}\gtrsim n^{(2-\varepsilon)\delta}$,
$\sum_{n=1}^{\infty}a_{n}^{-2}\mathbf{E}[X_{n}^{4}]\lesssim\sum_{n=1}^{\infty}\frac{(\log n)^{4+4q}}{n^{(2-\varepsilon)\delta}}$
converges as $(2-\varepsilon)\delta>1$. Therefore, by Theorem 4.1, there exists a sequence of independent centred Gaussian variables $(Z_{k})$ such that $\mathbf{E}[Z_{k}^{2}]=\mathbf{E}[X_{k}^{2}]$ and
$\sup_{1\leq k\leq n}\biggl|\sum_{i=1}^{k}X_{i}-\sum_{i=1}^{k}Z_{i}\biggr|=o\bigl((a_{n}(\log(\sigma_{n}^{2}/a_{n})+\log\log a_{n}))^{1/2}\bigr)\quad\text{$\mathbf{P}$-a.s.}$
As in the proof of Theorem 2.3, we obtain (6.18) but with an extra factor of $(\log\sigma_{n})^{q}$ appearing; that is, we obtain that
$\sum_{k=1}^{n}\mathbf{E}[Z_{k}^{2}]=\sigma_{n}^{2}+\mathcal{O}(\sigma_{n}(\log\sigma_{n})^{1+q}).$
Since
$a_{n}\bigl(\log(\hat{\sigma}_{n}^{2}/a_{n})+\log\log a_{n}\bigr)^{1/2}\lesssim\hat{\sigma}_{n}^{1-\varepsilon/2}\log\hat{\sigma}_{n},$
we have thus shown that (ASIP) holds with $\beta=\frac{\varepsilon}{2}$. Since $\varepsilon$ is restricted, we obtain (ASIP) for all $\beta<\min\{\frac{1}{2},1-\frac{1}{2\delta}\}$. ∎
7. Proof of ASIP for Axiom A Diffeomorphisms
Recall that in the statement of Theorem 2.5, $T\colon X\to X$ is an Axiom A diffeomorphism restricted to a mixing component and $\mu$ is a Gibbs measure corresponding to a Hölder continuous observable on $X$. In this setting, it is well known (see [4]) that there exists a (finite) Markov partition of $X$ that allows one to encode $(X,T,\mu)$ as a sub-shift of finite type (SFT), which we denote as $(\Sigma,\sigma,\mu^{*})$. Let $\pi\colon\Sigma\to X$ denote the quotient map corresponding to the aforementioned choice of Markov partition. Then $\pi\circ\sigma=T\circ\pi$ and $\mu=\pi^{*}\mu^{*}$. Recall as well that $(\varphi_{n})$ is a sequence of $\alpha$-Hölder continuous functions such that $\|\varphi_{n}\|_{\infty}\lesssim(\log n)^{q}$ and $\|\varphi_{n}\|_{\alpha}\lesssim n^{p}$ for some constants $p,q>0$. We note the following result as stated by Bowen [4, Lemma 4.1], but it is due to Hirsch and Pugh [16, p. 140].
Lemma 7.1.
There exist $\varepsilon>0$ and $\alpha\in(0,1)$ for which the following is true: if $x,y\in X$ and $d(T^{k}x,T^{k}y)\leq\varepsilon$ for all $k\in[-N,N]$, then $d(x,y)<\alpha^{N}$.
We remark for later that Lemma 7.1 implies that, for any $\alpha\in(0,1]$, there exists $\theta\in(0,1)$ such that, given any $\alpha$-Hölder continuous function $\varphi\colon X\to\mathbb{R}$, one has that $\varphi^{*}=\varphi\circ\pi$ is an element of $\mathcal{F}_{\theta}$ with $\|\varphi^{*}\|_{\theta}\lesssim\|\varphi\|_{\alpha}$.
Our first lemma is a modification of the ‘Sinai trick’ (explained in [4, 29] for instance), which allows one to reduce proofs of certain properties of two-sided SFTs (such as existence of Gibbs measures) to those of one-sided SFTs. The usual Sinai trick is employed to obtain, from a function $f\colon\Sigma\to\mathbb{R}$, a function $f^{+}\colon\Sigma\to\mathbb{R}$ depending only on future coordinates, i.e., $f^{+}(\underline{x})=f^{+}(\underline{y})$ whenever $\underline{x},\underline{y}\in\Sigma$ with $x_{i}=y_{i}$ for all $i\geq 0$. The modification we make accounts for the non-stationary aspect of our problem; we consider a sequence of functions $(\varphi_{n})$ as opposed just a single function $\varphi$. In the case that $(\varphi_{n})$ is uniformly bounded in $\|\cdot\|_{\alpha}$, Haydn et al [13, Corollary 6.2] prove a modified Sinai trick. In our case, in order to apply our results to the shrinking target problem, we do not assume that the sequence $(\varphi_{n})$ is uniformly bounded in $\|\cdot\|_{\alpha}$. We also keep track of the expectation and variance corresponding to the new sequence of functions. Once we establish our modified Sinai trick, we can reduce the problem to the one-sided sub-shift and then use Conditions D.1, M.1 and M.2 since they hold for $(\Sigma^{+},\sigma,\mu^{*})$.
Lemma 7.2 (Non-autonomous Sinai Trick).
Let $\theta\in(0,1)$ and suppose that $(\varphi_{n}^{*})\subset\mathcal{F}_{\theta}$ is a sequence of observables $\varphi_{n}^{*}\colon\Sigma\to\mathbb{R}$ such that $\|\varphi_{n}^{*}\|_{\infty}\lesssim(\log n)^{q}$ and $\|\varphi_{n}^{*}\|_{\theta}\lesssim n^{p}$ for some constants $p,q>0$. Denote the variance corresponding to the sequence $(\varphi_{n}^{*})$ as $\sigma_{n}^{2}=\operatorname{Var}(\sum_{k=1}^{n}\varphi_{k}^{*}\circ\sigma^{k})$. Then there exists $\beta\in(0,1)$ and a sequence $(f_{n})\subset\mathcal{F}_{\beta}$ of observables $f_{n}\colon\Sigma\to\mathbb{R}$ depending only on future coordinates such that $\|f_{n}\|_{\infty}\lesssim(\log n)^{1+q}$, $\|f_{n}\|_{\beta}\lesssim n^{p}$ and
(7.1) $\biggl|\sum_{k=1}^{n}f_{k}\circ\sigma^{k}-\sum_{k=1}^{n}\varphi_{k}^{*}\circ\sigma^{k}\biggr|\lesssim(\log n)^{1+q}\quad\text{$\mu^{*}$-a.e.},$
(7.2) $\biggl|\sum_{k=1}^{n}\mu^{*}(f_{k})-\sum_{k=1}^{n}\mu^{*}(\varphi_{k}^{*})\biggr|\lesssim(\log n)^{1+q},$
(7.3) $\biggl|\operatorname{Var}\biggl(\sum_{k=1}^{n}f_{k}\circ\sigma^{k}\biggr)-\sigma_{n}^{2}\biggr|\lesssim\sigma_{n}(\log n)^{1+q}+(\log n)^{2+q}.$
We modify the proof of the usual Sinai trick (see for instance [4, 29]) and incorporate a modification due to [13, Corollary 6.2].
Proof.
We begin by constructing the sequence of functions $(f_{n})$. For each allowable digit $i$ in $\Sigma$, fix an element $\underline{x}^{i}\in\Sigma$ with $x_{0}^{i}=i$. Define the function $G\colon\Sigma\to\Sigma$ by $\underline{y}=G(\underline{x})$, where
$y_{k}=\begin{cases}x_{k}&k\geq 0,\\ x_{k}^{i}&k<0\text{ and }x_{0}=i.\end{cases}$
This function clearly only depends on future coordinates. For each $n\geq 1$, define
$v_{n}(\underline{x})=\sum_{k=n}^{\infty}\varphi_{k}^{*}(\sigma^{k-n}\underline{x})-\varphi_{k}^{*}(\sigma^{k-n}G\underline{x}).$
Notice that the series converges everywhere and that
(7.4) $\|v_{n}\|_{\infty}\lesssim(\log n)^{1+q}.$
Indeed, let $\beta_{n}=D\log n$ for some constant $D>-p/\log\theta$. Then, for all $\underline{x}\in\Sigma$,
$|v_{n}(\underline{x})|\leq 2\beta_{n}\log(n+\beta_{n})^{q}+\sum_{k=n+\beta_{n}}^{\infty}|\varphi_{k}^{*}(\sigma^{k-n}\underline{x})-\varphi_{k}^{*}(\sigma^{k-n}G\underline{x})|$
since $\|\varphi_{k}^{*}\|_{\infty}\leq\log(n+\beta_{n})^{q}$ for $k\leq n+\beta_{n}$. By the definition of $\beta_{n}$, we have that $\log(n+\beta_{n})\lesssim\beta_{n}$, which, combined with the fact that $(\varphi_{k}^{*})\subset\mathcal{F}_{\theta}$, implies that
$\begin{split}|v_{n}(\underline{x})|&\lesssim\beta_{n}^{1+q}+\sum_{k=n+\beta_{n}}^{\infty}\|\varphi_{k}^{*}\|_{\theta}\theta^{k-n}\\ &\lesssim\beta_{n}^{1+q}+\sum_{k=n+\beta_{n}}k^{p}\theta^{k-n}\\ &\lesssim\beta_{n}^{1+q}+n^{p}\theta^{\beta_{n}}.\end{split}$
Since $n^{p}\theta^{\beta_{n}}=n^{p+D\log\theta}\to 0$ by the choice of $D$, we therefore obtain (7.4). Now, for each $n\geq 1$ and $\underline{x}\in\Sigma$, define
(7.5) $\begin{split}f_{n}(\underline{x})&=\varphi_{n}^{*}(\underline{x})-v_{n}(\underline{x})+v_{n+1}(\sigma\underline{x})\\ &=\varphi_{n}^{*}(G\underline{x})+\sum_{k=n+1}^{\infty}\varphi_{k}^{*}(\sigma^{k-n}G\underline{x})-\varphi_{k}^{*}(\sigma^{k-n-1}G\sigma\underline{x})\end{split}$
and note that $f_{n}$ depends only on future coordinates.
Let $\beta=\theta^{1/3}$. We now show that $f_{n}\in\mathcal{F}_{\beta}$ with $\|f_{n}\|_{\beta}\lesssim n^{p}$ for which it suffices to show that $v_{n}\in\mathcal{F}_{\beta}$ and $\|v_{n}\|_{\beta}\lesssim n^{p}$. We wish to show that
(7.6) $\operatorname{var}_{3N}v_{n}\leq C_{0}n^{p}\theta^{N}$
for each $N\geq 0$. In which case, we obtain that
$\operatorname{var}_{3N+j}v_{n}\leq\Bigl(\frac{C_{0}}{\theta^{2/3}}\Bigr)n^{p}(\theta^{1/3})^{3N+j}\lesssim n^{p}\beta^{3N+j}$
for each $N\geq 0$ and $j=0,1,2$. As a consequence, $v_{n}\in\mathcal{F}_{\beta}$ and
(7.7) $\lvert v_{n}\rvert_{\beta}\lesssim n^{p}.$
Fix $N\geq 0$ and $n\geq 1$. Let $\underline{x},\underline{y}\in\Sigma$ such that $x_{i}=y_{i}$ for all $|i|\leq 3N$. Then
$|\varphi_{k}^{*}(\sigma^{k-n}\underline{x})-\varphi_{k}^{*}(\sigma^{k-n}\underline{y})|$ $\lesssim\|\varphi_{k}^{*}\|_{\theta}\theta^{3N-(k-n)}\quad\text{for all }n\leq k\leq n+N,$
$|\varphi_{k}^{*}(\sigma^{k-n}G\underline{x})-\varphi_{k}^{*}(\sigma^{k-n}G\underline{y})|$ $\lesssim\|\varphi_{k}^{*}\|_{\theta}\theta^{3N-(k-n)}\quad\text{for all }n\leq k\leq n+N,$
and
$|\varphi_{k}^{*}(\sigma^{k-n}\underline{x})-\varphi_{k}^{*}(\sigma^{k-n}G\underline{x})|$ $\leq\|\varphi_{k}^{*}\|_{\theta}\theta^{k-n}\quad\text{for all }k\geq n,$
$|\varphi_{k}^{*}(\sigma^{k-n}\underline{y})-\varphi_{k}^{*}(\sigma^{k-n}G\underline{y})|$ $\leq\|\varphi_{k}^{*}\|_{\theta}\theta^{k-n}\quad\text{for all }k\geq n.$
Therefore,
$\begin{split}|v_{n}(\underline{x})-v_{n}(\underline{y})|&\leq 2\sum_{k=n}^{n+N}\|\varphi_{k}^{*}\|_{\theta}\theta^{3N-(k-n)}+2\sum_{k=n+N}^{\infty}\|\varphi_{k}^{*}\|_{\theta}\theta^{k-n}\\ &\lesssim(n+N)^{p}\sum_{k=n}^{n+N}\theta^{3N-(k-n)}+\sum_{k=n+N}^{\infty}k^{p}\theta^{k-n}\\ &\lesssim n^{p}\theta^{2N}+N^{p}\theta^{2N}\\ &\lesssim n^{p}\theta^{2N}+\theta^{N}\\ &\lesssim n^{p}\theta^{N}.\end{split}$
We have therefore established (7.6), and thus (7.7). In conclusion, we obtain $\|f_{n}\|_{\infty}\lesssim(\log n)^{1+q}$ from (7.4) and the fact that $\|\varphi_{n}^{*}\|_{\infty}\lesssim(\log n)^{q}$. We also obtain that $\lvert f_{n}\rvert_{\beta}\lesssim n^{p}$ from (7.7), and therefore that $\|f_{n}\|_{\beta}\lesssim n^{p}$. This establishes the regularity properties of $(f_{n})$.
We now establish (7.1)–(7.3). We obtain from (7.5) that
$\sum_{k=1}^{n}f_{k}\circ\sigma^{k}-\sum_{k=1}^{n}\varphi_{k}^{*}\circ\sigma^{k}=v_{n+1}\circ\sigma^{n+1}-v_{1}\circ\sigma.$
The estimates in (7.1) and (7.2) now follow from a straightforward application of (7.4). Lastly, notice that
$\biggl(\sum_{k=1}^{n}\tilde{f}_{k}\circ\sigma^{k}\biggr)^{2}=\biggl(\sum_{k=1}^{n}\tilde{\varphi}_{k}^{*}\circ\sigma^{k}\biggr)^{2}+2\biggl(\sum_{k=1}^{n}\tilde{\varphi}_{k}^{*}\circ\sigma^{k}\biggr)(\tilde{v}_{n+1}\circ\sigma^{n+1}-\tilde{v}_{1})\\ +(\tilde{v}_{n+1}\circ\sigma^{n+1}-\tilde{v}_{1})^{2},$
which implies, together with (7.4), that
$\biggl|\operatorname{Var}\biggl(\sum_{k=1}^{n}\tilde{f}_{k}\circ\sigma^{k}\biggr)-\sigma_{n}^{2}\biggr|\lesssim\sigma_{n}(\log n)^{1+q}+(\log n)^{2+q},$
where we have used Hölder’s inequality to establish
$\int\biggl|\sum_{k=1}^{n}\tilde{\varphi}_{k}^{*}\circ\sigma^{k}\biggr|\mathop{}\!d\mu^{*}\leq\sigma_{n}.$
We have thus established (7.3) and we conclude the proof. ∎
The proof of Theorem 2.5 now becomes a simple application of Lemma 7.2.
Proof of Theorem 2.5.
Given the sequence $(\varphi_{n})$ of $\alpha$-Hölder continuous functions $\varphi_{n}\colon X\to\mathbb{R}$, define $\varphi_{n}^{*}=\varphi_{k}\circ\pi$. As a direct consequence, we obtain that $\|\varphi_{n}^{*}\|_{\infty}\lesssim(\log n)^{q}$ and, by the remark after Lemma 7.1, there exists $\theta\in(0,1)$ such that $(\varphi_{n}^{*})\subset\mathcal{F}_{\theta}$ with $\|\varphi_{n}^{*}\|_{\theta}\lesssim n^{p}$. Applying Lemma 7.2 to the sequence $(\varphi_{n}^{*})$, we obtain the sequence $(f_{n})\subset\mathcal{F}_{\beta}$ of functions depending only on future coordinates satisfying $\|f_{n}\|_{\infty}\lesssim(\log n)^{1+q}$, $\|f_{n}\|_{\beta}\lesssim n^{p}$ and (7.1)–(7.3). Since $f_{n}$ depends only on future coordinates, we may view the functions $f_{n}$ as functions on $\Sigma^{+}$ instead.
The corresponding system $(\Sigma^{+},\sigma,\mu^{*})$ satisfies Conditions D.1, M.1 and M.2. Furthermore, the functions $(f_{n})$ satisfy the assumptions of Theorem 2.4 with $q$ in the statement replaced with $q+1$. Hence, $(\tilde{f}_{n}\circ\sigma^{n})$ satisfies (ASIP) for any $\beta<\min\{\frac{1}{2},1-\frac{1}{2\delta}\}$. By the estimates established in Lemma 7.2, $(\tilde{\varphi}_{n}\circ T^{n})$ satisfies (ASIP) for any $\beta<\min\{\frac{1}{2},1-\frac{1}{2\delta}\}$ and we conclude the proof. ∎
We can now apply Theorem 2.5 to the shrinking target problem for Axiom A systems.
Proof of Corollary 2.3.
The proof is identical to the proof of Corollary 2.2, except for the fact that Axiom A systems satisfy Condition D.2 rather than (2.1). However, by selecting a larger constant $D$ in the proof of Corollary 2.2 accounts for this. ∎
8. Proof of Variance Estimates
In this section, we prove the variance estimates stated in Section 3. Recall the sequences $(M_{n})$ and $(E_{n})$, and $(\hat{M}_{n})$ and $(\hat{E}_{n})$ corresponding, respectively, to the implicit and explicit radii problems. For the implicit problem, let $\mathcal{M}_{n}=\sum_{k=1}^{n}M_{k}$ and $\mathcal{E}_{n}=\sum_{k=1}^{n}\mu(E_{k})$, and for the explicit problem, let $\hat{\mathcal{M}}_{n}=\sum_{k=1}^{n}\hat{M}_{k}$ and $\hat{\mathcal{E}}_{n}=\sum_{k=1}^{n}\mu(\hat{E}_{k})$. By Lemma 4.2, we conclude that $\mathcal{M}_{n}$ and $\mathcal{E}_{n}$ differ at most by a (uniform) constant, and that $(\mu(E_{n}))$ satisfies Condition $\mathcal{S}_{\gamma,\upsilon}$ with the same constants $\gamma,\upsilon$ as the sequence $(M_{n})$, but perhaps with the monotonicity condition replaced with monotonicity up to a uniform multiplicative constant. The same reasoning applies for the sums $\hat{\mathcal{E}}_{n}$ and $\hat{\mathcal{M}}_{n}$, and the sequences $(\mu(\hat{E}_{n}))$ and $(\hat{M}_{n})$.
As it will be used throughout this section, we note that Condition $\mathcal{S}_{\gamma,\upsilon}$ implies that, for any $\delta>0$,
(8.1) $\lim_{n\to\infty}\frac{\sum_{k=1}^{n}a_{k}^{1+\delta}}{\sum_{k=1}^{n}a_{k}}=0$
where $(a_{k})$ is any of the sequences $(M_{k})$, $(\hat{M}_{k})$, $(\mu(E_{k}))$ or $(\mu(\hat{E}_{k}))$. For the proofs of the variance estimates, the following lemma estimating the correlations of $(E_{n})$ and $(\hat{E}_{n})$ will be useful. The lemma is similar to [23, Lemma 4.2], [34, Lemma 3] and [35, Proposition 4.2], relaxing the multiple decorrelation conditions in those works by Condition D.3, but obtaining worse estimates as a result.
Lemma 8.1.
Suppose that $(X,\mu,T,d)$ satisfies Conditions D.3, M.1 and M.2. Then, there exists $p_{0}>2$ and $q_{0},\eta>0$ such that, for all $k\leq j\leq n$,
$|\mu(E_{k}\cap E_{j})-\mu(E_{k})\mu(E_{j})|\lesssim n^{-p_{0}}+n^{q_{0}}\bigl(e^{-\eta k}+e^{-\eta(j-k)}\bigr)$
and
(8.2) $\Bigl|\mu(\hat{E}_{k}\cap\hat{E}_{j})-\int_{X}\mu(B(x,r_{k}))\mu(B(x,r_{j}))\mathop{}\!d\mu(x)\Bigr|\\ \lesssim n^{-p_{0}}+n^{q_{0}}\bigl(e^{-\eta k}+e^{-\eta(j-k)}\bigr)$
Remark 10.
We only require that Condition D.3 holds for $3$-fold correlations, i.e., the case when $m+m^{\prime}=3$.
Proof of Lemma 8.1.
We prove only the upper bounds in both inequalities since the lower bounds follow from similar calculations. To show that
$\mu(E_{k}\cap E_{j})\leq\mu(E_{k})\mu(E_{j})+\mathcal{O}\bigl(n^{-p_{0}}+n^{q_{0}}(e^{-\eta_{0}k}+e^{-\eta_{0}(j-k)})\bigr),$
let $\kappa>\max\{4,8/\alpha_{0}\}$ and $\varepsilon_{n}=2n^{-\kappa/2}$. Obtain $g_{n,k}\colon X\times X\to(0,1)$ using Lemma 4.4 applied to $r=r_{k}(x)$ and $\varepsilon=\varepsilon_{n}$. Then
$\mu(E_{k}\cap E_{j})\leq\int_{X}g_{n,k}(x,T^{k}x)g_{n,j}(x,T^{j}x)\mathop{}\!d\mu(x)$
By the second part of Lemma 4.5 with $p=\kappa/2$, we have that
(8.3) $\begin{split}\int g_{n,k}(x,T^{k}x)g_{n,j}(x,T^{j}x)\mathop{}\!d\mu(x)&\leq\iint g_{n,k}(y,T^{k}x)g_{n,j}(y,T^{j}x)\mathop{}\!d\mu(x)\mathop{}\!d\mu(y)\\ &\quad+\mathcal{O}\bigl(n^{-\kappa/2}+n^{(d+3)\kappa/\alpha}e^{-\tau_{r}k}\bigr).\end{split}$
By the first part of Lemma 4.5 applied to the observables $g(x,T^{j-k}x)=g_{n,k}(y,x)g_{n,j}(y,T^{j-k}x)$, we obtain
(8.4) $\begin{split}\iint g_{n,k}(y,T^{k}x)g_{n,j}(y,T^{j}x)&\mathop{}\!d\mu(x)\mathop{}\!d\mu(y)\\ &\leq\iiint g_{n,k}(y,x)g_{n,j}(y,z)\mathop{}\!d\mu(x)\mathop{}\!d\mu(z)\mathop{}\!d\mu(y)\\ &\quad+\mathcal{O}\bigl(n^{-\kappa/2}+n^{(d+3)\kappa/\alpha}e^{-\tau_{r}(j-k)}\bigr).\end{split}$
Now,
$\begin{split}\int_{X}|g_{n,k}(y,x)-\mathbf{1}_{B(y,r_{k}(y))}&(x)|\mathop{}\!d\mu(x)\\ &\leq\int_{X}|\mathbf{1}_{B(y,r_{k}(y)+\varepsilon_{n})}(x)-\mathbf{1}_{B(y,r_{k}(y))}(x)|\mathop{}\!d\mu(x)\\ &\lesssim\varepsilon_{n}^{\alpha_{0}}\lesssim n^{-\alpha_{0}\kappa/4}\end{split}$
by Condition M.2. Hence,
$\int g_{n,k}(x,T^{k}x)g_{n,j}(x,T^{j}x)\mathop{}\!d\mu(x)\leq\int\mu\bigl(B(y,r_{k}(y))\bigr)\mu\bigl(B(y,r_{j}(y))\bigr)\mathop{}\!d\mu(y)\\ +\mathcal{O}\bigl(n^{-\kappa/2}+n^{-\alpha_{0}\kappa/4}+n^{(d+3)\kappa/\alpha}(e^{-\tau_{r}k}+e^{-\tau_{0}(j-k)})\bigr).$
Since $\mu(B(y,r_{k}(y)))=M_{k}$, we obtain that
$\mu(E_{k}\cap E_{j})\leq M_{k}M_{j}+\mathcal{O}\bigl(n^{-\kappa/2}+n^{-\alpha_{0}\kappa/4}+n^{(d+3)\kappa/\alpha}(e^{-\tau_{r}k}+e^{-\tau_{0}(j-k)})\bigr).$
By Lemma 4.2, $|\mu(E_{k})-M_{k}|\lesssim e^{-\eta_{0}k}$, hence
$\mu(E_{k}\cap E_{j})\leq\mu(E_{k})\mu(E_{j})+\mathcal{O}\bigl(n^{-p_{0}}+n^{q_{0}}e^{-\eta(j-k)}\bigr),$
where $\eta=\min\{\tau_{0},\tau_{r},\eta_{0}\}$, $p_{0}=\min\{\kappa/2,\alpha_{0}\kappa/4\}>2$ (by the choice of $\kappa$) and $q_{0}=(d+3)\kappa/\alpha$.
To establish
$\mu(\hat{E}_{k}\cap\hat{E}_{j})\leq\int\mu(B(x,r_{k}))\mu(B(x,r_{j}))\mathop{}\!d\mu(x)+\mathcal{O}\bigl(n^{-p_{0}}+n^{q_{0}}(e^{-\eta k}+e^{-\eta(j-k)}\bigr)$
and therefore (8.2), let $\kappa>\max\{2,4/\alpha_{0}\}$ and $\varepsilon_{n}=2n^{-\kappa/2}$. Obtain $g_{n,k}\colon X\times X\to(0,1)$ using Lemma 4.4 applied to $r=r_{k}$ and $\varepsilon=\varepsilon_{n}$. Then one obtains, similarly to how one establishes (8.3) and (8.4), that
$\mu(\hat{E}_{k}\cap\hat{E}_{j})\leq\iiint g_{n,k}(y,x)g_{n,j}(y,z)\mathop{}\!d\mu(x)\mathop{}\!d\mu(z)\mathop{}\!d\mu(y)\\ +\mathcal{O}\bigl(n^{-\kappa/2}+n^{(d+3)\kappa/\alpha}(e^{-\tau_{r}k}+e^{-\tau_{0}(j-k)})\bigr).$
Since
$\begin{split}\int|g_{n,k}(y,x)-\mathbf{1}_{B(y,r_{k})}(x)|\mathop{}\!d\mu(x)&\leq\int|\mathbf{1}_{B(y,r_{k}+\varepsilon_{n})}(x)-\mathbf{1}_{B(y,r_{k})}(x)|\mathop{}\!d\mu(x)\\ &\lesssim\varepsilon_{n}^{\alpha_{0}}\lesssim n^{-\alpha_{0}\kappa/2}\end{split}$
by Condition M.2, we have that
$\mu(\hat{E}_{k}\cap\hat{E}_{j})\leq\int_{X}\mu(B(y,r_{k}))\mu(B(y,r_{j}))\mathop{}\!d\mu(y)\\ +\mathcal{O}\bigl(n^{-p_{0}}+n^{q_{0}}(e^{-\eta k}+e^{-\eta(j-k)})\bigr).$
Choosing $\eta=\min\{\tau_{0},\tau_{r}\}$, $p_{0}=\min\{\kappa/2,\alpha_{0}\kappa/4\}>2$ (by the choice of $\kappa$) and $q_{0}=(d+3)\kappa/\alpha$. ∎
8.1. Proof of Lemma 3.1
We first prove the variance estimate for the shrinking target problem.
Proof of Lemma 3.1.
We show the part of the lemma involving $s_{n}^{2}$ since the first part follows from a similar and more straightforward argument. Recall that we may, without loss of generality, assume that $r_{k}(y)\leq\rho_{0}$ for all $k$ and $y\in\operatorname{supp}\mu$. By expanding the expression for $s_{n}^{2}(y)$, we obtain that
$s_{n}^{2}(y)=\sum_{k=1}^{n}(M_{k}-M_{k}^{2})+\sum_{k=1}^{n-1}\sum_{j=k+1}^{n}\operatorname{Cov}(\mathbf{1}_{B(y,r_{k}(y))},\mathbf{1}_{B(y,r_{j}(y))}\circ T^{j-k}).$
Now
(8.5) $\sum_{k=1}^{n}(M_{k}-M_{k}^{2})=\mathcal{M}_{n}-o(\mathcal{M}_{n})$
by (8.1). We split the double sum into two parts: when $j\leq k+\beta_{k}$ and when $j>k+\beta_{k}$.
Let
(8.6) $\kappa>\max\Bigl\{2,\frac{\gamma}{s_{0}},\frac{2}{\alpha_{0}}\Bigr\}$
and $\beta_{n}=D\log n$ for some constant
$D>\frac{1+4(d+3)\kappa}{\tau_{0}},$
where $d$ is the dimension of $X$. First,
(8.7) $\begin{split}\sum_{k=1}^{n-1}\sum_{j=k+1}^{k+\beta_{n}}\operatorname{Cov}(\mathbf{1}_{B(y,r_{k}(y))},\mathbf{1}_{B(y,r_{j}(y))}\circ T^{j-k})&\geq-\sum_{k=1}^{n-1}\sum_{j=k+1}^{k+\beta_{n}}M_{k}M_{j}\\ &\gtrsim-\sum_{k=1}^{n-1}M_{k}^{2}\log n.\end{split}$
Recall that $M_{n}\gtrsim n^{-\gamma}$ and thus that $\mathcal{M}_{n}\gtrsim n^{1-\gamma}$. As a result, for any $s\in(0,1-\gamma)$, we have that
$\sum_{k=1}^{n^{s}}M_{k}^{2}\log n\leq n^{s}\log n=o(\mathcal{M}_{n}).$
For $k\geq n^{s}$, we have that $\log k\geq s\log n$ and therefore
$\sum_{k=n^{s}+1}^{n-1}M_{k}^{2}\log n\leq\frac{1}{s}\sum_{k=n^{s}+1}^{n-1}M_{k}^{2}\log k=o(\mathcal{M}_{n})$
since $M_{k}\lesssim(\log k)^{-\upsilon}$ for some $\upsilon>1$. By (8.7), we therefore obtain that
(8.8) $\sum_{k=1}^{n-1}\sum_{j=k+1}^{k+\beta_{n}}\operatorname{Cov}(\mathbf{1}_{B(y,r_{k}(y))},\mathbf{1}_{B(y,r_{j}(y))}\circ T^{j-k})\geq-o(\mathcal{M}_{n}).$
We leverage Condition D.2 in order to estimate the other part of the double sum. Let $\varepsilon_{n}=n^{-\kappa}$ (see (8.6)) and note that $r_{k}(y)\gtrsim M_{k}^{1/s_{0}}\gtrsim k^{-\gamma/s_{0}}$ by Condition M.1. Therefore, $\varepsilon_{n}\leq r_{k}(y)\leq\rho_{0}$ for all sufficiently large $n$, all $k\leq n$ and $y\in\operatorname{supp}\mu$. Thus, we may apply Lemma 4.4 with $\varepsilon=\varepsilon_{n}$ and $r=r_{k}$ to obtain $\alpha$-Hölder continuous functions $g_{n,k}\colon X\times X\to\mathbb{R}$ such that $\|g_{n,k}\|_{\alpha}\lesssim\varepsilon_{n}^{-1}$ and
(8.9) $\mathbf{1}_{B(y,r_{k}(y)-\varepsilon_{n})}(x)\leq g_{n,k}(x,y)\leq\mathbf{1}_{B(y,r_{k}(y))}(x)\quad\forall x,y\in X,$
(8.10) $\biggl|\int_{X}g_{n,k}(x,y)\mathop{}\!d\mu(x)-M_{k}\biggr|\lesssim\varepsilon_{n}^{\alpha_{0}}\lesssim n^{-\alpha_{0}\kappa}\quad\forall y\in\operatorname{supp}\mu.$
Thus,
$\operatorname{Cov}(\mathbf{1}_{B(y,r_{k}(y))},\mathbf{1}_{B(y,r_{j}(y))}\circ T^{j-k})\\ \geq\int_{X}g_{n,k}(y,x)g_{n,j}(y,T^{j-k}x)\mathop{}\!d\mu(x)-M_{k}M_{j}$
and, by the first part of Lemma 4.5 applied to the function $g(x,T^{j-k}x)=g_{n,k}(y,x)g_{n,j}(y,T^{j-k}x)$,
$\int_{X}g_{n,k}(y,x)g_{n,j}(y,T^{j-k}x)\mathop{}\!d\mu(x)-M_{k}M_{j}\\ \gtrsim-n^{-\alpha_{0}\kappa}-n^{-\kappa}-n^{4(d+3)\kappa/\alpha}e^{-\tau_{0}(j-k)}$
for all $y\in\operatorname{supp}\mu$ since (8.10) implies that
$\begin{split}\iint g_{n,k}(y,x)g_{n,j}(y,z)\mathop{}\!d\mu(x)\mathop{}\!d\mu(z)&=(M_{k}+\mathcal{O}(n^{-\alpha_{0}\kappa}))(M_{j}+\mathcal{O}(n^{-\alpha_{0}\kappa}))\\ &=M_{k}M_{j}+\mathcal{O}(n^{-\alpha_{0}\kappa})\end{split}$
for all $y\in\operatorname{supp}\mu$. Hence,
(8.11) $\begin{split}\sum_{k=1}^{n-1}\sum_{j=k+\beta_{n}+1}^{n}&\operatorname{Cov}(\mathbf{1}_{B(y,r_{k}(y))},\mathbf{1}_{B(y,r_{j}(y))}\circ T^{j-k})\\ &\gtrsim-\sum_{k=1}^{n-1}\sum_{j=k+\beta_{n}+1}^{n}\bigl(n^{-\alpha_{0}\kappa}+n^{-\kappa}+n^{4(d+3)\kappa/\alpha}e^{-\tau_{0}(j-k)}\bigr)\\ &\gtrsim n^{2-\alpha_{0}\kappa}+n^{2-\kappa}+n^{1+4(d+3)\kappa/\alpha}e^{-\tau_{0}\beta_{n}}\\ &=-o(\mathcal{M}_{n})\end{split}$
for all $y\in\operatorname{supp}\mu$ since the terms in the third line converge to 0 by the choice of the constants $\kappa$, and $D$ in $\beta_{n}=D\log n$. We stress that the multiplicative constants appearing in (8.11) do not depend on $y$. Therefore, by combining (8.5), (8.8) and (8.11), we conclude that, for all $y\in\operatorname{supp}\mu$,
$s_{n}^{2}(y)\geq\mathcal{M}_{n}-o(\mathcal{M}_{n}),$
where the constant appearing in $o(\mathcal{M}_{n})$ is independent of $y$. This implies the main statement. Finally, by the condition $M_{n}\gtrsim n^{-\gamma}$, we obtain that $s_{n}^{2}(y)\gtrsim\mathcal{M}_{n}\gtrsim n^{1-\gamma}$ for all $y\in\operatorname{supp}\mu$. ∎
8.2. Proof of variance estimates for implicit radii
In this section, we prove the results in §3.1. Recall that we may assume without loss of generality that $r_{n}(x)\leq\rho_{0}$ for all $n\geq 1$ and $x\in\operatorname{supp}\mu$, where $\rho_{0}>0$ appears in Condition M.2.
Proof of Lemma 3.2.
Let $D>1/\eta_{0}$, where $\eta_{0}$ is the constant appearing in Lemma 4.2, $\beta_{n}=D\log n$ and
$\sigma_{n,\beta_{n}}^{2}=\int_{X}\biggl(\sum_{k=1}^{n}\tilde{\mathbf{1}}_{E_{k+\beta_{n}}}\biggr)^{2}\mathop{}\!d\mu.$
By shifting time, we have that
$\sum_{k=1}^{n}\tilde{\mathbf{1}}_{E_{k+\beta_{n}}}=\sum_{k=1}^{n}\tilde{\mathbf{1}}_{E_{k}}+\mathcal{O}(\beta_{n}).$
By squaring and integrating both sides of the equation, we obtain that
(8.12) $|\sigma_{n,\beta_{n}}^{2}-\sigma_{n}^{2}|\lesssim\beta_{n}\int_{X}\biggl|\sum_{k=1}^{n}\tilde{\mathbf{1}}_{E_{k}}\biggr|\mathop{}\!d\mu+\beta_{n}^{2}\lesssim\beta_{n}\sigma_{n}+\beta_{n}^{2},$
and hence that $\sigma_{n,\beta_{n}}^{2}\sim\sigma_{n}^{2}$.
It suffices therefore to show that
(8.13) $\liminf_{n\to\infty}\frac{\sigma_{n,\beta_{n}}^{2}}{\mathcal{E}_{n}}\geq 1$
for the first part of the lemma. Expanding the expression for $\sigma_{n,\beta_{n}}^{2}$, we obtain that
$\sigma_{n,\beta_{n}}^{2}=\sum_{k=1}^{n}\mu(E_{k+\beta_{n}})-\mu(E_{k+\beta_{n}})^{2}+\sum_{k=1}^{n-1}\sum_{j=k+1}^{n}\operatorname{Cov}(\mathbf{1}_{E_{k+\beta_{n}}},\mathbf{1}_{E_{j+\beta_{n}}})$
It is straightforward to see that
$\sum_{k=1}^{n}\mu(E_{k+\beta_{n}})-\mu(E_{k+\beta_{n}})^{2}=\mathcal{E}_{n}-o(\mathcal{E}_{n})+\mathcal{O}(\beta_{n})=\mathcal{E}_{n}-o(\mathcal{E}_{n})$
by shifting time, using that $\mu(E_{k})$ satisfies (8.1) and the assumption that $\beta_{n}\lesssim\log n=o(\mathcal{E}_{n})$. We therefore focus on the double sum, which we split into two parts: when $j\leq k+\beta_{n}$ and when $j>k+\beta_{n}$.
In the first case, we have that
$\begin{split}\sum_{k=1}^{n-1}\sum_{j=k+1}^{k+\beta_{n}}\operatorname{Cov}(\mathbf{1}_{E_{k+\beta_{n}}},\mathbf{1}_{E_{j+\beta_{n}}})&\geq-\sum_{k=1}^{n-1}\sum_{j=k+1}^{k+\beta_{n}}\mu(E_{k+\beta_{n}})\mu(E_{j+\beta_{n}})\\ &\gtrsim-\beta_{n}\sum_{k=1}^{n-1}\mu(E_{k+\beta_{n}})^{2},\end{split}$
where we have used that $\mu(E_{k})\lesssim\mu(E_{j})$ for $k\leq j$. By shifting time and splitting the sum up into $k\leq n^{s}$ and $k\geq n^{s}$ for some $s\in(0,1-\gamma)$, we obtain that
$\begin{split}\beta_{n}\sum_{k=1}^{n-1}\mu(E_{k+\beta_{n}})^{2}&\lesssim\beta_{n}^{2}+\beta_{n}\sum_{k=1}^{n-1}\mu(E_{k})^{2}\\ &\lesssim\beta_{n}^{2}+\beta_{n}n^{s}+(\log n)\sum_{k=n^{s}}^{n-1}\mu(E_{k})^{2}\\ &\lesssim\beta_{n}^{2}+\beta_{n}n^{s}+\sum_{k=n^{s}}^{n-1}\mu(E_{k})^{2}\log k\\ &=o(\mathcal{E}_{n}),\end{split}$
where we have used that $\mu(E_{k})$ satisfies (8.1) and $\mu(E_{k})\lesssim(\log k)^{-\upsilon}$, and $\mathcal{E}_{n}\gtrsim n^{1-\gamma}$ and $(\log n)n^{s}=o(n^{1-\gamma})$ by the choice of $s\in(0,1-\gamma)$.
For the second case, when $j>k+\beta_{n}$, we use Lemma 8.1 and obtain, for some $p_{0}>2$ and $\eta,q_{0}>0$, that
$\begin{split}\sum_{k=1}^{n-1}\sum_{j=k+\beta_{n}+1}^{n}&\operatorname{Cov}(\mathbf{1}_{E_{k+\beta_{n}}},\mathbf{1}_{E_{j+\beta_{n}}})\\ &\gtrsim-\sum_{k=1}^{n-1}\sum_{j=k+\beta_{n}+1}^{n}\bigl(n^{-p_{0}}+n^{q_{0}}(e^{-\eta(k+\beta_{n})}+e^{-\eta(j-k)})\bigr)\\ &\gtrsim-n^{2-p_{0}}-n^{1+q_{0}}\sum_{k=1}^{n-1}\bigl(e^{-\eta(k+\beta_{n})}+e^{-\eta\beta_{n}}\bigr)\\ &\gtrsim-n^{2-p_{0}}-n^{2+q_{0}}e^{-\eta\beta_{n}}\\ &=-n^{2-p_{0}}-n^{2+q_{0}-D\eta}\\ &=o(\mathcal{E}_{n})\end{split}$
since $p_{0}>2$ and by choosing $D>(2+q_{0})/\eta$. In conclusion, $\sigma_{n,\beta_{n}}^{2}\geq\mathcal{E}_{n}-o(\mathcal{E}_{n})$, which implies (8.13). Since $\mu(E_{n})\gtrsim n^{-\gamma}$, we obtain as a consequence that $\sigma_{n}^{2}\gtrsim\mathcal{E}_{n}\gtrsim n^{1-\gamma}$ for all sufficiently large $n$. ∎
Proof of Lemma 3.3.
We show that
(8.14) $\liminf_{n\to\infty}\int_{X}\frac{s_{n}^{2}(y)}{\sigma_{n}^{2}}\mathop{}\!d\mu(y)\geq 1.$
That the limit superior holds with the reversed inequality is established in a similar manner with only minor modifications and we therefore omit its proof.
Let
(8.15) $\kappa>\max\Bigl\{4,\frac{4}{\alpha_{0}},\frac{2\gamma}{s_{0}}\Bigr\}$
and $\beta_{n}=D\log n$ for some constant $D$ such that $D>\eta_{0}^{-1}$, where $\eta_{0}$ appears in Lemma 4.2, and such that
$\lim_{n\to\infty}n^{2(d+3)\kappa}P_{r}(n)e^{-\tau_{r}\beta_{n}}=0,$
where $\tau_{r}$ and $P_{r}$ appear in Condition D.3. As it will be used later, we recall (8.12) in the proof of Lemma 3.2, namely that
(8.16) $|\sigma_{n}^{2}-\sigma_{n,\beta_{n}}^{2}|\lesssim\beta_{n}\sigma_{n}+\beta_{n}^{2},$
where
(8.17) $\sigma_{n,\beta_{n}}^{2}=\int_{X}\biggl(\sum_{k=1}^{n}\mathbf{1}_{E_{k+\beta_{n}}}\biggr)^{2}\mathop{}\!d\mu-\biggl(\sum_{k=1}^{n}\mu(E_{k+\beta_{n}})\biggr)^{2}.$
Obtain $\mathcal{Q}_{n}$, $Y_{n}$ and $\{\rho_{n,Q}\}_{Q\in\mathcal{Q}_{n}}$ from Lemma 4.3 with $\kappa_{0}=\kappa$. Let $\varepsilon_{n}=n^{-\kappa/2}$ and notice that $r_{n}\gtrsim n^{-\gamma/s_{0}}$. Hence $\varepsilon_{n}\leq r_{n}$ for sufficiently large $n$ and we may obtain Lipschitz continuous functions $g_{n,k}\colon X\times X\to[0,1]$ from Lemma 4.4 applied to $\varepsilon=\varepsilon_{n}$ and $r=r_{n}$. That is, $\|g_{n,k}\|_{\textup{Lip}}\lesssim\varepsilon_{n}^{-1}\lesssim n^{\kappa/2}$ and $g_{n,k}$ satisfies (4.3) and (4.4). Then $\mathbf{1}_{E_{k}}(x)\leq g_{n,k}(x,T^{k}x)$ and, for each $Q\in\mathcal{Q}_{n}$ and all $x,y_{Q}\in Q$,
(8.18) $\biggl|\biggl(\sum_{k=1+\beta_{n}}^{n+\beta_{n}}g_{n,k}(x,T^{k}x)\biggr)^{2}-\biggl(\sum_{k=1+\beta_{n}}^{n+\beta_{n}}g_{n,k}(x_{Q},T^{k}x)\biggr)^{2}\biggr|\lesssim n^{2}\varepsilon_{n}^{-1}\lvert\mathcal{Q}_{n}\rvert$,
where we notice that $n^{2}\varepsilon_{n}^{-1}\lvert\mathcal{Q}_{n}\rvert\lesssim n^{2-\kappa/2}\lesssim 1$ by the choice of $\kappa$ (see (8.15)). Hence,
$\int_{X}\biggl(\sum_{k=1}^{n}\mathbf{1}_{E_{k+\beta_{n}}}\biggr)^{2}\mathop{}\!d\mu\leq\sum_{Q\in\mathcal{Q}_{n}}\int_{X}\mathbf{1}_{Q}(x)\biggl(\sum_{k=1}^{n}g_{n,k+\beta_{n}}(x_{Q},T^{k+\beta_{n}}x)\biggr)^{2}\mathop{}\!d\mu(x)\\ +\mathcal{O}(1+n^{2}\mu(Y_{n})),$
where we have used that $X\setminus\bigcup_{Q\in\mathcal{Q}_{n}}Q=Y_{n}$. Recall that $\mu(Y_{n})\lesssim n^{-\kappa}$, and therefore that $n^{2}\mu(Y_{n})\lesssim 1$ since $\kappa>2$. Furthermore, by employing the densities $\{\rho_{Q,n}\}_{Q\in\mathcal{Q}_{n}}$ and their property that $\|\rho_{Q,n}-\mu(Q)^{-1}\mathbf{1}_{Q}\|_{L^{1}}\lesssim n^{-\kappa}$, we obtain that
$\int_{X}\biggl(\sum_{k=1}^{n}\mathbf{1}_{E_{k+\beta_{n}}}\biggr)^{2}\mathop{}\!d\mu\\ \leq\sum_{Q\in\mathcal{Q}_{n}}\mu(Q)\int_{X}\rho_{Q,n}\biggl(\sum_{k=1}^{n}g_{n,k+\beta_{n}}(x_{Q},T^{k+\beta_{n}}x)\biggr)^{2}\mathop{}\!d\mu(x)+\mathcal{O}(1+n^{2-\kappa}).$
Again, since $\kappa>2$, we have that $n^{2-\kappa}\lesssim 1$. Applying Condition D.3 to the right-hand side, one obtains that
(8.19) $\int_{X}\biggl(\sum_{k=1}^{n}\mathbf{1}_{E_{k+\beta_{n}}}\biggr)^{2}\mathop{}\!d\mu\leq\sum_{Q\in\mathcal{Q}_{n}}\mu(Q)\int_{X}\biggl(\sum_{k=1}^{n}g_{n,k+\beta_{n}}(x_{Q},T^{k}x)\biggr)^{2}\mathop{}\!d\mu(x)\\ +\mathcal{O}(1+n^{2-\kappa}+n^{2(d+2)\kappa}\varepsilon_{n}^{-2}P_{r}(n)e^{-\tau_{r}\beta_{n}}),$
where we have used that $\mu(\rho_{Q,n})=1+\mathcal{O}(n^{-\kappa})$ and $\|\rho_{Q,n}\|_{\alpha}\lesssim n^{2(d+2)\kappa}$. The choice of $\kappa$ and $D$ in $\beta_{n}$ ensures that the error term is bounded above by a constant. Using (8.18) again, we can write
$\biggl|\mu(Q)\biggl(\sum_{k=1}^{n}g_{n,k+\beta_{n}}(x_{Q},T^{k}x)\biggr)^{2}-\int_{Q}\biggl(\sum_{k=1}^{n}g_{n,k+\beta_{n}}(y,T^{k}x)\biggr)^{2}\mathop{}\!d\mu(y)\biggr|\lesssim 1.$
In view of (8.19), we obtain that
$\int_{X}\biggl(\sum_{k=1}^{n}\mathbf{1}_{E_{k+\beta_{n}}}\biggr)^{2}\mathop{}\!d\mu\\ \leq\sum_{Q\in\mathcal{Q}_{n}}\mu(Q)\int_{Q}\int_{X}\biggl(\sum_{k=1}^{n}g_{n,k+\beta_{n}}(y,T^{k}x)\biggr)^{2}\mathop{}\!d\mu(x)\mathop{}\!d\mu(y)+\mathcal{O}(1).$
Using again that $X\setminus\bigcup_{Q\in\mathcal{Q}_{n}}Q=Y_{n}$, where $\mu(Y_{n})\lesssim n^{-\kappa}$, we obtain that
$\int_{X}\biggl(\sum_{k=1}^{n}\mathbf{1}_{E_{k+\beta_{n}}}\biggr)^{2}\mathop{}\!d\mu\leq\iint_{X\times X}\biggl(\sum_{k=1}^{n}g_{n,k+\beta_{n}}(y,T^{k}x)\biggr)^{2}\mathop{}\!d\mu(x)\mathop{}\!d\mu(y)+\mathcal{O}(1).$
By the construction of $g_{n,k+\beta_{n}}$, we have that
$\biggl|\iint_{X\times X}\biggl(\sum_{k=1}^{n}g_{n,k+\beta_{n}}(T^{k}\cdot,y)\biggr)^{2}\mathop{}\!d\mu(x)\mathop{}\!d\mu(y)\\ -\iint_{X\times X}\biggl(\sum_{k=1}^{n}\mathbf{1}_{B(y,r_{k+\beta_{n}}(y))}\circ T^{k}\biggr)^{2}\mathop{}\!d\mu(x)\mathop{}\!d\mu(y)\biggr|\\ \lesssim n^{2}\varepsilon_{n}^{\alpha_{0}}\lesssim 1$
since $n^{2}\varepsilon_{n}^{\alpha_{0}}\lesssim n^{2-\alpha_{0}\kappa/2}\lesssim 1$ by $\kappa>4/\alpha_{0}$. Thus, by combining the previous two equations, we obtain that
$\int\biggl(\sum_{k=1}^{n}\mathbf{1}_{E_{k+\beta_{n}}}\biggr)^{2}\mathop{}\!d\mu-\iint\biggl(\sum_{k=1}^{n}\mathbf{1}_{B(y,r_{k+\beta_{n}}(y))}(T^{k}x)\biggr)^{2}\mathop{}\!d\mu(x)\mathop{}\!d\mu(y)\lesssim 1.$
Now, since
$\begin{split}\biggl|\biggl(\sum_{k=1}^{n}\mu(E_{k+\beta_{n}})\biggr)^{2}-\biggl(\sum_{k=1}^{n}M_{k+\beta_{n}}\biggr)^{2}\biggr|&\lesssim n\sum_{k=1}^{n}\bigl|\mu(E_{k+\beta_{n}})-M_{k+\beta_{n}}\bigr|\\ &\lesssim ne^{-\eta_{0}\beta_{n}}\\ &\lesssim 1\end{split}$
by Lemma 4.2 and the choice of $D$ in $\beta_{n}=D\log n$, we obtain, from (8.17), that
$\sigma_{n,\beta_{n}}^{2}\leq\iint\biggl(\sum_{k=1}^{n}\mathbf{1}_{B(y,r_{k+\beta_{n}}(y))}(T^{k}x)\biggr)^{2}\mathop{}\!d\mu(x)\mathop{}\!d\mu(y)-\biggl(\sum_{k=1}^{n}M_{k+\beta_{n}}\biggr)^{2}+\mathcal{O}(1).$
Mimicking the notation of $\sigma_{n}^{2}$ and $\sigma_{n,\beta_{n}}^{2}$, we can write this equation as
(8.20) $\sigma_{n,\beta_{n}}^{2}-\int_{X}s_{n,\beta_{n}}^{2}(y)\mathop{}\!d\mu(y)\lesssim 1,$
where
$s_{n,\beta_{n}}^{2}(y)=\int_{X}\biggl(\sum_{k=1}^{n}\mathbf{1}_{B(y,r_{k+\beta_{n}}(y))}(T^{k}(x))\biggr)^{2}\mathop{}\!d\mu(x)-\biggl(\sum_{k=1}^{n}M_{k+\beta_{n}}\biggr)^{2}.$
In a similar manner to how (8.16) was obtained, one establishes that
$|s_{n,\beta_{n}}^{2}(y)-s_{n}^{2}(y)|\lesssim\beta_{n}s_{n}(y)+\beta_{n}^{2}$
for all $y\in X$. We therefore have, by (8.16) and (8.20), that
(8.21) $\sigma_{n}^{2}-\int_{X}s_{n}^{2}(y)\mathop{}\!d\mu(y)\lesssim\beta_{n}\int_{X}s_{n}(y)\mathop{}\!d\mu(y)+\beta_{n}\sigma_{n}+\beta_{n}^{2}.$
From this equation, we can conclude (8.14), namely that
$\liminf_{n\to\infty}\frac{\int s_{n}^{2}(y)\mathop{}\!d\mu(y)}{\sigma_{n}^{2}}\geq 1.$
Indeed, by Lemma 3.1, we have, for all sufficiently large $n$ and all $y\in\operatorname{supp}\mu$, that $s_{n}(y)\geq\frac{1}{2}\sqrt{\mathcal{M}_{n}}$. Hence,
$\lim_{n\to\infty}\frac{\beta_{n}\int s_{n}(y)\mathop{}\!d\mu(y)}{\int s_{n}^{2}(y)\mathop{}\!d\mu(y)}=0$
by an application of Jensen’s inequality:
$0\leq\limsup_{n\to\infty}\frac{\beta_{n}\int s_{n}(y)\mathop{}\!d\mu(y)}{\int s_{n}^{2}(y)\mathop{}\!d\mu(y)}\leq\limsup_{n\to\infty}\frac{D\log n}{\int s_{n}(y)\mathop{}\!d\mu(y)}\leq\limsup_{n\to\infty}\frac{2D\log n}{\sqrt{\mathcal{M}_{n}}}=0$
since $\mathcal{M}_{n}\gtrsim n^{1-\gamma}$. Furthermore, $\sigma_{n}^{-1}\log n\to 0$ by Lemma 3.2. We can therefore rewrite (8.21) as
$a_{n}(1+b_{n})+c_{n}\geq 1$
for
$a_{n}=\liminf_{n\to\infty}\frac{\int s_{n}^{2}(y)\mathop{}\!d\mu(y)}{\sigma_{n}^{2}}$
and $b_{n},c_{n}\to 0$. Therefore, we obtain that $\liminf_{n\to\infty}a_{n}(1+b_{n})\geq 1$, which implies that $\liminf_{n\to\infty}a_{n}\geq 1$ as desired.
To establish that
$\limsup_{n\to\infty}\frac{\int s_{n}^{2}(y)\mathop{}\!d\mu(y)}{\sigma_{n}^{2}}\leq 1,$
approximate the functions $\mathbf{1}_{E_{k}}$ from below using Lemma 4.4. The desired result then follows analogously. ∎
Proof of Lemma 3.4.
We begin by noticing that, by Lemma 3.3,
$\lim_{n\to\infty}\int_{X}\frac{s_{n}^{2}}{\sigma_{n}^{2}}\mathop{}\!d\mu=1.$
Therefore, to establish the lemma, it suffices to show that
$\lim_{n\to\infty}\mu\biggl\{y:\Bigl|\frac{s_{n}^{2}(y)}{\sigma_{n}^{2}}-\int_{X}\frac{s_{n}^{2}}{\sigma_{n}^{2}}\mathop{}\!d\mu\Bigr|>\varepsilon\biggr\}=0$
for all $\varepsilon>0$. By Chebyshev’s inequality, we have that
$\lim_{n\to\infty}\mu\biggl\{y:\Bigl|\frac{s_{n}^{2}(y)}{\sigma_{n}^{2}}-\int_{X}\frac{s_{n}^{2}}{\sigma_{n}^{2}}\mathop{}\!d\mu\Bigr|>\varepsilon\biggr\}\leq\frac{1}{\varepsilon^{2}}\cdot\frac{1}{\sigma_{n}^{4}}\operatorname{Var}(s_{n}^{2}).$
Consequently, establishing that $\operatorname{Var}(s_{n}^{2})=o(\sigma_{n}^{4})$ implies the lemma.
By the proof of Lemma 3.1, in particular (8.11) and the corresponding upper bound that one obtains in a similar fashion, we have that
$s_{n}^{2}(y)=\mathcal{M}_{n}+o(\mathcal{M}_{n}^{\varepsilon_{0}})+\sum_{k=1}^{n-1}\sum_{j=k+1}^{k+\beta_{n}}\mu\bigl(B(y,r_{k}(y))\cap T^{-(j-k)}B(y,r_{j}(y))\bigr)$
some $\varepsilon_{0}\in(0,1)$ and where $\beta_{n}=D\log n$ for some large constant $D>0$. For notational convenience, let
$A_{k,j}(y)=\mu\bigl(B(y,r_{k}(y))\cap T^{-(j-k)}B(y,r_{j}(y))\bigr)$
for $k\leq j$ and notice for later that $\mu(A_{k,j}(y))\leq\mu(B(y,r_{k}(y)))=M_{k}$. We can then write
$s_{n}^{4}(y)=\mathcal{M}_{n}^{2}+o(\beta_{n}\mathcal{M}_{n}^{1+\varepsilon_{0}})+2\mathcal{M}_{n}\sum_{k=1}^{n-1}\sum_{j=k+1}^{k+\beta_{n}}A_{k,j}(y)\\ +\sum_{k,l=1}^{n-1}\sum_{j=k+1}^{k+\beta_{n}}\sum_{i=l+1}^{l+\beta_{n}}A_{k,j}(y)A_{l,i}(y)$
and
$\biggl(\int_{X}s_{n}^{2}\mathop{}\!d\mu\biggr)^{2}=\mathcal{M}_{n}^{2}+o(\beta_{n}\mathcal{M}_{n}^{1+\varepsilon_{0}})+2\mathcal{M}_{n}\sum_{k=1}^{n-1}\sum_{j=k+1}^{k+\beta_{n}}\int_{X}A_{k,j}(y)\mathop{}\!d\mu(y)\\ +\sum_{k,l=1}^{n-1}\sum_{j=k+1}^{k+\beta_{n}}\sum_{i=l+1}^{l+\beta_{n}}\int_{X}A_{k,j}(y)\mathop{}\!d\mu(y)\int_{X}A_{l,i}(z)\mathop{}\!d\mu(z).$
Hence,
$\operatorname{Var}(s_{n}^{2})=\sum_{k,l,j,i}\int_{X}A_{k,j}(y)A_{l,i}(y)\mathop{}\!d\mu(y)-\int_{X}A_{k,j}(y)\mathop{}\!d\mu(y)\int_{X}A_{l,i}(z)\mathop{}\!d\mu(z)\\ +o(\beta_{n}\mathcal{M}_{n}^{1+\varepsilon_{0}}),$
where the sum is taken over all $1\leq k,l\leq n-1$, and all $k<j\leq k+\beta_{n}$ and $l<i\leq l+\beta_{n}$. Using the trivial estimate $\mu(A_{l,i}(y))\leq\mu(B(y,r_{l}(y)))=M_{l}$, we obtain that
(8.22) $\operatorname{Var}(s_{n}^{2})\lesssim\beta_{n}\mathcal{M}_{n}\sum_{k=1}^{n-1}\sum_{j=k+1}^{k+\beta_{n}}\int_{X}A_{k,j}(y)\mathop{}\!d\mu(y)+o(\beta_{n}\mathcal{M}_{n}^{1+\varepsilon_{0}}).$
For $k\leq n^{s}$, where $s\in(0,1-\gamma)$, we have that
$\beta_{n}\mathcal{M}_{n}\sum_{k=1}^{n^{s}}\sum_{j=k+1}^{k+\beta_{n}}\int_{X}A_{k,j}(y)\mathop{}\!d\mu(y)\lesssim n^{s}(\log n)^{2}\mathcal{M}_{n}=o(\sigma_{n}^{4})$
since $\sigma_{n}^{4}\gtrsim\mathcal{E}_{n}^{2}\gtrsim n^{1-\gamma}\mathcal{M}_{n}$ by Lemmas 3.2 and 4.2. For $k\geq n^{s}$, we have that $\beta_{n}=D\log n\lesssim\log k$. Hence, for some constant $C>0$,
$\begin{split}\beta_{n}\mathcal{M}_{n}\sum_{k=n^{s}}^{n-1}\sum_{j=k+1}^{k+\beta_{n}}\int_{X}A_{k,j}(y)\mathop{}\!d\mu(y)&\lesssim\beta_{n}\mathcal{M}_{n}\sum_{k=n^{s}}^{n-1}\sum_{j=k+1}^{k+C\log k}\int_{X}A_{k,j}(y)\mathop{}\!d\mu(y)\\ &=o(\beta_{n}\mathcal{M}_{n}\sigma_{n}^{2-\delta})\\ &=o(\sigma_{n}^{4})\end{split}$
by employing the short returns estimate (3.1) in the second line, and $\sigma_{n}^{2}\gtrsim\mathcal{M}_{n}$ and $\log n\lesssim\log\sigma_{n}$. By (8.22), we have therefore established that $\operatorname{Var}(s_{n}^{2})=o(\sigma_{n}^{4})$, which concludes the proof. ∎
8.3. Proof of variance estimates for explicit radii
In this section, we prove the results in §3.2. Recall that we assume that $r_{n}\to 0$.
Proof of Lemma 3.5.
Expanding the expression for the variance, we obtain
$\hat{\sigma}_{n}^{2}=\sum_{k=1}^{n}A_{k}+\sum_{k=1}^{n-1}\sum_{j=k+1}^{n}B_{k,j},$
where
$A_{k}=\mu(\hat{E}_{k})-2\int_{X}\mathbf{1}_{B(x,r_{k})}(T^{k}x)\mu(B(x,r_{k}))\mathop{}\!d\mu(x)+\int_{X}\mu(B(x,r_{k}))^{2}\mathop{}\!d\mu(x)$
and
$B_{k}=\mu(\hat{E}_{k}\cap\hat{E}_{j})-\int_{X}\mathbf{1}_{B(x,r_{k})}(T^{k}x)\mu(B(x,r_{j}))\mathop{}\!d\mu(x)\\ -\int_{X}\mathbf{1}_{B(x,r_{j})}(T^{j}x)\mu(B(x,r_{k}))\mathop{}\!d\mu(x)\\ +\int_{X}\mu(B(x,r_{k}))\mu(B(x,r_{j}))\mathop{}\!d\mu(x).$
As in the proof of Lemma 3.2, we consider
$\hat{\sigma}_{n,\beta_{n}}^{2}=\sum_{k=1}^{n}A_{k+\beta_{n}}+\sum_{k=1}^{n-1}\sum_{j=k+1}^{n}B_{k+\beta_{n},j+\beta_{n}}$
for $\beta_{n}=D\log n$ for a constant $D>0$ that we will choose sufficiently large. Similar to the explicit radii setting (see (8.12)), we obtain that
$|\hat{\sigma}_{n,\beta_{n}}^{2}-\hat{\sigma}_{n}^{2}|\lesssim\beta_{n}\hat{\sigma}_{n}+\beta_{n}^{2}.$
Since $\hat{\sigma}_{n}^{-1}\log n\to 0$ by Lemma 3.5, it suffices to show that $\hat{\sigma}_{n,\beta_{n}}^{2}=\mathcal{E}_{n}-o(\mathcal{E}_{n})$.
We begin by estimating $A_{k+\beta_{n}}$. Let
$\kappa>\max\Bigl\{4,\frac{2}{\alpha_{0}},\frac{2\gamma}{s_{0}}\Bigr\}$
and $\varepsilon_{n}=n^{-\kappa/2}$. Obtain the functions $g_{n,k}\colon X\times X\to[0,1]$ by applying Lemma 4.4 to $r=r_{n}$ and $\varepsilon=\varepsilon_{n}$. We therefore have that
$\int\mathbf{1}_{B(x,r_{k})}(T^{k}x)\mu(B(x,r_{k}))\mathop{}\!d\mu(x)\leq\iint g_{n,k}(x,T^{k}x)g_{n,k}(x,y)\mathop{}\!d\mu(y)\mathop{}\!d\mu(x).$
Applying Lemma 4.5 to the function $g=g_{n,k}(x,T^{k}x)\int g_{n,k}(x,y)\mathop{}\!d\mu(y)$, we obtain that
$\begin{split}\iint g_{n,k}(x,T^{k}x)g_{n,k}(x,y)&\mathop{}\!d\mu(y)\mathop{}\!d\mu(x)\\ &\leq\iiint g_{n,k}(x,z)g_{n,k}(x,y)\mathop{}\!d\mu(y)\mathop{}\!d\mu(z)\mathop{}\!d\mu(x)\\ &\quad+\mathcal{O}\bigl(n^{-\kappa/2}+n^{2(d+3)\kappa/\alpha}e^{-\tau_{0}k}\bigr).\end{split}$
Using that
$\Bigl|\int_{X}g_{n,k}(x,y)\mathop{}\!d\mu(y)-\mu(B(x,r_{k})\Bigr|\lesssim\varepsilon_{n}^{\alpha_{0}}\lesssim n^{-\alpha_{0}\kappa/2}$
for $\mu$-a.e. $x$, we obtain that
$\iiint g_{n,k}(x,z)g_{n,k}(x,y)\mathop{}\!d\mu(y)\mathop{}\!d\mu(z)\mathop{}\!d\mu(x)\\ \leq\int_{X}\mu(B(x,r_{k}))^{2}\mathop{}\!d\mu(x)+\mathcal{O}(n^{-\alpha_{0}\kappa/2}).$
We therefore obtain that
(8.23) $\int_{X}\mathbf{1}_{B(x,r_{k})}(T^{k}x)\mu(B(x,r_{k}))\mathop{}\!d\mu(x)\\ \leq\int_{X}\mu(B(x,r_{k}))^{2}\mathop{}\!d\mu(x)+\mathcal{O}\bigl(n^{-\alpha_{0}\kappa/2}+n^{-\kappa/2}+n^{2(d+3)\kappa/\alpha}e^{-\tau_{r}k}\bigr).$
Hence,
$\sum_{k=1}^{n}A_{k+\beta_{n}}\geq\sum_{k=1}^{n}\mu(\hat{E}_{k+\beta_{n}})-\int_{X}\mu(B(x,r_{k+\beta_{n}}))^{2}\mathop{}\!d\mu(x)\\ +\mathcal{O}\bigl(n^{1-\alpha_{0}\kappa/2}+n^{1-\kappa/2}+n^{2(d+3)\kappa/\alpha}e^{-\tau_{r}\beta_{n}}\bigr).$
By the choices of $\kappa$ and $D$ in $\beta_{n}=D\log n$, we have that the error term converges to 0 as $n\to\infty$. Furthermore, by shifting time, and using that $\mu(B(x,r))\lesssim r^{s_{0}}$ and $\hat{M}_{k}=\int\mu(B(x,r_{k}))\mathop{}\!d\mu(x)$, we obtain that
$\sum_{k=1}^{n}A_{k+\beta_{n}}\geq\sum_{k=1}^{n}\bigl(\mu(\hat{E}_{k})-r_{k}^{s_{0}}\hat{M}_{k}\bigr)-\beta_{n}+\mathcal{O}(1).$
Since $r_{k}\to 0$ and $|\mu(\hat{E}_{k})-\hat{M}_{k}|\lesssim e^{-\eta_{0}k}$ by Lemma 4.2, we obtain that
(8.24) $\sum_{k=1}^{n}A_{k+\beta_{n}}\geq\hat{\mathcal{E}}_{n}-o(\hat{\mathcal{E}}_{n}),$
where we have also used that $\beta_{n}=o(\hat{\mathcal{E}}_{n})$.
For the double sum involving $B_{k,j}$, we consider the cases $j\leq k+\beta_{n}$ and $j>k+\beta_{n}$ separately. In the first case, we have that
$\int_{X}\mathbf{1}_{B(x,r_{k})}(T^{k}x)\mu(B(x,r_{j}))\mathop{}\!d\mu(x)\lesssim r_{j}^{s_{0}}\mu(\hat{E}_{k})\lesssim r_{k}^{s_{0}}\mu(\hat{E}_{k})$
by using the monotonicity of $(r_{k})$. Now let $s\in(0,1-\gamma)$. By shifting time again,
(8.25) $\begin{split}\sum_{k=1}^{n-1}\sum_{j=k+1}^{k+\beta_{n}}B_{k+\beta_{n},j+\beta_{n}}&\gtrsim-\beta_{n}-\beta_{n}\sum_{k=1}^{n-1}r_{k}^{s_{0}}\mu(\hat{E}_{k})\\ &\gtrsim-\beta_{n}-\beta_{n}n^{s}-\sum_{k=n^{s}}^{n-1}r_{k}^{s_{0}}\mu(\hat{E}_{k})\log k\\ &=o(\hat{\mathcal{E}}_{n}),\end{split}$
where we have used that $\hat{\mathcal{E}}_{n}\gtrsim n^{1-\gamma}$, $r_{k}\gtrsim(\log k)^{\upsilon/s_{0}}$ for some $\upsilon>1$, and $\beta_{n}n^{s}=o(n^{1-\gamma})$ by the choice of $s$.
When $j>k+\beta_{n}$, notice that, by Lemma 8.1 and a variation of (8.23),
$\sum_{k=1}^{n-1}\sum_{j=k+\beta_{n}+1}^{n}B_{k+\beta_{n},j+\beta_{n}}\\ =\sum_{k=1}^{n-1}\sum_{j=k+\beta_{n}+1}^{n}\Bigl(\mu(\hat{E}_{k+\beta_{n}}\cap\hat{E}_{k+\beta_{n}})-\int_{X}\mu(B(x,r_{k+\beta_{n}}))\mu(B(x,r_{j+\beta_{n}}))\mathop{}\!d\mu(x)\\ \quad+\mathcal{O}\bigl(n^{-\kappa/2}+n^{2(d+3)\kappa}e^{-\tau_{r}(k+\beta_{n})}\bigr)\Bigr).$
By Lemma 8.1, there exist $p_{0}>2$ and $q_{0},\eta>0$ such that
$\Bigl|\mu(\hat{E}_{k+\beta_{n}}\cap\hat{E}_{k+\beta_{n}})-\int_{X}\mu(B(x,r_{k+\beta_{n}}))\mu(B(x,r_{j+\beta_{n}}))\mathop{}\!d\mu(x)\Bigr|\\ \lesssim n^{-p_{0}}+n^{q_{0}}\bigl(e^{-\eta(k+\beta_{n})}+e^{-\eta(j-k)}\bigr).$
Hence, by combining the previous two equations and letting $p=\min\{p_{0},\kappa/2\}$ and $\eta^{\prime}=\min\{\tau_{r},\eta\}$, we obtain that
$\begin{split}\sum_{k=1}^{n-1}\sum_{j=k+\beta_{n}+1}^{n}|B_{k+\beta_{n},j+\beta_{n}}|&\lesssim\sum_{k=1}^{n-1}\sum_{j=k+\beta_{n}+1}^{n}\bigl(n^{-p}+n^{q}\bigl(e^{-\eta^{\prime}(k+\beta_{n})}+e^{-\eta^{\prime}(j-k)}\bigr)\bigr)\\ &\lesssim n^{2-p}+n^{q}\sum_{k=1}^{n-1}\bigl(e^{-\eta^{\prime}(k+\beta_{n})}+e^{-\eta^{\prime}\beta_{n}}\bigr)\\ &\lesssim n^{2-p}+n^{q+1}e^{-\eta^{\prime}\beta_{n}}\end{split}$
for some constant $q>0$. Noticing that $p>2$ and by choosing $D$ in $\beta_{n}=D\log n$ sufficiently large, we therefore obtain that
(8.26) $\sum_{k=1}^{n-1}\sum_{j=k+\beta_{n}+1}^{n}B_{k+\beta_{n},j+\beta_{n}}=\mathcal{O}(1).$
Thus, in combination with (8.25), we obtain that
$\sum_{k=1}^{n-1}\sum_{j=k+1}^{n}B_{k+\beta_{n},j+\beta_{n}}=o(\hat{\mathcal{E}}_{n}).$
In view of (8.24), we obtain that
$\hat{\sigma}_{n,\beta_{n}}^{2}=\hat{\mathcal{E}}_{n}-o(\hat{\mathcal{E}}_{n})$
which implies the result. ∎
We now establish Lemma 3.6. We note that, under the assumptions of the lemma, we have that
(8.27) $\lim_{n\to\infty}\int_{X}\frac{\hat{s}_{n}^{2}(y)}{\hat{\mathcal{M}}_{n}}\mathop{}\!d\mu(y)=1.$
and
(8.28) $\lim_{n\to\infty}\frac{\hat{\sigma}_{n}^{2}}{\hat{\mathcal{E}}_{n}}=1.$
These estimates in turn imply that
(8.29) $\lim_{n\to\infty}\int_{X}\frac{\hat{s}_{n}^{2}(y)}{\hat{\sigma}_{n}^{2}}\mathop{}\!d\mu(y)=1.$
Indeed,
$\lim_{n\to\infty}\int_{X}\frac{\hat{s}_{n}^{2}(y)}{\hat{\sigma}_{n}^{2}}\mathop{}\!d\mu(y)=\lim_{n\to\infty}\frac{\hat{\mathcal{E}}_{n}}{\hat{\sigma}_{n}^{2}}\cdot\lim_{n\to\infty}\frac{\hat{\mathcal{M}}_{n}}{\hat{\mathcal{E}}_{n}}\cdot\lim_{n\to\infty}\int_{X}\frac{\hat{s}_{n}^{2}(y)}{\hat{\mathcal{M}}_{n}}\mathop{}\!d\mu(y)=1,$
where we have used that $\hat{\mathcal{E}}_{n}$ and $\hat{\mathcal{M}}_{n}$ differ only by a constant by Lemma 4.2 and diverge by the assumption on $\hat{M}_{k}$.
We show how to obtain (8.28) since (8.27) follows from a more straightforward yet similar argument. By Lemma 3.5, it suffices to show that
$\limsup_{n\to\infty}\frac{\hat{\sigma}_{n}^{2}}{\hat{\mathcal{E}}_{n}}\leq 1.$
We make use of the following lemma of which we momentarily delay the proof.
Lemma 8.2.
Under the assumptions of Lemma 3.6, we have that
(8.30) $\mu(\hat{E}_{k}\cap\hat{E}_{k+l})\lesssim\mu(\hat{E}_{k})^{1+\delta_{0}}$
for all sufficiently large $k$ and all $l=1,\ldots,(\log k)^{p_{0}}$.
Remark 11 (Short return times).
Compare (8.30) with [34] in which Persson and the author use the weaker condition
$\mu(\hat{E}_{k}\cap\hat{E}_{k+l})\leq C\mu(\hat{E}_{k})^{1+\delta_{0}}+\mu(\hat{E}_{k})\psi(l),$
where $\psi(l)$ is a summable function. The condition is established for some expanding piecewise smooth expanding maps of the interval with Gibbs measure and linear Anosov diffeomorphisms of the $2$-torus. In those cases, $\psi(l)$ is an exponential function. Such an estimate is not enough for the methods in this paper since $\sum_{k=1}^{n}\sum_{l=1}^{D\log k}\mu(\hat{E}_{k})\psi(l)\lesssim\hat{\mathcal{M}}_{n}$, which is not dominated by $\hat{\sigma}_{n}^{2}$.
By the proof of Lemma 3.5, namely (8.24) and (8.26), it is sufficient to show that
(8.31) $\sum_{k=1}^{n-1}\sum_{j=k+1}^{k+\beta_{n}}\mu(\hat{E}_{k}\cap\hat{E}_{j})+\int_{X}\mu(B(x,r_{k}))\mu(B(x,r_{j}))\mathop{}\!d\mu(x)=o(\hat{\mathcal{E}}_{n}),$
where $\beta_{n}=D\log n$ for some $D>0$ sufficiently large. Let $s\in(0,1-\gamma)$ and consider separately the cases $k\leq n^{s}$ and $k\geq n^{s}$.
For $k\leq n^{s}$, we have
$\sum_{k=1}^{n^{s}}\sum_{j=k+1}^{k+\beta_{n}}\mu(\hat{E}_{k}\cap\hat{E}_{j})+\int_{X}\mu(B(x,r_{k}))\mu(B(x,r_{j}))\mathop{}\!d\mu(x)\lesssim n^{s}\log n=o(\hat{\mathcal{E}}_{n})$
since $\hat{\mathcal{E}}_{n}\gtrsim n^{1-\gamma}$. For $k\geq n^{s}$, we have $\beta_{n}\lesssim\log k$. Thus, for some $C>0$, we have that $\beta_{n}\leq C\log k$ and thus, for sufficiently large $n$ (and hence $k$), $C\log k\leq(\log k)^{p_{0}}$. We can therefore apply Lemma 8.2 and Condition M.1 to obtain that
$\begin{split}\sum_{k=n^{s}}^{n-1}\sum_{j=k+1}^{k+\beta_{n}}\mu(\hat{E}_{k}\cap\hat{E}_{j})+\int_{X}&\mu(B(x,r_{k}))\mu(B(x,r_{j}))\mathop{}\!d\mu(x)\\ &\lesssim\sum_{k=n^{s}}^{n-1}\sum_{j=k+1}^{k+C\log k}\hat{M}_{k}^{1+\delta}+r_{k}^{s_{0}}\hat{M}_{k}\\ &\lesssim\sum_{k=n^{s}}^{n-1}(\hat{M}_{k}^{1+\delta}+r_{k}^{s_{0}}\hat{M}_{k})\log k\\ &=o(\hat{\mathcal{E}}_{n}),\end{split}$
where the last line follows from the fact that $\hat{M}_{k}\lesssim(\log k)^{1-\upsilon\delta}$ and $r_{k}\lesssim(\log k)^{1-\upsilon}$ ($s_{0}=d$) for some $\upsilon>\delta^{-1}>1$. We have therefore established (8.31).
Proof of Lemma 8.2.
Let $\kappa>\max\{4\gamma,8\gamma/\alpha_{0}\}$ and $\varepsilon_{k}=k^{-\kappa/2}$, and obtain, for $k\in\mathbb{N}$, functions $g_{k}\colon X\to(0,1)$ by applying Lemma 4.4 to $\varepsilon=\varepsilon_{k}$ and $r=r_{k}$. Then
$\mu(\hat{E}_{k}\cap\hat{E}_{k+l})\leq\int_{X}g_{k}(x,T^{k}x)g_{k+l}(x,T^{k+l}x)\mathop{}\!d\mu(x).$
Using the second part of Lemma 4.5 with $n=k+l\lesssim k$ (since $l\leq(\log k)^{p_{0}}$), we obtain that
$\int_{X}g_{k}(x,T^{k}x)g_{k+l}(x,T^{k+l}x)\mathop{}\!d\mu(x)=\iint g_{k}(y,x)g_{k+l}(y,T^{l}x)\mathop{}\!d\mu(x)\mathop{}\!d\mu(y)\\ +\mathcal{O}(k^{-\kappa/2}+k^{2(d+3)\kappa/\alpha}e^{-\tau_{r}k}).$
By Condition M.2, we obtain that
$\int_{X}|g_{k}(y,x)g_{k+l}(y,T^{l}x)-\mathbf{1}_{B(y,r_{k})}(x)\mathbf{1}_{B(y,r_{k+l})}(T^{l}x)\mathop{}\!d\mu(x)\lesssim\varepsilon_{k}^{\alpha_{0}/2}\lesssim k^{-\alpha_{0}\kappa/4}$
for $\mu$-a.e. $y$. Hence,
$\mu(\hat{E}_{k}\cap\hat{E}_{k+l})\leq\int_{X}\mu\bigl(B(y,r_{k})\cap T^{-l}B(y,r_{k+l})\bigr)\mathop{}\!d\mu(y)\\ +\mathcal{O}(k^{-\alpha_{0}\kappa/4}+k^{-\kappa/2}+k^{2(d+3)\kappa/\alpha}e^{-\tau_{r}k}).$
Using the short returns estimate (3.2), we have that
$\mu(\hat{E}_{k}\cap\hat{E}_{k+l})\lesssim\hat{M}_{k}^{1+\delta}+\mathcal{O}(k^{-\alpha_{0}\kappa/4}+k^{-\kappa/2}+k^{2(d+3)\kappa/\alpha}e^{-\tau_{r}k}).$
By the choice of $\kappa$, we obtain that
$\mu(\hat{E}_{k}\cap\hat{E}_{k+l})\lesssim\hat{M}_{k}^{1+\delta}$
for sufficiently large $k$. ∎
We now prove Lemma 3.6.
Proof of Lemma 3.6.
We have, by the Lebesgue differentiation theorem, that
$F_{k}(y):=\frac{\mu(B(y,r_{k}))}{m(B(y,r_{k}))}=\frac{1}{m(B(y,r_{k}))}\int_{B(y,r_{k}))}h\mathop{}\!dm\xrightarrow[k\to\infty]{}h(y)\quad\text{$m$-a.e.\ $y$}.$
We will now show that $\lim_{n\to\infty}\int F_{k}\mathop{}\!d\mu=\int h\mathop{}\!d\mu$ by using the Hardy–Littlewood maximal function $\mathfrak{M}$, which is defined by
$(\mathfrak{M}f)(x)=\sup_{r>0}\frac{1}{m(B(x,r))}\int_{B(x,r)}|f|\mathop{}\!dm$
for any locally integrable function $f$. It is clear that $F_{k}\leq\mathfrak{M}h$ for all $k\geq 0$. If $f\in L^{p}(m)$ for some $p>1$, then
$\|\mathfrak{M}f\|_{L^{p}}\lesssim_{p}\|f\|_{L^{p}}$
see for example [37, Theorem 1]. In particular, since $h\in L^{p}(m)$ for all $p\in(1,2]$, we have that
$\|F_{k}\|_{L^{p}}\leq\|\mathfrak{M}h\|_{L^{p}}\lesssim_{p}\|h\|_{L^{p}}$
for all $p\in(1,2]$. Since $F_{k}\to h$ $\mu$-a.e. and $(X,m)$ is a finite measure space, this implies that $\lim_{k\to\infty}\|F_{k}-h\|_{L^{1}(m)}=0$. By the Cauchy–Schwarz inequality, we obtain that $\lim_{k\to\infty}\|F_{k}-h\|_{L^{2}(\mu)}=0$.
At the same time
$\hat{M}_{k}=\int_{X}\mu\bigl(B(x,r_{k})\bigr)\mathop{}\!d\mu(x)=\int_{X}m(B(x,r_{k}))F_{k}(x)\mathop{}\!d\mu(x).$
Since $(X,d)$ is a compact Riemannian manifold, we have that $m(B(x,r))=m(B(y,r))(1+\mathcal{O}(r))$ for all $x,y\in X$ and all sufficiently small $r>0$. Thus, for sufficiently large $k$ and any $x_{0}\in X$, we have that
$\hat{M}_{k}=m(B(x_{0},r_{k}))(1+\mathcal{O}(r_{k}))\int_{X}F_{k}(x)\mathop{}\!d\mu(x).$
Let $\varepsilon>0$. Since $\mu(F_{k})\to\mu(h)$ by the previous paragraph, we have that
$\hat{M}_{k}=m(B(x_{0},r_{k}))(1+\mathcal{O}(r_{k}))\Bigl(\int_{X}h\mathop{}\!d\mu+\mathcal{O}(\varepsilon)\Bigr)$
for all sufficiently large $k$. Summing over $k$,
$\hat{\mathcal{M}}_{n}=\Bigl(\int_{X}h\mathop{}\!d\mu+\mathcal{O}(\varepsilon)\Bigr)\sum_{k=1}^{n}m(B(x_{0},r_{k}))(1+\mathcal{O}(r_{k})).$
Notice that the sum on the right-hand side is asymptotically equal to $\sum_{k=1}^{n}m(B(x_{0},r_{k}))$ since the sum diverges (as $r_{k}\gtrsim n^{-\gamma/d}$ for some $\gamma<d$) and $r_{k}\to 0$. Therefore, dividing by $\sum_{k=1}^{n}m(B(x_{0},r_{k}))$, taking $n\to\infty$ and then $\varepsilon\to 0$, we obtain that
$\lim_{n\to\infty}\frac{\sum_{k=1}^{n}\hat{M}_{k}}{\sum_{k=1}^{n}m(B(y,r_{k}))}=\int_{X}h\mathop{}\!d\mu$
for all $y\in X$. A similar argument shows that
(8.32) $\lim_{n\to\infty}\frac{\sum_{k=1}^{n}\mu(B(y,r_{k}))}{\sum_{k=1}^{n}\hat{M}_{k}}=\frac{h(y)}{\mu(h)}\quad\text{$\mu$-a.e.\ $y$}$
The estimate in (8.28) implies that $\hat{\sigma}_{n}^{2}\sim\hat{\mathcal{M}}_{n}=\sum_{k=1}^{n}\hat{M}_{k}$. Since $\mu$ is finite, (8.32) implies convergence in measure: for all $\varepsilon>0$,
(8.33) $\lim_{n\to\infty}\mu\Bigl\{y:\Bigl|\frac{\sum_{k=1}^{n}\mu(B(y,r_{k}))}{\hat{\sigma}_{n}^{2}}-\frac{h(y)}{\mu(h)}\Bigr|>\varepsilon\Bigr\}=0.$
On the other hand, by Markov’s inequality,
$\mu\biggl\{y\in X:\biggl|\frac{1}{\hat{\sigma}_{n}^{2}}\sum_{k=1}^{n-1}\sum_{j=k+1}^{k+\beta_{k}}\operatorname{Cov}\bigl(\mathbf{1}_{B(y,r_{k})},\mathbf{1}_{B(y,r_{j})}\circ T^{j-k}\bigr)\biggr|>\varepsilon\biggr\}\\ \lesssim\frac{1}{\varepsilon^{2}\hat{\sigma}_{n}^{2}}\sum_{k=1}^{n-1}\sum_{j=k+1}^{k+\beta_{k}}\int_{X}\mu\bigl(B(y,r_{k})\cap T^{-(j-k)}B(y,r_{j})\bigr)\mathop{}\!d\mu(y)$
By the short returns estimate (3.2), we have that
$\sum_{k=1}^{n-1}\sum_{j=k+1}^{k+\beta_{k}}\int_{X}\mu\bigl(B(y,r_{k})\cap T^{-(j-k)}B(y,r_{j})\bigr)\mathop{}\!d\mu(y)\lesssim\sum_{k=1}^{n}\beta_{k}\hat{M}_{k}^{1+\delta_{0}}=o(\hat{\mathcal{M}}_{n})$
by Condition $\mathcal{S}_{\gamma,\upsilon}$ with $\upsilon>2(\delta_{0}\dim X)^{-1}$. Hence,
(8.34) $\lim_{n\to\infty}\mu\biggl\{y\in X:\biggl|\frac{1}{\hat{\sigma}_{n}^{2}}\sum_{k=1}^{n-1}\sum_{j=k+1}^{k+\beta_{k}}\operatorname{Cov}\bigl(\mathbf{1}_{B(y,r_{k})},\mathbf{1}_{B(y,r_{j})}\circ T^{j-k}\bigr)\biggr|>\varepsilon\biggr\}=0.$
The remaining terms can be estimated using decay of correlations:
(8.35) $\lim_{n\to\infty}\mu\biggl\{y\in X:\biggl|\frac{1}{\hat{\sigma}_{n}^{2}}\sum_{k=1}^{n-1}\sum_{j=k+\beta_{k}}^{n}\operatorname{Cov}(\mathbf{1}_{B(y,r_{k})},\mathbf{1}_{B(y,r_{j})}\circ T^{j-k})\biggr|>\varepsilon\biggr\}=0.$
Since
$\begin{split}\hat{s}_{n}^{2}(y)&=\sum_{k=1}^{n}\mu(B(y,r_{k}))-\mu(B(y,r_{k}))^{2}\\ &\quad+\sum_{k=1}^{n-1}\sum_{j=k+1}^{n}\operatorname{Cov}\bigl(\mathbf{1}_{B(y,r_{k})},\mathbf{1}_{B(y,r_{j})}\circ T^{j-k}\bigr)\end{split}$
and $\hat{\sigma}_{n}^{-2}\sum_{k=1}^{n}\mu(B(y,r_{k}))^{2}\to 0$, the limits in (8.33)–(8.35) imply exactly that $\hat{\sigma}_{n}^{-2}\hat{s}_{n}^{2}\xrightarrow[]{\mu}\mu(h)^{-1}h$. By (8.29), we have that
$\lim_{n\to\infty}\int_{X}\frac{s_{n}^{2}}{\sigma_{n}^{2}}\mathop{}\!d\mu=1=\int_{X}\frac{h}{\mu(h)}\mathop{}\!d\mu.$
This implies (see e.g. [22, Proposition 4.12]) that $\hat{\sigma}_{n}^{-2}\hat{s}_{n}^{2}\to\mu(h)^{-1}h$ in $L^{1}$. In particular, we have that
$\int_{X}\biggl|\frac{\hat{s}_{n}}{\hat{\sigma}_{n}}-\frac{\sqrt{h}}{\sqrt{\mu(h)}}\biggr|\mathop{}\!d\mu\lesssim\int_{X}\Bigl|\frac{\hat{s}_{n}^{2}}{\hat{\sigma}_{n}^{2}}-\frac{h}{\mu(h)}\Bigr|\mathop{}\!d\mu$
since
$\biggr|\frac{\hat{s}_{n}}{\hat{\sigma}_{n}}+\frac{\sqrt{h}}{\sqrt{\mu(h)}}\biggr|\geq\frac{\sqrt{h}}{\sqrt{\mu(h)}}\geq\sqrt{\frac{C_{h}}{\mu(h)}}>0$
since $h\geq C_{h}$ by assumption. Hence, $\hat{\sigma}_{n}^{-1}\hat{s}_{n}\to\sqrt{\mu(h)^{-1}h}$ in $L^{1}$ as desired. ∎
9. Discussion of the Short Returns Estimate
We devote this section to a discussion of the short return estimates, (3.1) and (3.2), present in Sections 2.2 and 3. We compare (3.2) with known results in order to motivate (3.1). We assume that $\mu$ is absolutely continuous with respect to $m$. The method goes back to Collet [8] (see Corollary 2.4, Lemma 2.5 therein) who establishes an a.e. version of (3.1) for some non-uniformly hyperbolic maps on the interval. More precisely the estimate is of the following form: for $\mu$-a.e. $x$, there exist $C,N,p,\delta>0$ such that, for all $k>0$ and $l=1,\ldots,(\log k)^{1+p}$,
(9.1) $\mu\bigl(B(x,r_{k})\cap T^{-l}B(x,r_{k})\bigr)\leq C\mu(B(x,r_{k}))^{1+\delta}.$
Similar results can be found in [18, Lemma 4.1], [19, §6.2], [17, §13], which modify Collet’s argument.
Our estimate (3.2) is weaker than (9.1) in the sense that we can allow for a positive measure set in which (9.1) fails. However, we require the constants $C,N,p,\delta$ to be uniform in $k,l$.
Lemma 9.1.
Suppose that $\mathop{}\!d\mu=h\mathop{}\!dm$ for a density $h$ such that $c\leq h\leq C$ for some constants $c,C>0$. Let $\hat{E}_{k,l}=\{x\in X:d(x,T^{l}x)<2r_{k}\}$ and suppose that there exists $\delta_{0}>0$ such that, for all $k,l>0$,
(9.2) $\mu(\hat{E}_{k,l})\lesssim\hat{M}_{k}^{\delta_{0}}.$
If $\hat{M}_{k}\lesssim(\log k)^{-\upsilon}$ for some $\upsilon>2\delta_{0}^{-1}$, then there exist $p_{0}>1$ and $\delta_{1},\eta_{1}>0$ such that, for all $k>0$ and $l=1,\ldots,(\log k)^{p_{0}}$,
(9.3) $\mu\bigl\{x\in X:\mu\bigl(B(x,r_{k})\cap T^{-l}B(x,r_{k})\bigr)>\hat{M}_{k}^{1+\delta_{1}}\bigr\}\lesssim\hat{M}_{k}^{\eta_{1}}.$
Proof.
Let $p_{0}\in(1,\delta_{0}\upsilon/2)$ and define
$\tilde{E}_{k}=\{x\in X:d(T^{l}x,x)\leq 2r_{k}\text{ for some }l=1,\ldots,(\log k)^{p_{0}}\}.$
Notice that $\tilde{E}_{k}\subset\bigcup_{l=1}^{(\log k)^{p_{0}}}\hat{E}_{k,l}$. Hence, by (9.2), $\mu(\tilde{E}_{k})\lesssim\hat{M}_{k}^{\delta}$ for some $\delta<\delta_{0}$, where we use that $\hat{M}_{k}\lesssim(\log k)^{-\upsilon}$ and $p_{0}<\delta_{0}\upsilon/2$. Let $\delta_{1}<\delta$ and define
$F_{k}:=\bigl\{x\in X:\mu\bigl(B(x,r_{k})\cap\tilde{E}_{k}\bigr)\geq\hat{M}_{k}^{1+\delta_{1}}\bigr\}.$
Define the Hardy–Littlewood maximal function $\mathfrak{M}_{k}$ for $\mathbf{1}_{\tilde{E}_{k}}h$ by
$\mathfrak{M}_{k}(x)=\sup_{r>0}\frac{1}{m(B(x,r))}\int_{B(x,r)}\mathbf{1}_{\tilde{E}_{k}}h\mathop{}\!dm.$
Notice that if $x\in F_{k}$, then $\mathfrak{M}_{k}(x)\gtrsim m(B(x,r_{k}))^{-1}\hat{M}_{k}^{1+\delta_{1}}\geq C_{0}\hat{M}_{k}^{\delta_{1}}$ for some constant $C_{0}>0$ by the assumptions on the density $h$. A theorem of Hardy and Littlewood (see e.g. [37, Theorem 1]) states that
$m\{x:\mathfrak{M}_{k}(x)>c\}\lesssim\frac{\|\mathbf{1}_{E_{k}}h\|_{L^{1}(m)}}{c}.$
As a consequence,
$\mu(F_{k})\lesssim m\{x:\mathfrak{M}_{k}(x)\geq C_{0}\hat{M}_{k}^{\delta_{1}}\}\lesssim\mu(\tilde{E}_{k})\hat{M}_{k}^{-\delta_{1}}\lesssim\hat{M}_{k}^{\delta-\delta_{1}}.$
Let $\eta_{1}=\delta-\delta_{1}>0$. Observe that if $y\in B(x,r_{k})\cap T^{-l}B(x,r_{k})$ for some $k>0$ and $l=1,\ldots,(\log k)^{p_{0}}$, then $y\in T^{-l}B(y,2r_{k})$ and $y\in B(x,r_{k})\cap\tilde{E}_{k}$. Hence,
$\mu\bigl\{x\in X:\mu\bigl(B(x,r_{k})\cap T^{-l}B(x,r_{k})\bigr)>\hat{M}_{k}^{1+\delta_{1}}\bigr\}\leq\mu(F_{k})\lesssim\hat{M}_{k}^{\eta_{1}},$
which establishes (9.3). ∎
Remark 12.
The estimate in (9.2) has been verified for several systems, see [18, Lemmas 3.4, 3.10] for Gibbs–Markov maps and non-uniformly expanding interval maps, and [23, Section 6.2] for Linear Anosov diffeomorphisms.
We now show that, by an application of Lemma 9.1, one can obtain (3.2), i.e.,
$\int_{X}\mu\bigl(B(y,r_{k})\cap T^{-l}B(y,r_{k})\bigr)\mathop{}\!d\mu(y)\lesssim\hat{M}_{k}^{1+\delta}$
for some $\delta>0$ and all $k>0$ and $l=1,\ldots,(\log k)^{p_{0}}$. Write $A_{k,l}(y)=\mu(B(y,r_{k})\cap T^{-l}B(y,r_{k}))$ and notice that $\mu(A_{k,l})\leq\mu(B(y,r_{k}))\lesssim\hat{M}_{k}$ by the assumptions on the density $h$. Then, for all $k>0$ and $l=1,\ldots,(\log k)^{p_{0}}$,
$\begin{split}\int_{X}A_{k,l}(y)\mathop{}\!d\mu(y)&\leq\hat{M}_{k}^{1+\delta_{1}}+\int_{\bigl\{A_{k,l}>\hat{M}_{k}^{1+\delta_{1}}\bigr\}}A_{k,l}(y)\mathop{}\!d\mu(y)\\ &\lesssim\hat{M}_{k}^{1+\delta_{1}}+\hat{M}_{k}\cdot\mu\{x:A_{k,l}(x)>\hat{M}_{k}^{1+\delta_{1}}\}\\ &\lesssim\hat{M}_{k}^{1+\delta_{1}}+\hat{M}_{k}^{1+\eta_{1}}.\end{split}$
We therefore obtain (3.2) for $\delta=\min\{\delta_{1},\eta_{1}\}$.
Acknowledgements
I am deeply grateful to my supervisor, Tomas Persson, for his insightful discussions and careful review of numerous drafts.